A Bibliography of Publications about SGML, the Standard Generalized Markup Language: 2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

17 May 2023 Version 2.48

Title word cross-reference

3 [Abi01]. $99 [Kro00a]. R_{XML} [Eri01].

.NET [BFS^{+}02, SS02, SH02].

1.0 [Cas06, Gra00a, Kuz02, Len05].

2 [DS00, Lad01]. 2.0 [Cas06, Kay08, SK02]. 2000 [ACN01, Kro00a, Mar01]. 2002 [ACM02a, B^{+}02]. 2003 [ACM03b, Eri03, FLA^{+}03]. 21st [MKR^{+}01]. 29th [FLA^{+}03]. 2nd [Ano00].

3.0 [Hei01]. 3D [Rah01].

4 [Cas00, Hol00b, RR00]. 4.0 [Del02].

7 [LKB^{+}02, WK03, WK06].

= [Ram03].

ABCs [Des00]. Abridged [FB04]. Academic [KSSS02]. Acceptance [Kit02]. Access [Ano02b, CIK02, DDPS02, Dix01, Gun01a, LMY02, YSLJ02, AD04, BGBJ05, SA03]. Accessibility [Mas02, YSLJ02]. Accessing [NQ02]. According [BGMT02]. Accuracy [Mas02]. ACM [ACM02a, FMA02, SM01]. Acquisition [KSSS02, Lin03]. across [SGW01]. Action [CPJ05, BPW02]. Active [ABM^{+}02, BCP01, Kro00b, SB02, BCP02, ABB^{+}03]. adaptive [CMS02]. Add [Bar01]. Adding [CP02]. Adds [Kro00b]. Adobe [Ano02c]. Advanced [PAB02]. Advances [FLMS05]. Affordable [Kro00b]. Agent [GMRU02]. Agere [Ano02c]. Aggregation [CFF^{+}02]. Aggregators [MRRWW04]. Agile [HBH^{+}03]. Agreement [Bar01]. AI [Coc01]. aided [Fie00, LSS01]. Ails [Eri01]. Ajax [CPJ05]. Algebra [JLST02, KSK^{+}02, LN02]. Algebraic [Gut04]. Algorithm [Bar01, HSJJ02a, XWW^{+}02]. Algorithms [Nør02, SM02]. Almost [dTU04]. Alone [Pad02]. Altera [Ano02b]. Amaya [QV02]. ambient [Jon03b]. amidst [Sid02]. AMPS [Lin03]. Analysis [BPW02, Bar01, GLS^{+}02, HKYU02b, NLB^{+}02, ACM01, Ano03, Fie00, TMK05, WK03, Hun02]. Analytical [BL05]. Analyzing [HBH^{+}03]. Anatomy [FHK^{+}02]. animation [Dan00]. Annotation [LKB^{+}02, TMYU02]. Announced [Coc01]. ANSI [Dav03]. APEX [CMS02]. API [Mus01, Ano03, Mun00]. Applets [Hei03, Fre01]. appliance [Ano03]. Application [Abi01, KYU02a, Kro00b, LR02, MlF01, NR02, PWK02, Rah01, Sch02, TEM^{+}01, Bea02, ME01, Roc01, SK02, Wil02b]. Applications [Abb02, Ano02c, BFS^{+}02, Cer02, KL02, Kro00a, Lea00, MKR^{+}01, Mor00, FP00, Gra00b, Lar03, LCZ04, Luc00, Mue00, SH02, SM02, MRRWW04]. Applied [SS02]. Applying [AA04, HCC^{+}02, LB03]. Approach [DHA^{+}02, FMP02, For08, Koc03, LMMT02, NQ02, AL03, AL05, Le00, TW05, YASU01, YWL^{+}03]. Approximate [GJK^{+}02, LP02, AYFSX03a, GJK^{+}06]. Arabic [NR02]. Architects [Bea03]. Architecture [AvM02, Lut02, BGBJ05, DM03]. Archive [Bor02]. Archives [WS02]. Aren’t [MKR^{+}01]. Art [Coc01]. ASP [BBB^{+}00]. ASP.NET [Wah02]. Assertion [JSSM04]. Asset [Kro00a]. Association [BCKL02]. Astronomy [Bar01]. Atlantic [Bar00]. ATLAS [CL04]. ATM [Lut02]. Audio [Rol00]. August [B^{+}02, BCH^{+}05]. aural [Rol00]. Auswertung [Fie00]. Author [Min02b, BCF01]. author-X [BCF01]. Authoring [HCC^{+}02, Kuz02]. Authorizations [CIK02]. authors [App00, Wil02b]. Automata [Koc03, Nev02, YWL^{+}03, GGM^{+}04]. Automata-based [Koc03]. Automatic [dTU04]. Automatically [ETL02, JLP04]. AutoWeb [FP00]. AVoN [RS00]. Award [Eri03]. aware [WK06]. AXL [RS00].

Backup [Ano00]. Bad [MKR^{+}01]. Balancing [HBH^{+}03]. Bandwagon [SSC^{+}00]. Bang [Sea00]. Barbara [SM01]. Base [VAS02, FHK^{+}02]. Based [ABFS02, AvM02, BBSW03, CLCC02, Cha02b, DLS^{+}03, FMPL^{+}03, Gös03, HKYU02b, HCC^{+}02, HS02b, JSSM04, KL02, KKK02, Kun02, LMY02, LMP02, LWY^{+}02, Mas02, May02, MLLA03, NQ02, NZ02, Pan02, PAB02, PSK02, QN02, SC02, UIN02, WL04, Wil03c, XWP02, XYW02, YKDC02, AD04, BEH^{+}06, BMKL02, Beh00, BGBJ05, Fal00a, Fal00b, FCD02, GSH03, Koc03, KP04, Kro00a, Law04, Le00, LCZ04, Luc00, Mam01, PWK02, SG02a, SCG01, SDC04, WD02, YASU01, Zhe03]. Bases [B^{+}02, FLA^{+}03]. Basic [BD00, SH02]. basiertes [Beh00]. Basis [Fie00]. BBQ [Mun00]. BEA [Ano03]. Beats [Bar01]. been [Whi01]. beginner [Mer01, Nie01]. Beginning [BGR^{+}00, CH00, HC^{+}00, H^{+}07]. Being [Fox02]. Bell [Bar00]. Ben [Ano00]. Benchmark [SWK^{+}02, SY04]. Berlin [FLA^{+}03]. Bertinoro [ABD03]. best [Dan00]. bestselling [Nie02b]. better [Gra00b]. between [DJM02, LZZ03, Nay02a, Nay02b, Wil02b]. Beyond [Abb02, BSL00, CKN03, Nie01]. Bible [Ano00, Har01]. bibliographical [Jak04]. Binding [Ano02c, Bro03, FMPL^{+}03, GLFO^{+}03, TL04, McL02, VRW^{+}03]. biomedical [SRCV06]. Bisimulation [Ram03]. black [Hol00a]. Blake [Sem02]. Block [FMPL^{+}03]. Block-Structured [FMPL^{+}03]. Blocks [Sid02]. blueprint [Hud08a, Hud08b]. Board [Bar01]. Book [Ano00, Kle03, Law04, Mur00, Sem02, Hol00a]. books [Mur00, Nie02b]. Bookshelf [Cro01, Lut02, Ass02, Wil00, Wil01b, Wil01c, Wil03a, Wil03b, Nie02b]. bound [IHW02]. bounds [Seg03]. Breeze [Ano02c]. Bridging [FKS^{+}02c, FKS^{+}03]. Briefs [Lea00]. Bringing [Doo02b]. Broadband [Lut02]. Brokering [DF03]. Browser [Hun02, Kro00b, RS02]. Browsers [Car02a]. Browsing [TMYU02, Mun00]. BSML [VRW^{+}03]. Build [Kro00a]. Building [Ano00, Bar00, Bav00, Edg01, FA00, Flo00, Mue00, PPV02, Sid02, TMK05, Gra00b, Hud08b, Lut02, Kle03]. bulkload [KM06]. bus [Zhe03]. bus-type [Zhe03]. Business [Bar01, Dau03, DJM02, GLFO^{+}03, KW02].

C [Arc02, Bav00, CL04, MKR^{+}01]. C/ATLAS [CL04]. Caching [Tur02a, YLH03, CRW02]. calculus [Jac03]. Calif [ACM03b, SB00]. California [ACM03a, SM01]. calling [RS00]. Can [Dav01, Dav03]. Cantabria [Gut04]. Capturing [Liu02]. cards [VR06]. Cascading [Mey00, Mey01a]. Case [CL04, NZ02, UIN02, Roc01]. Case-Based [UIN02]. Castle [FLMS05]. catalogue [VR06]. Catches [Bar01]. Categorial [JLP04]. CD [Nie02b, Ass02]. CD-ROM [Nie02b]. CDF [LP02]. CDuce [BCF03]. CE [SSC^{+}00]. Center [Ano02a, HG01]. Centric [KSK^{+}02, BCF03, RHC^{+}06]. Century [MKR^{+}01]. CGI [Mor00]. challenged [Kro00a]. challenges [BL05, KP05]. Chancen [LSS01]. Change [XWW^{+}02]. changing [Wil02b]. channel [SA03]. Channels [Kro00b]. Charles [Ano00, GP01a, GP02, GP04]. Checking [DLS^{+}03, LYT^{+}05]. Chemical [GMRRW01, MRR01, MRR03, MRRWW04, GMRRW01]. Chicago [Ano02a]. China [B^{+}02]. Clarification [GWT^{+}01, HBH^{+}03]. Clark [Kim01]. Clash [Lut02]. Classics [Wil00]. Clearly [Sta00, Sta01]. clustering [KM06]. CML [CB04, MRR03]. CMLDOM [MRR01]. Code [Geo02, Kro00b]. Coding [End00]. Colby [Coc01]. ColdFusion [Ano02c]. Collaborative [BCD^{+}02, HRB^{+}02, QN02]. Collecting [Joh02]. Collections [KSH02]. College [Bar01]. Combination [Sea00]. Combinatory [JLP04]. Combines [Lea00]. commentaries [Wil02a]. commentary [Bur02a, Hay02, Pay02]. communication [Int00, Str02]. companion [Bra00]. comparison [AD04, Fre01]. Compass [GMRU02]. compilation [PMK^{+}06]. Complete [Hab02, Sta03, Sta06, Sta07, Bin03, Ehl00, McF00, Pow01, Wil01a]. complexity [GKPS05, MN05, Seg03]. Component [Hei01, LN02, TEM^{+}01]. Composing [LBN03]. Comprehensive [Goo02]. Compressed [ABC^{+}03b, BGK03, YSLJ02]. compression [MPC03]. Computation [Gut04]. Computational [TB04, WJ02]. Computer [Bar01, Fie00, LSS01]. Computer-aided [Fie00, LSS01]. computergestutzte [Fie00]. computergestutzter [LSS01]. Computers [Coc01]. Computing [Bar01, CDF01, KP04, Kro00a]. Concept [Kro00b]. concepts [TB00]. Conceptual [DJM02, EWH^{+}02, LAG02, MLMT02, FCD03, FCD04]. Concerns [VN03]. condition [BPW02]. conditions [Wil02b]. Conference [ACM03a, Ano02a, ABD03, B^{+}02, FMA02, FLA^{+}03, SM01]. Configuration [Sin00, SDC04]. Conflict [KLL02]. Conformance [Int00]. conforming [ZWG^{+}03]. connections [RHC^{+}06]. Connexions [HRB^{+}02]. consequences [ACMS06]. Conserving [SW02b]. Consistency [Ng02, AFL02b].

Consortium [Bar01]. constraint [FKS02a]. Constraints [AFL02a, HSJJ02a, RRB03, DT05, FL02, LC00, LYT^{+}05]. Constraints-Preserving [HSJJ02a]. Constructing [JLP04]. containment [DT05]. Contemporary [Bor02]. Content [BLS02, Cha02b, Doo02b, For08, HRW02, KSSS02, Man02b, Sim02b, Tur02a, XYW02, GSH03, SCG01, VR06]. Content-Based [Cha02b]. Content-Faithful [HRW02]. Contents [HKYU02b]. Contest [Coc01]. Context [PD02, ACMS06, Beh00, MB06]. Control [Cox01b, DDPS02, LMY02, SSC^{+}00, YSLJ02, AD04, BGBJ05]. Controlled [Coc01]. Controlling [Kro00a]. Controversy [Kro00b]. conversational [Luc00]. Conversion [SW02b, HG01]. Converting [Joh02]. cookbook [Man02a]. Cooperative [MEC02]. CORBA [Die01, EF02, Lut02, TEM^{+}01, Zhe03]. Core [Lut02, SR00]. Corporate [Gur02, FA00]. Correction [TEM^{+}01]. correctness [CGMS04]. Correlating [GK03]. Cost [BEH^{+}06, PRP02]. Cost-based [BEH^{+}06]. COTS [BCD^{+}02]. count [FHR^{+}02]. Course [Hei03]. Courseware [QN02]. COVAX [Bor02]. Covering [Ram03]. CPRM [GWT^{+}01]. Creating [Jan01, PSK02, Ray01, Aga02, Car00, CK01, McF00]. Credit [CNB^{+}02]. Cross [Car02a, See02]. Cross-Platform [Car02a]. Crowder [Ano00, Ano00]. Crusoe [Lea00]. Crypto [CNB^{+}02]. CSS [Goo02, Hud08a, Hud08b, Mey01b, SR00, Tea01, Kle03]. cultural [ACMS06]. Culture [Bor02, Lut02]. Curl [Coc01, Müf01]. Curley [Ano00]. Current [Car02a]. Curriculum [HRB^{+}02]. Custom [Kro00b]. Customizable [Kro00b]. Customized [EWH^{+}02, Le00]. Customizing [BFH^{+}02]. CYNTHIA [RS02].

D [Abi01]. Dagstuhl [FLMS05]. Daniel [Ano00]. Dashboard [Kro00b]. Data [ACM03a, ABS00, ABM^{+}02, AA04, Ano02c, ABC^{+}03b, AvM02, Bea03, BBSW03, B^{+}02, BCKL02, Bro03, BS02, CFF^{+}02, CP02, Coy02, Cro01, ETL02, FMPL^{+}03, FMA02, FB04, FLA^{+}03, GLFO^{+}03, HS02b, KYU02b, KSK^{+}02, KP02a, Koc03, KPSS02, LL02b, LZZ03, NQ02, NLB^{+}02, NJ02, PWK02, PG02b, Ray01, RP02, SSC^{+}00, SWK^{+}02, SM01, SMM02, TL04, WL04, XYW02, AL03, AL05, BP05, BMKL02, BM06, BCH^{+}06, CMS02, CDF01, Fal00a, Fal00b, FKS^{+}02b, GK03, GJK^{+}06, HRL^{+}05, HS02a, IHW02, Jon03b, KP05, LFG^{+}01, Liu00, LKB^{+}02, McL02, MMJ^{+}01, MAA^{+}03, MAA^{+}05, MPC03, Myl02, NACP01, RM06, SSB^{+}01, TMK05, TW02, VRW^{+}03, WPFY03]. Data-Binding [Ano02c]. data-intensive [HRL^{+}05]. Database [ACM02a, ACM03b, BCH^{+}05, Coc01, GMW00, LWY^{+}02, Sye02, WL02, AKYJ03, Fei05, HRL^{+}05, JAKC^{+}02, Qui00, SVMAM04, TVB^{+}02, WK03, WK06, BCH^{+}05]. Databases [AJEM02, BHK^{+}03, Clu03, CKS02b, KLL02, KC02, Lew02, Nør02, Oba03, Psa02a, AMN^{+}03, BP05, CZ01, GA03, PG02a, Thu02, VFMMP06, Whi01, YASU01]. Days [Cro01, LCT01]. DB [MB03, Psa02b, TMK05]. DB2 [BEH^{+}06, BCH^{+}06, EMS00, Sel02]. Db4XML [SVMAM04]. DBMS [Whi01]. DBMSs [RP02]. Deadlock [GWT^{+}01]. Debuts [Ano02c]. December [FLMS05]. Decision [TD02]. Declarative [BM06, BS02, LL02b]. Defect [Kro00b]. Defect-Tracking [Kro00b]. Defective [Dav01]. Defending [HBH^{+}03]. defends [Ano03]. Defined [KiY02]. defining [AD04]. definition [LC00]. Definitive [Goo02, MK02, Wal02, Mey00, MK00]. Deisgn [Ano02b]. Deliver [WJ02]. Delivers [Ano02b]. Delphi [TEM^{+}01, Hei01]. demand [Tan02]. Demo [Kuz02]. Demonstration [Kun02, BCF01]. denormalized [BP05]. Deployment [Man02b]. Deriving [WS02]. Describing [Ray01]. description [TW05]. descriptions [WK03, WK06]. Design [Gra00a, NZ02, SG02a, Abi01, Bur02a, CL04, FCD02, Hay02, JKA02, Nie01, Nie02b, Pay02, Rah01, SY04, Str02, Wil02a, Wil02b]. designers [App00, Wil02b]. Designing [Bea03, CLL02, SVMAM04, Hud08a]. Desktop [WE02, HM01, HM04]. Detecting [GWT^{+}01, WH02]. Detection [XWW^{+}02]. deterministic [GGM^{+}04]. deutschsprachiger [Fie00]. Developer [Bar01, Cag00, A.01, Mar01, LR02]. developers [Tra00, Wah02, Wil02b]. Developing [LR02, Stu00, Aye00, Lar03, Roc01, Lut02]. Development [Ano02b, CNB^{+}02, Gun01a, HBH^{+}03, HRB^{+}02, HS02b, Kro00a, Lin03, Bea02, DS00, FP00, Gra00b, Qui00, See02, SK02, VFMMP06]. Devices [Kro00a, Por03]. DHTML [Kle03, Dan00, Fre01, Gil00, Tea01]. Dials [Kro00b]. dictionaries [LSS01]. Diego [ACM03a, ACM03b]. Dies [Coc01]. Diff [XWW^{+}02]. Different [LZZ03]. Digital [GLS^{+}02, Kro00a, Kro00b, Mas02, PAB02, LSS01]. digitaler [LSS01]. direct [PMK^{+}06]. Discovery [KP04]. Disk [Kro00b]. display [VR06]. Displaying [BLS02, Sye02]. disputes [Wil02b]. Dissemination [BF02, CFF^{+}02]. distance [GK03, GK05]. Distributed [ABB^{+}03, Cer02, Die01, Gun01a, JSSM04, LMY02, Lut02, HRL^{+}05, Luc00]. Distributing [Bar01]. distribution [ABC^{+}03a]. DM [TMK05]. Dobb [Eri03]. DocBook [BP01, Sta03, Sta06, Sta07]. Document [Cha02b, Int00, KSK^{+}02, KiY02, KSH02, Kuz02, LN02, LCC^{+}02, MEC02, Mor00, KM06, LC00, LYT^{+}05, YLM^{+}05, Fie00]. Document-Authoring [Kuz02]. Document-Centric [KSK^{+}02]. documentation [End00]. Documents [AJEM02, Bav00, BGMT02, BF02, BFH^{+}02, Car02b, CIK02, CVZ^{+}02, CKS02a, DDPS02, HKYU02b, HKYU02a, JWLY03, JLP04, JOKA02, KC02, Law02, LL02a, Liu02, MS03, Nac02, Psa02b, Sim02b, UIN02, WH02, WD02, XWW^{+}02, ABC^{+}03a, AL02, BPV04, BCF01, BL05, CFGR02, CH06, CTZ02, FCD02, Fie00, GA03, GSBS03, MdlFD03, MAC03, Rol00, SV02, Seg03, SSB^{+}01, YF04, YASU01]. doing [KW02]. Dokumentation [End00]. DOM [Goo02, Har03, LWY^{+}02]. DOM-Based [LWY^{+}02]. Domain [WS02, YKDC02]. Domain-Specific [WS02]. Domino [LZZ03, Tam00]. Dournaee [Sem02]. down [MN05]. Dr [Eri03]. Dreamweaver [WE02]. Driven [Hou01, Mas02, FP00, VFMMP06]. DTD [JOKA02, PCK02, WS02, ZWG^{+}03]. DTD-conforming [ZWG^{+}03]. DTDs [BGMT02, CK02, FL02, MLMT02]. dummies [RR00]. Dutch [BHW^{+}02]. Dynamic [ABC^{+}03a, BGMT02, DLS^{+}03, GF02, Goo02, LL02a, Min02a, Sye02, CKM02, Le00, WPFY03, Aye00, CK01]. dynamically [BMS01].

e-AMPS [Lin03]. E-Learning [QV02]. e-services [SGW01, BCP01]. Earned [HBH^{+}03]. Earned-Value [HBH^{+}03]. ebXML [KW02]. EDI [LW04]. edit [GK03, GK05]. Edition [Ano00, Hol01b, Lad01]. Editor [Kro00b, Doo02b]. Editorial [Eri01]. EDK [Ano02b]. Effective [Myl02, Hud08a, SK02]. Efficient [CFGR02, CVZ^{+}02, CTZ02, CKK03, FMS01, Jac03, Koc03, Lut02, MAC03, NLB^{+}02, WL02, XWW^{+}02, YLH03, YSLJ02, ZZY^{+}02, KM06, PMK^{+}06, SM02]. Efficiently [SSB^{+}01, CAYLS03]. Eighth [B^{+}02]. EJB [EF02, TEM^{+}01]. Electronic [HG01, Lin03, Str02]. Elements [For08, GF02, St.00]. Eliminate [Bar01]. Embeddable [Jon03a]. Embedded [Ano02b, Sea00]. Embedding [RW02b]. embeddings [GK03, GK05]. Embrace [CNB^{+}02]. Emerging [HBZ06]. Empowered [Gur02]. Enable [Car02a]. Enabled [Edg01, RS00, SGW01]. Enabler [PE02]. Enabling [Abb02, Ano02c, Hei03]. Encoded [LL02a]. Encrypt [Dav01]. encrypted [FJ04]. Engine [CMKS03, Pad02, PWK02, SC02, XYW02, IHW02, SVMAM04, TW02]. Engineering [CNB^{+}02, HRB^{+}02, KKK02, LCC^{+}02, Tra00, ACM01, SA03]. English [ED00, Edd00, ED01]. Enhanced [Kle03]. enhances [Ano03]. Enhancing [Cas06]. Enough [CNB^{+}02]. Enterprise [MlF01, Sta01, BGBJ05, Tan02, Sta00]. enterprise-wide [BGBJ05]. Entity [Psa02a, Psa02b]. Entity-Relationship [Psa02a, Psa02b]. Environment [Hab02, LL02a, LZZ03, Müf01, Nay02a, Phi01]. environments [CL04, VRW^{+}03]. equality [TW05]. ERX [Psa02a, Psa02b]. ERX-QL [Psa02b]. Essence [SW02a, SW03]. Essential [Aye00, BSL00, Man02b, SG02b, Tur02b]. Essentials [Cer02, Eis02, Fit01, SW01]. Estimation [Kro00a, LWP^{+}02, PRP02]. evaluating [EP05]. Evaluation [GS03, LWY^{+}02, DAF^{+}03, FMS01, FLMS05, GKPS05]. Evaluator [Kun02]. event [BPW02]. event-condition-action [BPW02]. Evolution [QN02]. Evolves [Lea00]. Evolving [BGMT02]. Exact [KPSS02, AYFSX03a]. example [Mar00]. Excellence [Eri03, Lut02]. excellent [GT00]. Exchange [LZZ03, BDG^{+}03]. Exchanging [MAA^{+}03, MAA^{+}05]. execution [HRL^{+}05]. Expand [Lea00]. Expanded [AY08]. Expeditious [YHL02]. Experience [Man02b, Psa02a]. expert [Hud08b]. Explained [Sta00, Sta01]. Explanations [NR02]. Explosive [Sea00]. Expression [Doo02b, ZZY^{+}02, HVP00, HP01]. Expressions [Joh02, CFGR02, YF04].

Expressive [Koc03]. Extend [DL08]. Extender [EMS00]. eXtensible [Fie00, Des00, Fie00, LS03]. extension [CH06]. external [MMJ^{+}01]. Extracting [CK02, ETL02, JOKA02, LL02b, NQ02, YF04]. Extraction [HKYU02b, KiY02, Lew02, WL04, YKDC02, MdlFD03, MAC03, Myl02].

F [GP01a, GP02, GP04]. F2 [AJEM02]. F2/XML [AJEM02]. Factor [Ano02c]. Failures [Bar01]. Faithful [HRW02]. Fashion [GF02]. Fast [Aye00, RM06]. February [ABD03, SB00]. Federations [PRP02]. Feedback [HBH^{+}03]. Feel [Kro00a]. fetching [Bur02b]. Fifth [SB00]. file [Beh00]. files [Mar04, Mar05]. filtering [CFGR02, DAF^{+}03]. Filters [KP04]. Fine [DDPS02]. Fine-Grained [DDPS02]. Firewalls [Ano00, HBH^{+}03]. First [ACM02a, Bur02a, FMP02]. Fist [Kro00b]. Fix [TEM^{+}01]. Flash [Dan00]. Flexibility [CP02, dTU04]. Flexible [CKS02b, KP02b, SDC04]. Flynn [Wig00]. FO [Paw02]. Foresight [HBH^{+}03]. form [AL02]. Formal [BB02, NR02]. formalism [Jac03]. formalism-only [Jac03]. Formalized [Coc01]. Forms [Joh02, AL03, AL05, PPV02]. formulation [Le00]. Forum [CNB^{+}02]. foundations [Die01]. FPGA [Ano02b]. FrameMaker [Ano02c]. Framework [JLP04, KPSS02, Kro00a, SC02, BGBJ05, FKS^{+}02b, Sah01]. Free [Ano02b]. Freedom [DL08]. Frenzy [GWT^{+}01]. Full [Cas06]. Full-Text [Cas06]. fully [Dav03]. Function [PMC02, CH06]. Fundamentals [BFS^{+}02]. Future [Whi01].

Gains [VN03]. Gateway [OJCH02]. Gauges [Kro00b]. Gene [TMYU02]. GeneAround [TMYU02]. general [BCF03]. general-purpose [BCF03]. Generalized [Int00]. generate [Tan02]. Generated [JLP04, Tur02a, BMS01]. Generating [JOKA02, QN02, VR06]. Generation [Geo02, Gra00a, LN02, PMK^{+}06, YKDC02, BM06]. generator [BMKL02]. generators [Cle01a, Cle01b]. generic [SY04]. Genetic [End00]. Geometric [KPSS02]. German [Beh00, End00, Fie00, LSS01]. German-language [Fie00]. Germany [FLA^{+}03, FLMS05]. Give [CNB^{+}02]. global [KW02, dTU04]. GNOME [GWT^{+}01]. go [LFG^{+}01]. goes [BCH^{+}06]. going [Whi01]. Goldfarb [GP01a, GP02, GP04]. Good [Paw02, EM02]. goodies [Bur02c]. Gopher [Mam01]. Grained [DDPS02]. Grammar [JLP04]. grammars [BB02, PMK^{+}06]. graph [DL04, PG02a]. graph-structured [PG02a]. graphical [CDF01]. graphics [Nie01]. Graphs [KiY02, PG02b, TD02]. Grids [SC02]. Grouping [PAKJ^{+}02]. GTRBAC [BGBJ05]. GUI [Hei03]. Guide [Gol09, KKK02, MK02, Sta03, Sta06, Sta07, App00, A.01, Des00, GR02, Har03, HS02a, Leh02, Mar01, McF00, Mer01, Mey00, MK00, Nie01, Tur02b].

Handbook [Ano00, Cag00, Edg01, GP00, Cha00, CSK01, GP01a, GP01b, GP02, GP04, San03]. Handicapped [RS02]. Handling [MLMT02]. hands [App00]. hands-on [App00]. Handwriting [WW02]. Hard [AFL02a]. harness [SH02]. Healthcare [Eri01]. Help [Kle03]. Heterogeneous [KP02a, NQ02, SGW01]. Hickory [Ano02a]. hierarchical [Dav03]. High [DF03, Mas02, SW02b, DAF^{+}03, SVMAM04]. High-Level [SW02b]. high-performance [DAF^{+}03, SVMAM04]. High-Volume [DF03]. Highly [XWW^{+}02]. Histogram [LWP^{+}02]. History [FMPL^{+}03]. Hits [Kro00b]. Holistic [BKS02, JWLY03]. Hong [B^{+}02]. Hopes [Bar01]. Hosted [MKR^{+}01]. Hotel [Abi01, Rah01]. HTML [Kle03, Aga02, AF02, App00, Aye00, Bav00, BMS01, Bur02b, Bur02c, Cal00, Car00, CK01, Cas00, CNB^{+}02, Del02, Ehl00, ETL02, FMP02, Fun00, Goo02, GT00, Gra00a, Hol00a, Hol00b, Hud08a, Jac03, JKA02, Jan01, Kit02, Knu01, Le00, Lea00, Leh02, LCT01, Liu02, LHO08, MEC02, Mer02, Mor00, MK00, MK02, Nie00, Nie01, Nie02a, Pow01, PW01, RR00, RS02, RW02b, TB00, UIN02, WH02]. HTML-based [Le00]. Humanities [HG01]. hybrid [BCH^{+}06]. HyO [DL04]. HyO-XTM [DL04]. hyper [DL04]. hyper-graph [DL04]. Hyperbolic [LP02]. Hypermedia [WD02, SA03]. Hypertext [LSS01, LSS01]. Hypertextualization [Mas02].

IBM [Doo02b, Sel02]. IDE [Kro00b]. idiot [McF00]. IEC [Int00]. II [Ang00]. III [GMRRW01]. im [Beh00]. Immediate [KLL02]. impact [SRCV06]. Impaired [PE02]. Implementation [Mam01, NZ02, SG02a, WD02, Die01, LB03]. implementations [CZ01]. Implementing [CLCC02, Har02, SS02]. Importance [Fox02, KM06]. improved [Qui00]. incomplete [ASV06]. Incremental [BPV04]. Index [ZZY^{+}02, ACN01, CMS02, WPFY03]. Indexed [CVZ^{+}02, JWLY03]. Indexes [Ram03, GGM^{+}04, TW02]. Indexing [Gun01b, HLM03, LL02a, LWY^{+}02, Sye02, CK02]. industries [HBZ06]. Industry [VN03]. INEX [FLMS05]. Inference [TD02]. Information [For08, HKYU02a, Int00, JLP04, JOKA02, Kro00a, Law04, Sel02, YKDC02, ASV06, AL03, AL05, Beh00, CK02, FLMS05, HBZ06, HYC04, MdlFD03, SRCV06, Tan02, MRR01]. information-theoretic [AL03, AL05]. Informationen [Beh00]. infrastructure [CL04]. Ingredients [Man02b]. initiative [FLMS05]. Inlining [PMC02]. Installations [Kro00a]. Instant [Sea00]. Instrument [Cox01b]. Integral [Kun02]. integrate [Dav03]. Integrated [Ng02, Mun00, Ano02c, Lut02]. Integrating [BCH^{+}06, EMS00, GJK^{+}06, Psa02a]. Integration [ABM^{+}02, Bea03, KP02a, Kun02, MlF01, PLM^{+}02, Sel02, dTU04, HYC04]. Integrity [Ano02b, FL02]. Intelligence [WL04]. Intelligent [LN02, YKDC02]. intensional [MAA^{+}03, MAA^{+}05]. intensive [HRL^{+}05]. Interaction [OJCH02, Rol00]. Interactions [Nay02a]. Interactive [LKB^{+}02, WW02, Aye00, Knu01, Müf01]. Interchange [AvM02, VRW^{+}03]. Interconnecting [NQ02]. interface [Mun00]. Interfaces [Jan01]. International [ACM03a, Ano02a, ABD03, B^{+}02, BCH^{+}05, FMA02, FLA^{+}03, Gut04, SM01, FLMS05, YLM^{+}05]. Internationalization [Sav01]. Internet [KW02, Law04, LSS01, Hou01, Knu01, Kro00a, LSS01, Müf01, SM02]. Internet-challenged [Kro00a]. Interoperability [DJM02, TEM^{+}01, SRCV06]. Interpreter [NZ02]. Interview [Wig00]. Intonation [JLP04]. IntraText [Mas02]. Introduction [Knu01, GT00, Lar03]. introductory [Car00, CK01]. Inversion [LP02]. IP [Lut02]. IR [HKYU02b]. IR-Based [HKYU02b]. ISAAC [Gut04]. ISO [Int00]. ISO/IEC [Int00]. issues [KP05]. Italy [ABD03].

Jabber [Ada02]. James [Kim01]. Java [Ahm01, GDB02, Fre01, AF02, Ano02b, Ano02c, Ano03, Bar01, Bur01, CLCC02, CZ01, Cle01a, Cle01b, DS00, Die01, Dwe00a, Dwe00b, EF02, Fal00a, Fal00b, Gri02, Har03, Hei03, JSSM04, Kro00a, Kun02, Lad01, LCZ04, Lin03, LZZ03, Mam01, McL00, McL01, McL02, McL07, MlF01, Roc01, SG02a, Tam00, WL04]. Java-Based [JSSM04, Fal00a, Fal00b, SG02a]. Java-XML [Lin03]. JavaScript [AF02, Gil00, Goo02, GT00, Knu01, Tam00, TEM^{+}01, TB00]. JAXP [Gri02, Har03].

JDOM [Har03]. Jelly [Gös03]. Joins

[CVZ^{+}02, JWLY03, TD02, BKS02, GJK^{+}02, GJK^{+}06]. JSP [QN02, Roc01]. July [Gut04]. June [ACM01, ACM02a, ACM03a, ACM03b, Ano02a, FMA02]. Just [Sim01b, Sim02a, Sah01].

KDE [GWT^{+}01]. Kenneth [Coc01]. key [LYT^{+}05]. Keys [BDF^{+}02a, BDF^{+}02b]. Keyword [BHK^{+}03, KYU02a, WL02, GSBS03]. KF [XWW^{+}02]. KF-Diff [XWW^{+}02]. Kit [Ano02b]. Knowledge [ABD03, HCC^{+}02, VAS02, YKDC02, DL04]. Kong [B^{+}02]. Kontext [Beh00]. Krause [Ano00]. Kuenringer [Fie00, Fie00]. Kweelt [Sah01]. kyokasho [ASY02].

Labeling [CKM02]. LAN [Ano02c].

Language [AY08, BS02, CKS02b, DJM02, Fie00, Gös03, Gra00a, Int00, JSSM04, May02, NR02, Pan02, PAB02, PSK02, RRB03, SB00, Uni01, BCF03, Cha02a, Cha03, Des00, HP03, LB03, VRW^{+}03, LS03]. Languages [FMPL^{+}03, Kim01, AD04, BB02, TA04]. Large [B^{+}02, FLA^{+}03, ACM^{+}02b]. large-scale [ACM^{+}02b]. LaTeX2HTML [Yua01]. Latin [HG01]. Lazenby [Ano00]. Learn [Cal00]. Learning [Fit03, Hab02, HCC^{+}02, Nie01, QV02, Ray01]. Leave [LAG02]. Legal [BHW^{+}02, MdlFD03]. LegoDB [BFH^{+}02]. Lehre [Beh00]. Lesson [FMPL^{+}03]. Letters [FMPL^{+}03, GLFO^{+}03, GWT^{+}01, HBH^{+}03, MKR^{+}01, SSC^{+}00, TEM^{+}01]. Level [SW02b]. Levelized [KiY02]. Levels [dTU04]. Leveraging [SA03]. Lexical [Mas02]. lexicography [LSS01]. Lexikographie [LSS01]. librarian [Des00]. Libraries [MKR^{+}01, PAB02]. Library [GLS^{+}02, JSSM04, Mas02, VR06]. Lightweight [Jon03a]. line [BDG^{+}03, LWP^{+}02]. Linear [Bar01]. Linux [Ano00, Kro00a, Sea00]. Lisp [XWP02]. Literate [Dwe00a]. Localization [Sav01]. Locating [Sim02b]. LockX [CAYLS03]. Log [GLS^{+}02]. Logging [GLS^{+}02]. Logic [Nev02, TW05]. Logical [RRB03]. Look [Kro00a, Paw02, CZ01, Mar04, Mar05]. Lore [GMW00]. Lotus [LZZ03]. LWP [Bur02b].

machine [Fal00a, Fal00b]. Macromedia [Ano02c]. Madison [ACM02a, FMA02]. Maintaining [Ng02]. Making [Lut02, Müf01, Paw02, RHC^{+}06, YLM^{+}05, EM02, FHR^{+}02]. Manage [LMMT02, Uni01]. Management [ACM03a, AA04, Ano02b, Ano02c, ABD03, Bar00, CLCC02, FMA02, GMW00, Hab02, HBH^{+}03, HCC^{+}02, KC02, Lut02, NLB^{+}02, SWK^{+}02, SM01, SC02, XYW02, DL04, FHK^{+}02, Jon03b, KP05, LFG^{+}01, MMJ^{+}01, SGW01, WK03]. Manager [Kro00a]. Managing [ABB^{+}03, FB04, Por03, CTZ02]. Map [YSLJ02, ZZY^{+}02, DL04]. Maple [BMH02, Kun02, Nay02a]. MapleNet [Man02b]. Mapping [GF02, HSJJ02a, HSJJ02b, Jak04, SMM02, RHC^{+}06]. Mappings [Nay02b]. Mark [Coc01]. Markov [LWP^{+}02]. Markup [BSL00, DJM02, Doo02b, Fie00, GMRRW01, Int00, JSSM04, Kim01, LS03, MRR01, MRR03, MRRWW04, Nay02b, Des00, TA04, VRW^{+}03, YLM^{+}05]. MARS [DT03]. Mason [RW02b]. Mastering [GDB02, NWB00, Tit02, Tid07]. Match [YWL^{+}03]. Matching [AYFSX03b, Dwe00b, AYFSX03a, BKS02, HP01, RM06]. Materialized [ACN01, ZDW^{+}03]. Math [AY08, Min02a, Sye02]. Mathematica [Har02, Sch02, WJ02]. Mathematical [ABD03, Joh02, Man02b, NR02, WW02, YF04]. Mathematics [BLS02, Sye02]. MathML [Doo02b, Ano02a, AK02, BMH02, BLS02, Car02a, Doo02b, Har02, HRB^{+}02, HRW02, Hun02, Joh02, KSSS02, NR02, Nay02a, Pad02, PD02, QV02, RW02a, San03, Sid02, SW02b, XWP02, XW02]. MathPlayer [Min02b]. May [SM01]. me [CNB^{+}02]. means [End00, MdlFD03, Whi01]. Measurement [Ano02b, Ano02c]. Mechanism [KLL02]. MED [MMJ^{+}02]. media [WK03, WK06]. Mediator [ABFS02]. Medical [Mam01, End00]. Medienarchivs [Beh00]. medium [Beh00]. medizinischen [End00]. Mesh [SCG01]. Mesh-based [SCG01]. Message [DF03, PSK02]. Messaging [Sea00]. Meta [LKB^{+}02]. Meta-data [LKB^{+}02]. Metadata [AvM02, MRRWW04, Tan02, FJ04]. metamodels [Tan02]. Metaphor [CNB^{+}02]. Method [LP02, WPFY03]. methodology [FCD02]. Methods [HBH^{+}03, OJCH02]. Metrics [KSH02]. MFC [Kro00b]. Microsoft [ACN01, Ano02c, Mar01]. middle [FMS01]. middle-ware [FMS01]. Middleware [Kro00a]. Middleware1 [BCD^{+}02]. MIME [Dav01]. Mining [BCKL02, MEC02, YLH03, TMK05]. Mithra [GLFO^{+}03]. mittels [End00]. MIX [Liu00]. Mixed [DT03, HBG^{+}03]. MKM [ABD03]. MM [ME01]. MML [TD02]. Mobile [Lea00, PWK02]. Mode [HBG^{+}03]. Model [FP00, GF02, KSK^{+}02, Lin03, Min02a, PLM^{+}02, VFMMP06, Beh00, BFRW02, FKS02a, LW04, Zhe03]. Model-driven [FP00]. Modeling [Dau03, DJM02, EWH^{+}02, PRP02, SB00, AD04, HBZ06]. Modell [Beh00]. Modellierung [Beh00]. Modelling [DHA^{+}02, Beh00]. Models [LMMT02, MLMT02, Psa02a, RRB03, FCD03, FCD04, SM02]. Modification [KP02b]. Modular [CB04, XW02]. Monitoring [Cox01b, NACP01]. Monterey [SB00]. Motif [Kro00b]. Moving [HG01, SD00]. MOWGLI [AK02]. Mozilla [Sid02]. MPEG [LKB^{+}02, WK03, WK06]. MPEG-7 [LKB^{+}02, WK03, WK06]. MSL [BFRW02]. MSXML [TEM^{+}01, Hei01]. Multi [TD02, SA03, ZDW^{+}03]. multi-channel [SA03]. Multi-way [TD02]. multi-XQuery [ZDW^{+}03]. Multidimensional [LMMT02, MLMT02]. Multimedia [HCC^{+}02, PSK02, ME01]. multisignature [LC04]. multiversion [CTZ02]. MX [Ano02c].

Naming [Law02]. Native [Fei05, MLLA03, BCH^{+}06, Dav03, FHK^{+}02, JAKC^{+}02, LFG^{+}01, PAKC^{+}03, SVMAM04]. Natural [NR02, TL04, LB03]. Navigation [GMRU02, SMM02, RS00]. need [Wil02b]. needs [Clu03]. Negation [Ram03]. negotiations [Str02]. Netscape [SSC^{+}00]. Network [CLCC02, Kro00a, Ano03, FCD02, IHW02]. Network-based [Kro00a, FCD02]. network-bound [IHW02]. Networks [Lut02]. News [Bar01, Coc01, Lea00]. Next [Gra00a, Sea00]. NIST [Fal00a, Fal00b]. Node [Koc03]. Node-Selecting [Koc03]. normal [AL02, AL03, AL05]. Norway [BCH^{+}05]. Notations [Mas02]. Note [FKS^{+}03, Mam01, PSK02]. Nothing [SSC^{+}00]. November [Tra00]. Numbering [KYU02b]. numerical [EP05]. Nutshell [WE02, HM01, HM02, HM04].

Object [AJEM02, Dix01, Fox02, GF02, Gun01a, KC02, Min02a, RRB03, RP02, CZ01, CL04, FCD03, FCD04]. Object-Oriented [GF02, CL04, FCD03, FCD04]. Object-Relational [KC02, RP02]. Objects [Dau03, MRR01, Nay02b, SSC^{+}00]. Observation [Wil03b]. Obtain [Psa02b]. OLAP [PRP02]. OLAP-XML [PRP02]. Old [Wil00]. OLE [TMK05]. On-line [LWP^{+}02, BDG^{+}03]. Online [Mas02]. only [Jac03]. Ontologically [ETL02]. ontologies [TW02]. Ontology [ABFS02]. Ontology-Based [ABFS02]. Open [Ano02c, Bar01, Coc01, Mam01, QN02, Qui00, Sid02, SSC^{+}00]. Open-source [Mam01]. OpenBSD [Ano00]. OpenMath [DL08, Nay02a]. OpenOffice.org [Ano02c]. Opera [Cro01]. Operational [Lut02]. Operations [KLL02, DL04]. Operators [Nør02]. opportunities [BL05, LSS01]. Optics [Lut02]. optimal [BKS02]. optimisation [BPW02]. Optimization [Kro00b, LWY^{+}02, BEH^{+}06, ZDW^{+}03]. Oracle [Ano02b, Cha00, Mue00, MB03]. Oracle9i [CSK01]. Order [Mam01]. ordered [TVB^{+}02]. Oriented [Ano02c, GF02, LMMT02, RRB03, CL04, FCD03, FCD04]. OrientStore [MLLA03]. Our [Cro01]. outline [Mer02]. Outsourcing [GLFO^{+}03]. Overview [FB04].

packages [ME01]. Page [Lat02, McF00]. Pages [Ang00, MEC02, Aga02, Bur02b, Car00, CK01, Hud08a]. Panda [Ano03]. panel [LFG^{+}01]. Paper [LCC^{+}02]. papers [FLMS05]. paradigm [BCP01]. Parallel [CNB^{+}02]. Parametric [HFC05]. Parasoft [Ano03]. Parser [NZ02, SG02a]. parsers [PMK^{+}06]. Parsing [Cox01a, GWT^{+}01, Bur02b, Jac03]. Part [Ang00, GMRRW01, MRR03]. Partial [HKYU02b, KLL02]. Passing [MKR^{+}01, PSK02]. Past [Whi01]. PASTE’01 [ACM01]. Path [BGK03, DAF^{+}03, LWP^{+}02, Pan02, ZZY^{+}02, CMS02, CGMS04, YASU01]. path-based [YASU01]. Pattern [BBSW03, CFF^{+}02, Dwe00b, NZ02, BKS02, HP01, RM06]. Pattern-Based [BBSW03, NZ02]. Patterns [YLH03, HBZ06]. PC [WW02]. PCKS#7 [Dav01]. PDF [CNB^{+}02]. Peer [ABM^{+}02, KP05]. Peer-to-Peer [ABM^{+}02, KP05]. People [RS02]. perceptions [Wil02b]. Performance [LWY^{+}02, DAF^{+}03, SVMAM04]. perimeters [Ano03]. Perl [RW02b, AF02, Bur02b, GWT^{+}01, Gun01b, RM02]. Persistence [TL04]. Personal [Coc01]. Persons [PE02]. Perspective [HBH^{+}03]. perspectives [Car00, CK01, LSS01]. Perspektiven [LSS01]. Perturbing [EP05]. Pervasive [KP04]. Peter [Wig00]. PGP [Dav01]. Philological [Mas02]. Philological/Textual [Mas02]. Phone [Edg01]. Phone-Enabled [Edg01]. PHP [MKR^{+}01]. Phrase [AYFSX03b, AYFSX03a]. physical [SDC04]. PIX [AYFSX03a]. Plain [ED00, Joh02, Edd00, ED01]. Platform [Car02a, LCZ04, See02]. Platforms [LZZ03, SGW01]. Platinum [Lad01]. PlayStation2 [Kro00b]. Pocket [Nie02a, EC01, Len05, Mey01b, Nie00, SFE05, WW02]. PODS [ACM02a, ACM03b]. Policy [SC02, BGBJ05]. polymorphism [HFC05]. Portal [Kro00a]. Portals [Gur02, FA00, Tan02]. Portfolio [Ano02b]. Porting [SSC^{+}00]. PostScript [YF04]. power [SH02]. Powerful [Sye02]. Practical [KKK02, CZ01]. predicate [DAF^{+}03]. Preliminary [SW02a]. Preparations [FJ04]. preprocessing [CKK03]. presence [FL02]. Present [Whi01]. Presentation [Doo02b, Nay02b]. Presentations [PSK02]. presenting [Liu00]. Preserving [HSJJ02a, GA03, LC00]. preview [Ano03]. Principles [ACM02a, ACM03b]. Print [Paw02]. Prior [Coc01]. Prize [Bar01]. Probabilistic [NJ02, WS02]. problem [GSH03, VRW^{+}03]. Problems [VAS02]. Proceedings [ACM03a, FMA02, FLA^{+}03, SM01, SB00, B^{+}02, Gut04, ACM02a, ACM03b, BCH^{+}05]. Process [GLFO^{+}03]. Processing [DF03, GGM^{+}04, HBG^{+}03, Har03, KYU02a, Koc03, PCK02, TL04, YHL02, ZZY^{+}02, BL05, Dav03, GK05, HP03, Int00, McG00, YWL^{+}03]. Processor [Ano02b, Lea00, LMP02]. production [LSS01]. Productivity [Ano02c]. Products [Ano02b, Ano02c, Ano03, Kro00a, Kro00b]. Produktion [LSS01]. Professional [ABB^{+}00, BBB^{+}00, Bir01, EST^{+}07, Mar01, Wil02b, Ahm01, BD00]. Program [ACM01, Cle01a, Cle01b, Dei01]. Programmer [Cro01, Wil00, Wil01b, Wil01c, Wil03a, Wil03b, Kay01, Kay08, Mey01a, PW01, SG02b]. Programmers [Hoq00]. Programming [Ada02, Ano00, Coc01, Dwe00a, Eri03, FMPL^{+}03, Hei03, Müf01, STK02, SJD01, Tam00, Wil03c, AF02, FMR02, GDB02, Knu01, SM02, TB00]. Programs [Jan01, EP05]. progress [EM02]. Project [AK02, Bar01, CKS02a, Kro00a, Lin03, BCF01, HG01]. Projecting [MS03]. Prominence [VN03]. Proof [NR02]. Proofs [NR02]. properties [BB02]. Proposal [BHW^{+}02]. Proposed [Bar01]. ProTDB [NJ02]. Protect [Coc01]. Protocol [Lea00, Dix01, Gun01a]. Proximity [BHK^{+}03]. PTDOM [WK06]. publication [LSS01]. Publicon [Kuz02]. Publikation [LSS01]. Publish [Yua01]. Publishing [BP01, Car02b, DT03, CKN03, FKS^{+}02b, LCT01, LBN03, SSB^{+}01]. purpose [BCF03]. Pushing [ABC^{+}03b, BCP02]. Putting [Gun01a, HBH^{+}03]. Python [Ang00, JD02, McG00].

QL [Psa02b]. queriable [MPC03]. Queries [ABC^{+}03b, BGK03, Cha02b, CP02, CKS02a, GS03, KYU02a, Koc03, PMC02, Ram03, CRW02, CKK03, CGMS04, CDF01, DT05, FMS01, Le00]. Query [AY08, BS02, CKS02b, DF03, GA03, HBG^{+}03, LMP02, LWY^{+}02, NQ02, Nør02, Pan02, PCK02, PAB02, WL02, YLH03, Cha02a, Cha03, GKPS05, IHW02, PPV02, RM06, Sah01, YWL^{+}03]. Query-preserving [GA03]. Querying [AKYJ03, ABFS02, BBSW03, JFB05, KPSS02, Liu00, LAG02, May02, MB06, Nac02, Psa02b, RP02, Suc02, ASV06, BP05, CAYLS03, FJ04, Mun00, PAKC^{+}03, Seg03, TVB^{+}02, WPFY03]. Quick [WE02, HM01, HM04, RR00, SG02b]. QuickStart [Gol09]. QuickTime [App00].

R [HLM03]. Rainbow [ZDW^{+}03, Hab02]. Raises [VN03]. Ralph [Ano00]. ranked [GSBS03, TW02]. RDBMS [HLM03]. RDF [NZ02]. reactive [BCP02]. readers [Wil02b]. Reading [PE02]. Real [Hoq00, Hun02, SSC^{+}00]. Real-World [Hun02]. Reality [SB00]. Realization [LZZ03]. really [Dav01]. Reasoning [BDF^{+}02b, TW05]. Recognition [DHA^{+}02, UIN02]. Recognizer [WW02]. recompilation [EP05]. Recovery [Ano00]. Recursive [PMC02]. reducing [CDHZ03]. Redundancy [CDHZ03]. Redundant [DT03]. Reference [Goo02, Nie02a, WE02, Bin03, EC01, HM01, HM04, Kay01, Kay08, Len05, LHO08, Mey01b, Mey01a, Nie00, Pow01, PW01, RR00, SG02b, SFE05, Wil01a]. Reflective [Dwe00b]. Reflexive [BCD^{+}02]. reformulation [DT05]. regime [Bur02a, Hay02, Pay02, Wil02a, Wil02b]. Regular [HVP00, HP01]. related [Int00]. Relational [BFH^{+}02, CKS02b, HSJJ02a, HSJJ02b, KP02b, KC02, Lew02, Oba03, Psa02a, RP02, WL02, AMN^{+}03, AL03, AL05, BP05, BCH^{+}06, BDG^{+}03, CKN03, FKS^{+}02b, FKS^{+}02c, FKS^{+}03, GA03, LFG^{+}01, LC00, SSB^{+}01, TVB^{+}02, YASU01]. Relations [KP02b, ABS00, CDHZ03]. Relationship [Psa02a, Psa02b]. Release [Bar01]. releases [Ano03]. Rendering [Pad02, PWK02, XYW02]. replication [ABC^{+}03a]. Reply [Wil02a]. report [MMJ^{+}02]. reports [PPV02]. repositories [BCP02, Tan02]. repository [ACM^{+}02b, Fal00a, Fal00b]. representation [CL04, Str02, SM02]. Representing [ASV06]. ReScUE [LW04]. Research [PSK02, KP05]. Researcher [Coc01]. Resolution [KLL02]. Resource [Goo02]. resources [Qui00]. Retrieval [Cha02b, HKYU02a, LL02a, LKB^{+}02, Fei05, FLMS05, TW02, YASU01]. Reuse [Bea03, BCD^{+}02]. Reverse [JOKA02, LCC^{+}02]. Review [Kle03, Law04, Sem02]. Reviews [Ano00, Mur00]. Revised [FLMS05]. Revolution [Coy02]. Revolutionary [SRCV06]. Rewriting [CB04]. Ridge [Ano02a]. Riding [SSC^{+}00]. ROLEX [BDG^{+}03]. ROM [Nie02b]. roots [TA04]. Router [Coc01]. routing [GSH03, SCG01]. RPC [Cer02, Jon03a, Por03, SJD01]. RRXF [CDHZ03]. RSA [Ano02b]. RSS [MRRWW04]. Ruby [MKR^{+}01]. rUID [KYU02a]. Rule [DJM02, LS03, May02]. Rule-Based [May02]. Rules [BCKL02, BPW02, BCP01, BCP02].

S [Dav01]. §-calculus [Jac03]. S/MIME [Dav01]. SAML [JSSM04]. Sams [LCT01]. San [ACM03a, ACM03b]. Santa [SM01]. Santander [Gut04]. SAR [B^{+}02]. SAX [Har03]. SAX2 [TEM^{+}01, Bro02, Hei01, Mus01]. Scalable [CFF^{+}02]. scale [ACM^{+}02b]. Scan [LN02]. Scan-to-XML [LN02]. Scenarios [PWK02]. Schaum [Mer02]. Scheduling [Lin03, VAS02]. Schema [AFL02a, Coc01, Dau03, For08, HSJJ02a, HSJJ02b, KL02, KP02b, KiY02, MLLA03, MRR03, PLM^{+}02, XW02, vdV02, Bin03, LC00, Str02, VRW^{+}03, Wal02, WK06, BFRW02, PMK^{+}06, SG02b]. schema-aware [WK06]. Schemas [EWH^{+}02, LMMT02, MB03, SB02, MAC03]. Schemata [FCD03, FCD04]. Scheme [KYU02b, LC04]. schemes [CTZ02]. Scholarly [Mas02]. Science [Bar01, TB04, Yua01]. scientific [HRL^{+}05].

Scripting [Gös03, Mar01]. SDH [Lut02]. SDH/SONET [Lut02]. Search [BHK^{+}03, CMKS03, WL02, Cas06, GSBS03, TW02]. Searching [Gun01b, Suc02]. Second [ACM03b, ABD03, Hay02, Mar05]. Secondary [Koc03]. Secrets [TEM^{+}01]. Secure [BF02, Car02b, Lut02, CH06, CAYLS03, Jon03b, VFMMP06]. Securing [Ano00, BCF01, Her02, Sem02]. Security [JSSM04, Nae03, Dou02, Sem02, Ano02b]. Select [CKS02a]. Select-Project [CKS02a]. selected [FLMS05]. Selecting [Koc03]. selection [ACN01, GSH03]. Selective [BF02]. Selectivity [LWP^{+}02]. Self [LWP^{+}02, Ray01]. Self-Describing [Ray01]. Self-Tuning [LWP^{+}02]. Semantic [CMKS03, KP02a, Law04, LKB^{+}02, Nay02b, UIN02, dTU04, CRW02, FCD02, GMRRW01, LC00, BLS02, Thu02]. Semantics [CB04, Liu02, SW02b]. Semistructured [BBSW03, BS02, ABS00]. September [FLA^{+}03]. Sequence [Bar01]. Sequencing [RM06]. Server [ACN01, Ang00, Cro01, Jan01, Jon03a, Mar01, Sch02, WJ02]. Servers [TEM^{+}01]. Service [KP04, LCZ04, See02]. Services [ABM^{+}02, Cer02, Coc01, Coy02, Gur02, JSSM04, Nae03, PAB02, STK02, WJ02, BCP01, BCP02, LR02, SS02, SGW01, SJD01]. Servlets [EF02]. session [LFG^{+}01]. Set [AY08, BGMT02, Kro00b, DL04]. SETI [Bar01]. Setting [KSSS02]. SGML [Int00, LSS01, LSS01]. SGML/XML [LSS01, LSS01]. Shared [MKR^{+}01, TEM^{+}01]. Sharing [MEC02, DAF^{+}03]. Sheets [Mey00, Mey01a]. Ships [Ano02b]. sibling [KM06]. SIGACT [ACM02a, ACM03b]. SIGART [ACM02a, ACM03b]. SIGMOD [ACM02a, ACM03b, ACM03a, FMA02, SM01]. SIGMOD-SIGACT-SIGART [ACM03b]. Sign [Dav01, JSSM04]. Sign-and-Encrypt [Dav01]. Sign-On [JSSM04]. Signal [Ano02b]. Signature [PCK02]. signatures [CKK03]. signed [GMRRW01]. SIGPLAN [ACM01]. SIGSOFT [ACM01]. SilkRoute [FKS^{+}02b]. Similarity [CP02, For08, KPSS02]. Simple [Dix01, Gun01a, Mus01]. Simplicity [Kim01]. Simplifying [Gun01a]. Simulation [BCD^{+}02, BS02, Ram03]. simulations [SDC04]. Simulators [BCD^{+}02]. Single [JSSM04]. Site [Gun01b, App00, Aye00]. sites [Flo00, Gra00b, Hud08b]. six [Nie02b]. Size [Nay02b]. Skeletons [SMM02]. Slicing [GS03]. Small [Kro00a]. Sniff [Ano02b]. SNMP [OJCH02]. Snowbird [ACM01]. SOAP [Cer02, Cro01, Dix01, EF02, Gun01a, SS02, See02, SG02b, STK02]. social [Wil02b]. Software [Ano02b, AvM02, Coc01, LN02, Sin00, ACM01, Zhe03]. SOI [Ano02b]. SOISIC [Ano02b]. solutions [Stu00, TMK05, Tan02, WK03]. solving [VRW^{+}03]. Some [Wil01c, Seg03]. Sonargaon [Abi01, Rah01]. SONET [Lut02]. Sophisticated [Kro00a]. sorting [CKN03]. Source [Ano02c, Bar01, Kro00b, TEM^{+}01, Mam01, Qui00, SSC^{+}00]. Sourcebook [Gra00a, Gra00b]. Sources [ABFS02, KP02a, NQ02, Tur02a, dTU04, GJK^{+}06]. Spain [Gut04]. Spatio [HYC04]. Spatio-temporal [HYC04]. Special [Hol01b]. Specific [WS02]. Specification [Bar01, BGBJ05]. specifications [AFL02b, LB03]. Specified [ETL02]. Specifying [RRB03]. Speech [Ano02c, Lar03]. Speech-Enabling [Ano02c]. SpeechStudio [Ano02b]. spiders [Bur02b]. spruce [LFG^{+}01]. SQL [ACN01, Mar01, BCH^{+}06, Dav03, EM02, FMR02, ME01, MMJ^{+}01, MMJ^{+}02, MB06, Pan02]. SQL/MED [MMJ^{+}02]. SQL/MM [ME01]. SQL/XML [EM02, FMR02, MB06]. SQL4X [CKS02b]. SquareList [GLFO^{+}03]. SSML [JLP04]. Stand [Pad02]. Stand-Alone [Pad02]. Standard [GLS^{+}02, Int00, KW02, Myl02, BHW^{+}02]. Standardized [HCC^{+}02]. Standards [HBH^{+}03, Nae03, QN02, Sid02, Wig00, Jak04]. States [SM01]. Static [BMS01, DLS^{+}03]. statically [HP03]. Statistical [DHA^{+}02, PG02a]. StatiX [FHR^{+}02]. status [MMJ^{+}02]. Stereotypes [SMM02]. Storage [BFH^{+}02, DT03, HLM03, Hei03, Koc03, LWY^{+}02, MLLA03, CDHZ03, Fei05, YASU01]. Storing [AJEM02, BP05, KP02b, RP02, TVB^{+}02]. strategies [AD04, SK02]. stream [GK05, GGM^{+}04]. streaming [SV02]. Streams [Suc02, GK03, GGM^{+}04, JFB05]. Structural [CVZ^{+}02, DHA^{+}02, For08, KYU02b, PMC02, ZZY^{+}02, TW05]. Structurally [PMC02]. Structure [Cha02b, CP02, ETL02, HKYU02b, KYU02a, PG02b, LYT^{+}05]. Structure- [Cha02b]. Structured [FMPL^{+}03, Kuz02, LL02a, WD02, AKYJ03, CKK03, PG02a]. Structures [Ano02b, SSC^{+}00, UIN02, WPFY03]. Structuring [WS02, Beh00]. Strukturierung [Beh00]. Studio [LR02]. Study [NZ02, CL04, Roc01]. Style [Mey00, Mey01a, St.00]. Subset [XWP02]. Succeeding [CZ01]. Successful [Kun02]. Suite [Ano02c]. Support [Car02a, MEC02, CKN03, HRL^{+}05, MdlFD03]. Supporting [Cha02b, GMRU02, JSSM04, WL02]. Supports [DJM02]. Surfing [Coc01]. SVG [BLS02, Eis02]. Symbol [DHA^{+}02]. Symbolic [Gut04]. Symposium [ACM02a, ACM03b, BCH^{+}05, Gut04, SB00]. Synopses [PG02b, PG02a]. Syntactic [LAG02]. synthetic [BM06]. System [Ano02b, BHK^{+}03, CLCC02, Cha02b, DDPS02, DM03, DT03, GMW00, Hab02, HLM03, HKYU02a, HS02b, Kuz02, LZZ03, LMY02, MEC02, MLLA03, QN02, TMYU02, Tur02a, VAS02, XYW02, dTU04, CRW02, CAYLS03, FHK^{+}02, FP00, Liu00, PAKC^{+}03, SVMAM04, TVB^{+}02, WK06, Bar00, Lut02].

Systems [ACM02a, ACM03b, Ano02b, AvM02, DJM02, Lut02, WL02, Wil03c, LFG^{+}01, Ano02c, Ano03, Int00].

Tables [ETL02, Pan02, WH02]. tagging [CKN03]. Tariff [Bar00]. TAX [JLST02]. TCOZ [DLS^{+}03]. TCP [Lut02]. TCP/IP [Lut02]. teach [LCT01]. teaching [Beh00]. techexplorer [Doo02a]. Technical [ACMS06, FKS^{+}03]. Technique [PMC02]. Techniques [HCC^{+}02, Die01, JKA02, Qui00, SK02]. technological [Wil02b]. Technologies [Lut02, Psa02a, Sye02, TMYU02, BCH^{+}05, Myl02, Tur02b]. technology [FKS^{+}02c, FKS^{+}03, Int00, LB03, RHC^{+}06, VR06]. Tektronix [Ano02b]. Telelogic [Ano02b]. Telephony [Ano02b]. template [BMKL02, Le00]. template-based [BMKL02]. Templates [Geo02, Lat02, Mor00]. Temporal [GS03, Nør02, HYC04]. Terminology [HBH^{+}03]. Test [Bar01, MKR^{+}01, Ano03]. Tester [Ano02c]. Testing [Kit02, Int00]. TEX/LaTEX [RW02a]. Text [Cas06, EMS00, Joh02, Suc02, WS02, AKYJ03, Fie00]. texts [HG01]. Textual [Mas02]. Them [AA04]. theoretic [AL03, AL05]. Third [BCH^{+}05, Pay02, FLMS05]. Thought [Coc01]. Thought-Controlled [Coc01]. Three [Mur00]. TIMBER [JAKC^{+}02, PAKC^{+}03]. Time [LAG02, SSC^{+}00, Wil01b]. tips [JKA02]. Today [Sid02, CZ01]. Together [AA04]. Tomorrow [Sid02]. Tool [Ano02b, Ano02c, GLS^{+}02, HCC^{+}02, Kro00b, Mam01, RW02a, ACN01, Ano03, Fal00a, Fal00b]. toolkit [Qui00]. Tools [Ano02b, Ano02c, ACM01, EF02]. top [MN05]. top-down [MN05]. Topic [DL04]. ToXgene [BMKL02]. Tracing [TA04]. Tracking [Kro00b]. Transducer [LMP02]. Transducer-Based [LMP02]. Transformation [BS02, FCD03, FCD04].

Transformations [HRW02, GR02, LBN03, MN05, Tid07, ZWG^{+}03]. Transformed [CIK02]. Transforming [LC00, HS02a]. Translating [CIK02, RW02a]. Translation [OJCH02, SW02b]. Transmeta [Lea00]. TrAX [Har03]. Tree [CFF^{+}02, JLST02, Koc03, GK03, GK05, WPFY03]. tree-edit [GK03, GK05]. Trees [LAG02, Ng02, CKM02, KM06]. Trends [HBH^{+}03, VN03]. TREX [ZWG^{+}03]. Trifles [Wil03b]. Triumph [Kim01]. Trondheim [BCH^{+}05]. trust [Dav01, GMRRW01]. Tuning [LWP^{+}02]. Twenty [ACM02a, ACM03b, B^{+}02]. Twenty-Eighth [B^{+}02]. Twenty-First [ACM02a]. Twenty-Second [ACM03b]. Twig [JWLY03, BKS02]. twigs [RM06]. Two [Min02b]. Type [DLS^{+}03, LC00, Zhe03]. Typechecking [AMN^{+}03, MN05]. typed [HP03]. Types [CGMS04, HVP00]. Typing [Seg03, GKPS05].

UDDI [Cer02]. ultimate [HS02a, LHO08]. UML [GDB02, Lea00, SMM02]. Unauthorized [Ano02b]. Understanding [AA04, Edg01]. Unification [BS02]. unified [FKS02a]. United [SM01]. Universal [CLCC02]. University [Gut04]. Unix [Ano00]. Unknown [ETL02]. Unveils [Ano02c, Lea00]. Update [KLL02, TEM^{+}01]. Updating [May02, TIHW01]. Upgrades [Ano02b]. Urkunden [Fie00]. Urkundentext [Fie00]. USA [ACM01, Ano02a, FMA02]. Usage [Hun02, SC02]. Use [Hun02, Sye02, HBZ06, SA03]. User [GMRU02, Jan01, KiY02]. User-Defined [KiY02]. Using [ABFS02, BFS^{+}02, Car02a, DHA^{+}02, Geo02, HKYU02b, Hol00b, KL02, KiY02, LN02, Lin03, LZZ03, MLMT02, NR02, PE02, PCK02, TMYU02, Tur02a, YKDC02, BP05, BCP02, CKK03, Die01, EF02, FA00, GK03, GK05, GJK^{+}06, HLM03, Hol01b, Jac03, Knu01, Lad01, LYT^{+}05, LKB^{+}02, See02, SCG01, Tan02, TVB^{+}02, TW02, VR06, YASU01, ZDW^{+}03]. Utah [ACM01]. Utility [DL08, Fal00a, Fal00b].

Valid [CLL02]. Validating [LYT^{+}05, SV02]. Validation [Ano02c, KLL02, BPV04, BMS01]. Value [HBH^{+}03, PG02b]. Variable [Nay02b]. Variable-Size [Nay02b]. Various [dTU04]. VB.NET [SH02]. verifying [AFL02b]. Version [SSC^{+}00, SW02a]. versus [GWT^{+}01]. vertical [HBZ06]. Very [B^{+}02, FLA^{+}03]. view [ACN01, GSH03]. Views [ACM^{+}02b, Bar01, CLL02, Coc01, AMN^{+}03, LBN03, ZDW^{+}03]. Virtual [Bor02, SB00, Die01]. ViST [WPFY03]. Visual [BBSW03, BD00, Gol09, Hud08a, Hud08b, Mun00, SH02]. Visualization [DLS^{+}03, WJ02, Law04]. Visualizing [Law04]. Visually [PE02, RS02]. visXcerpt [BBSW03]. VLDB [FLA^{+}03]. VLDP [B^{+}02]. Voice [Abb02, Bea02, Hou01, Phi01, RS00, SK02]. Voice-Driven [Hou01]. voice-enabled [RS00]. Voice/Web [Phi01]. VoiceXML [Abb02, Bea02, Edg01, FMP02, Hou01, Lar03, Luc00, Phi01, SK02]. Volume [DF03]. VRML [Abi01, Die01, Rah01, SB00].

W3C [BFRW02]. ware [FMS01]. Warehouse [AvM02, HS02b]. Warps [Wil01b]. watermarking [GA03]. Way [LL02b, EP05, TD02]. Ways [Min02b]. WDM [Lut02]. WDM/Optics [Lut02]. Weaving [AF02]. Web [GMRRW01, Law04, SS02, See02, App00, Aga02, Abb02, Abi01, ABS00, ABM^{+}02, Ano02b, Ano02c, Aye00, BFS^{+}02, BLS02, Bur02b, Car00, CK01, Car02a, Cas00, Cer02, Coy02, Dan00, Doo02b, Edg01, Flo00, FP00, GMRU02, GMRRW01, Gra00b, Gun01b, Gur02, HS02a, HCC^{+}02, Hud08a, Hud08b, Hun02, JKA02, JSSM04, Joh02, KKK02, Kro00a, Kun02, LR02, Lea00, Leh02, LCT01, Luc00, LMMT02, Man02b, McF00, MRR01, MRR03, MRRWW04, Myl02, Nae03, NACP01, Nie01, Nie02b, PWK02, Phi01, Rah01, Roc01, RS00, Sid02, STK02, SMM02, SJD01, Sye02, Tam00, Tea01, Thu02, Tur02a, Uni01, WL04, WJ02, Yua01]. Web-Based [HCC^{+}02, Kun02, Luc00]. Web-Oriented [LMMT02]. Web3D [SB00]. WebDAV [QN02]. WebEQ [Kun02, Min02b]. Webmasters [App00]. Website [AF02]. webspace [Abi01, Rah01]. WebSphere [LR02]. weekend [Cal00]. Westbridge [Ano02b]. Where [Whi01]. while [LC00]. Who [Clu03]. WI [FMA02]. Wide [Cas00, MRRWW04, Tea01, Uni01, BGBJ05, GMRRW01, MRR01]. Wildcards [AY08]. Will [Coc01]. Windows [Kro00a]. Winners [Coc01]. Wireless [Ano02c, Coc01, LCZ04]. Wisconsin [ACM02a]. without [EP05]. wizard [Leh02]. WMI [Mar01]. wonglediff [EP05]. work [Dan00, Gun01a]. Workflow [Kro00b, SGW01]. workflows [HRL^{+}05]. Working [Dix01, Fun00]. workload [SY04]. workshop [ACM01, FLMS05]. Workspaces [ABB^{+}03]. World [GWT^{+}01, Hun02, Jon03b, RHC^{+}06, Cas00, GMRRW01, MRR01, MRRWW04, Tea01, Uni01]. World-Wide [GMRRW01, MRR01]. worlds [Die01]. Worldwide [MRR03]. Worries [SSC^{+}00]. Worterbucher [LSS01]. Wrapper [YKDC02]. Writing [Mam01, Bur02b]. WSDL [Cer02].

X [BCF01, BGBJ05, GMRU02, dTU04]. X-Compass [GMRU02]. X-Global [dTU04]. X-GTRBAC [BGBJ05]. X2RMap [KP02b]. XCache [CRW02]. Xcerpt [BBSW03]. XDuce [HP03]. XGuide [KKK02]. XHTML [Aga02, BGR^{+}00, Car02a, Edd00, Gra00a, Gra00b, Hol01b, Hud08a, Hud08b, Lad01, LCT01, MK00, MK02, SW01, SD00, Tit02].

Xilinx [Ano02b]. XISS [HLM03]. XISS/R [HLM03]. XMark [SWK^{+}02]. XMAS [Le00]. XMI [GDB02]. XML [Ano00, Beh00, BCH^{+}05, BFRW02, Bur02a, End00, EMS00, Fie00, FLMS05, Hay02, Hei01, Law04, LSS01, MB03, Pay02, Sem02, TEM^{+}01, Wil02a, ABS00, ABM^{+}02, ABC^{+}03a, ABB^{+}03, ASV06, ACMS06, ACM^{+}02b, Ahm01, AA04, AJEM02, AKYJ03, AMN^{+}03, ABFS02, AYFSX03a, AYFSX03b, ABB^{+}00, Ano02b, Ano02c, Ano03, A.01, Arc02, AD04, AL02, AFL02b, AFL02a, AL03, AL05, ABC^{+}03b, ASY02, AvM02, BBB^{+}00, BPW02, BHK^{+}03, BPV04, BP05, BEH^{+}06, Bar01, BMKL02, BM06, Bea03, Beh00, BFS^{+}02, BCF03, BBSW03, BB02, BCF01, BGMT02, BF02, BCH^{+}06, BGBJ05, Bin03, Bir01, BCD^{+}02, BHW^{+}02, BFH^{+}02, BDG^{+}03, BCP01, BCP02, BL05, Bor02, BSL00, Bra00, BCKL02, BD00, BP01, Bro03, BKS02, BS02, BDF^{+}02a, BDF^{+}02b, BGK03, CH00, Cag00, Car02b, Cas06, Cer02, CLCC02, Cha02a, Cha03, CFGR02]. XML [CFF^{+}02, Cha00, CSK01, Cha02b, CH06, CIK02, CZ01, CKN03, CLL02, CRW02, CDHZ03, CL04, CVZ^{+}02, CTZ02, CAYLS03, CMS02, CK02, CKK03, CP02, Cle01a, Cle01b, Clu03, Coc01, CKM02, CKS02b, CKS02a, CMKS03, CGMS04, CDF01, Cox01a, Cox01b, Coy02, DS00, DDPS02, Dau03, DM03, Dav03, Dav01, Dei01, DHA^{+}02, Des00, DT03, DT05, DAF^{+}03, DF03, DLS^{+}03, DL04, Dou02, Dwe00a, Dwe00b, EF02, EC01, ED00, ED01, EM02, EWH^{+}02, End00, EMS00, EST^{+}07, Fal00a, Fal00b, FKS02a, FL02, Fei05, FCD02, FCD03, FJ04, FCD04, FMS01, FKS^{+}02b, FMPL^{+}03, FHK^{+}02, FA00, Flo00, For08, Fox02, FHR^{+}02, FB04, FKS^{+}02c, FMR02, FKS^{+}03, Fun00, GLFO^{+}03, GS03, GR02, GF02, GK03, GK05, GMRU02, Geo02, GMRRW01, GWT^{+}01, Gol09, GP00, GP01a, GP01b, GP02, GP04, GMW00, GLS^{+}02]. XML [Gös03, GKPS05, GGM^{+}04, Gri02, GDB02, GA03, GJK^{+}02, GJK^{+}06, GSBS03, GSH03, Gur02, HBG^{+}03, HG01, HLM03, Har01, HM01, HM02, Har03, HM04, HRL^{+}05, HKYU02b, HKYU02a, Hei03, HBH^{+}03, Her02, HBZ06, Hol01a, HSJJ02a, HSJJ02b, Hoq00, HVP00, HP01, HP03, HFC05, HCC^{+}02, HS02b, HYC04, Hud08b, HC^{+}00, H^{+}07, IHW02, Jac03, JAKC^{+}02, JLST02, Jak04, JWLY03, JD02, Jon03a, Jon03b, JFB05, JOKA02, KM06, KL02, KKK02, KYU02a, KYU02b, Kim01, KSK^{+}02, KP02b, KLL02, KiY02, KP02a, KSH02, Koc03, KP04, KP05, KPSS02, Kro00a, Kro00b, KC02, Lad01, LN02, LFG^{+}01, LR02, Law02, Lea00, LC00, LCC^{+}02, LL02a, LB03, LSS01, Lew02, LBN03, LCZ04, LWP^{+}02, Lin03, Liu00, LL02b, LZZ03, LYT^{+}05, LMY02, LW04, LC04, LMP02, LAG02, LMMT02, Lut02, LKB^{+}02, LWY^{+}02, Mam01, Mar00, MS03, Mar04, Mar05]. XML [MN05, MKR^{+}01, MdlFD03, Mar01, Mas02, May02, McG00, McL00, McL01, McL02, McL07, MLMT02, MB06, MLLA03, Mer01, MAA^{+}03, MAA^{+}05, MAC03, MPC03, MlF01, Mue00, Mun00, Mur00, MRR01, MRR03, MRRWW04, MB03, Mus01, Myl02, Nac02, NQ02, Nae03, NLB^{+}02, NWB00, NZ02, Nev02, Ng02, NACP01, NJ02, Nør02, Ass02, Oba03, OJCH02, PE02, Pan02, PAKJ^{+}02, PAKC^{+}03, PMC02, PCK02, PLM^{+}02, Paw02, PRP02, PWK02, PAB02, PMK^{+}06, PPV02, PSK02, PG02a, PG02b, Por03, Psa02a, Psa02b, QN02, Qui00, RRB03, Ram03, RM06, Ray01, RM02, Roc01, RHC^{+}06, RP02, Sah01, SSC^{+}00, Sav01, SWK^{+}02, SB02, SS02, Sea00, See02, SV02, Seg03, Sel02, Sem02, SY04, SRCV06, SSB^{+}01, SGW01, SG02a, SW02a, SW03, Sim01a, Sim01b, Sim02b, Sin00, SVMAM04, SG02b, SCG01]. XML [SMM02, St.00, SD00, SJD01, SFE05, Sta00, Sta01, SH02, Str02, Stu00, Suc02, SC02, SM02, SDC04, SA03, Tam00, TMK05, Tan02, TMYU02, TIHW01, TVB^{+}02, TW02, TL04, TB04, Thu02, Tid07, TW05, Tra00, Tur02a, Tur02b, Uni01, VAS02, VN03, VFMMP06, VR06, Wah02, Wal02, WL02, WPFY03, WL04, WD02, WK03, WK06, Wig00, Wil02b, Wil01a, Wil01c, Wil03c, WS02, XW02, XWW^{+}02, XYW02, YKDC02, YLH03, YHL02, YASU01, YSLJ02, YWL^{+}03, YLM^{+}05, ZDW^{+}03, ZZY^{+}02, Zhe03, ZWG^{+}03, dTU04, vdV02, Rol00]. XML-based [Law04, AvM02, CLCC02, DLS^{+}03, FMPL^{+}03, Gös03, HS02b, KL02, KKK02, LMY02, PAB02, PSK02, SC02, Wil03c, XYW02, YKDC02, AD04, BGBJ05, KP04, Kro00a, Mam01, PWK02, SDC04, WD02]. XML-basiertes [Beh00]. XML-centric [BCF03, RHC^{+}06]. XML-Driven [Mas02]. XML-enabled [SGW01]. XML-Encoded [LL02a]. XML-Oriented [Ano02c]. XML-RPC [Cer02, Jon03a, Por03, SJD01]. XML-Schema [For08, HSJJ02a, HSJJ02b]. XML-SQL [Pan02]. XML/EDI [LW04]. XML/HTML [Jac03]. XML/JSP/WebDAV [QN02]. XML/SNMP [OJCH02]. XOCL [RRB03]. XPath [Cas06, CFGR02, GR02, GKPS05, Kay08, LYT^{+}05, MB06, Sim02b, SG02b]. XPathLearner [LWP^{+}02]. XPointer [Sim02b]. XPRESS [MPC03]. XQForms [PPV02]. Xquec [ABC^{+}03b]. XQuery [BCH^{+}06, Cas06, Cha02a, Cha03, FMR02, MB06, ZDW^{+}03]. XRANK [GSBS03]. XRel [YASU01]. XRL [PAB02]. XSEarch [CMKS03]. XSL [Fit01, Geo02, LBN03, LL02b, Paw02, Sim02a, Sta03, Sta06, Sta07]. XSL-FO [Paw02]. XSLT [Bur01, Car02a, CH06, Fit03, Fun00, GR02, HS02a, Hol02, Hud08b, Kay00, Kay01, Kay08, Len05, LMMT02, Man02a, Nac02, Roc01, SG02b, Tid01, Tid07, Tur02a]. XSym [BCH^{+}05]. XTABLES [FKS^{+}02c, FKS^{+}03]. XTM [DL04]. XXL [TW02].

Y2K [SSC^{+}00]. yourself [LCT01].

Zope [Lat02]. zur [Beh00].

References

ArciniegasA:2001:XDG

[A.01] Fabio Arciniegas A. XML developer’s guide. McGraw-Hill, New York, NY, USA, 2001. ISBN 0-07-212648-5. xviii + 682 pp. LCCN QA76.76.H94 A75 2001.

Aiken:2004:XDM

[AA04] Peter Aiken and M. David Allen. XML in Data Management: Understanding and Applying Them Together. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2004. ISBN 0-12-045599-4. 398 (est.) pp. LCCN ???? US$44.95.

Anderson:2000:PX

[ABB^{+}00] Richard Anderson, David Baliles, Mark Birbeck, Michael Kay, Steven Livingstone, Brian Loesgen, Didier Martin, Stephen Mohr, Nikola Ozu, Bruce Peat, Jonathan Pinnock, Peter Stark, and Kevin Williams. Professional XML. Wrox Press, Chicago, IL, USA, 2000. ISBN 1-86100-311-0. xxi + 1169 pp. LCCN QA76.76.H94 P764 2000. US$49.99.

Abbott:2002:VEW

[Abb02] Kenneth R. Abbott. Voice Enabling Web Applications: VoiceXML and Beyond. Apress, Berkeley, CA, USA, 2002.

ISBN 1-893115-73-9. xii + 250 pp. LCCN QA76.76.H94 A23 2002. US$34.95. URL http://

www.apress.com/.

Abiteboul:2003:MDW

[ABB^{+}03] Serge Abiteboul, Jérôme Baumgarten, Angela Bonifati, Gregory Cobena, Cosmin Cremarenco, Florin Dragan, Ioana Manolescu, Tova Milo, and Nicoleta Preda. Managing distributed workspaces with Active XML. In Freytag et al. [FLA^{+}03], pages 1061– 1064. ISBN 0-12-722442-4.

URL http://www.vldb.org/
dblp/db/indices/a-tree/a/
Abiteboul:Serge.html.

Abiteboul:2003:DXD

[ABC^{+}03a] Serge Abiteboul, Angela Bonifati, Grégory Cobéna, Ioana Manolescu, and Tova Milo. Dynamic XML documents with distribution and replication. In ACM [ACM03a], pages 527– 538. ISBN ???? LCCN ????

Arion:2003:XPQ

[ABC^{+}03b] Andrei Arion, Angela Bonifati, Gianni Costa, Sandra D’Aguanno, Ioana Manolescu, and Andrea Pugliese. Xquec: Pushing queries to compressed XML data. In Freytag et al. [FLA^{+}03], pages 1065– 1068. ISBN 0-12-722442-4.

URL http://www.vldb.org/
dblp/db/indices/a-tree/a/
Arion:Andrei.html.

Asperti:2003:MKM

[ABD03] Andrea Asperti, Bruno Buchberger, and James Harold Davenport, editors. Mathematical Knowledge Management: Second International Conference, MKM 2003 Bertinoro, Italy, February 16–18, 2003, volume 2594 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. ISBN 3-540-00568-4. 65.50 SFR; 40.66 EUR.

Amann:2002:QXS

[ABFS02] Bernd Amann, Catriel Beeri, Irini Fundulaki, and Michel Scholl. Querying XML sources using an ontology-based mediator. Lecture Notes in Computer Science, 2519:429– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2519/25190429.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2519/25190429.pdf.

AbidurRahman:2001:AVW

[Abi01] Sharif Md Abidur Rahman.

Application of VRML in Web design: a 3D webspace for Hotel Sonargaon. Project (M. Arch.), UCLA School of Architecture, Los Angeles, CA, USA, 2001. Comprehensive examination for UCLA School of Architecture.

Abiteboul:2002:AXP

[ABM^{+}02] Serge Abiteboul, Omar Benjelloun, Ioana Manolescu, Tova Milo, and Roger Weber. Active XML: Peer-to-peer data and Web services integration. In Bernstein et al. [B^{+}02], pages 1087–1090. ISBN 1-55860-869-9. LCCN ???? URL http:

                                        //www.vldb.org/conf/2002/
                                        S33P12.pdf.

Abiteboul:2000:DWR

[ABS00] S. (Serge) Abiteboul, Peter Buneman, and Dan Suciu. Data on the Web: from relations to semistructured data and XML. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2000. ISBN 1-55860-622-X. xiii + 257 pp. LCCN QA76.9.D3 A258 2000.

ACM:2001:ASS

[ACM01] ACM, editor. ACM SIGPLAN– SIGSOFT workshop on Program analysis for software tools and engineering: June 18–19, 2001, Snowbird, Utah, USA: PASTE’01. ACM Press, New York, NY, USA, 2001. ISBN 1-58113-413-4. LCCN QA76.758 .A345 2001. Supplement to ACM SIGPLAN Notices.

ACM:2002:PTF

[ACM02a] ACM, editor. Proceedings of the Twenty-First ACM SIGMOD– SIGACT–SIGART Symposium on Principles of Database Systems: PODS 2002: Madison, Wisconsin, June 3–5, 2002. ACM Press, New York, NY, USA, 2002. ISBN 1-58113-507-

6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Aguilera:2002:VLS

[ACM^{+}02b] Vincent Aguilera, Sophie Cluet, Tova Milo, Pierangelo Veltri, and Dan Vodislav. Views in a large-scale XML repository. VLDB Journal: Very Large Data Bases, 11(3):238– 255, November 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL http://link.springer.
de/link/service/journals/
00778/bibs/2011003/20110238.
htm;   http://link.springer.
de/link/service/journals/
00778/papers/2011003/20110238.

pdf. Special issue VLDB best papers 2001.

ACM:2003:PAS

[ACM03a] ACM, editor. Proceedings of the 2003 ACM SIGMOD International Conference on Management of Data 2003, San Diego, California, June 09–12, 2003. ACM Press, New York, NY, USA, 2003. ISBN ???? LCCN ????

ACM:2003:PTS

[ACM03b] ACM, editor. Proceedings of the Twenty-Second ACM SIGMOD-SIGACT-SIGART Symposium on Principles of Database Systems: PODS 2003: San Diego, Calif., June 9–11, 2003. ACM Press, New York, NY, USA, 2003. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Adler:2006:TCC

[ACMS06] S. Adler, R. Cochrane, J. F. Morar, and A. Spector. Technical context and cultural consequences of XML. IBM Systems Journal, 45(2):207–223, ???? 2006. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/452/adler.html.

Agrawal:2001:MVI

[ACN01] Sanjay Agrawal, Surajit Chaudhuri, and Vivek Narasayya. Materialized view and index selection tool for Microsoft SQL Server 2000. SIGMOD Record (ACM Special Interest Group on Management of Data), 30 (2):608, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ardagna:2004:CMS

[AD04] C. Ardagna and S. De capitani di Vimercati. A comparison of modeling strategies in defining XML-based access control languages. International Journal of Computer Systems Science and Engineering, 19(3):??, May 2004. CODEN CSSEEI. ISSN 0267-6192.

Adams:2002:PJ

[Ada02] D. J. Adams. Programming Jabber. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00202-5. xxi + 455 pp. LCCN 49876 Stacks SEMSTAX-NBKS. US$39.95.

URL http://safari.oreilly.
com/0596002025;   http://
www.oreilly.com/catalog/
jabber.

Anderson-Freed:2002:WWP

[AF02] Susan Anderson-Freed. Weaving a Website: programming in HTML, JavaScript, Perl and Java. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-028220-0. 712 pp. LCCN QA76.625 .A64 2002. BUY-A-BOOK gift of Michael E. Hayden thanks to Professor Frank Deremer.

Arenas:2002:WHA

[AFL02a] M. Arenas, W. Fan, and L. Libkin. What’s hard about XML schema constraints? Lecture Notes in Computer Science, 2453:269– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2453/24530269.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2453/24530269.
pdf.

Arenas:2002:VCX

[AFL02b] Marcelo Arenas, Wenfei Fan, and Leonid Libkin. On verifying consistency of XML specifications. In ACM [ACM02a], pages 259–270. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

AC:2002:HXC

[Aga02] Against the Clock. HTML and XHTML: creating Web pages. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-031054-9. xv + 356 pp. LCCN QA76.76.H94 H75 2002. Includes CD-ROM.

Ahmed:2001:PJX

[Ahm01] Kal Ahmed. Professional Java XML. Programmer to programmer. Wrox Press, Chicago, IL, USA, 2001. ISBN 1-86100-401-X. xv + 1159 pp. LCCN QA76.76.H94 P7626 2001.

Al-Jadir:2002:FXS

[AJEM02] Lina Al-Jadir and Fatmé El-Moukaddem. F2/XML: Storing XML documents in object databases. Lecture Notes in Computer Science, 2425:108– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2425/24250108.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2425/24250108.
                                        pdf.

Asperti:2002:MMP

[AK02] Andrea Asperti and Michael Kohlhase. MathML in the MOWGLI project. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ???? URL http:/

/www.mathmlconference.com/
2002/presentations/asperti/

.

Al-Khalifa:2003:QST

[AKYJ03] Shurug Al-Khalifa, Cong Yu, and H. V. Jagadish. Querying structured text in an XML database. In ACM [ACM03a], pages 4–15. ISBN ???? LCCN ????

Arenas:2002:NFX

[AL02] Marcelo Arenas and Leonid Libkin. A normal form for XML documents. In ACM [ACM02a], pages 85–96. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Arenas:2003:ITA

[AL03] Marcelo Arenas and Leonid Libkin. An informationtheoretic approach to normal forms for relational and XML data. In ACM [ACM03b], pages 15–26. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Arenas:2005:ITA

[AL05] Marcelo Arenas and Leonid Libkin. An informationtheoretic approach to normal forms for relational and XML data. Journal of the ACM, 52 (2):246–283, March 2005. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Alon:2003:TXV

[AMN^{+}03] Noga Alon, Tova Milo, Frank Neven, Dan Suciu, and Victor Vianu. Typechecking XML views of relational databases. ACM Transactions on Computational Logic, 4(3):315–354, July 2003. CODEN ???? ISSN 1529-3785 (print), 1557-945X (electronic).

Angell:2000:PSPb

[Ang00] Kirby W. Angell. Python server pages: Part II. Dr. Dobbs Journal, 25(2):54, 57–61, February 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2000/2000_02/
psp2.txt.

Anonymous:2000:BRU

[Ano00] Anonymous. Book reviews: Unix Backup and Recovery by Charles Curley; LaTEX for Linux by Ben Crowder; The XML Handbook 2nd Edition, by Daniel Lazenby; Securing Linux, by Charles Curley; Building Linux and OpenBSD Firewalls, by Ralph Krause; Linux Programming Bible by Ben Crowder. Linux Journal, 78:??, October 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.
com/lj-issues/issue78/3839.
html;   http://noframes.
linuxjournal.com/lj-issues/
issue78/3872.html.

Anonymous:2002:MIC

[Ano02a] Anonymous, editor. MathML International Conference: Hickory Ridge Conference Center, Chicago, IL, USA, June 28–30, 2000. ????, ????, 2002. ISBN ???? LCCN ????

Anonymous:2002:PSS

[Ano02b] Anonymous. Products: SOISIC ships deisgn kit for SOI structures; systems and software development tools from Telelogic; RSA Security’s Web access management system; Altera’s free embedded processor portfolio; signal integrity measurement tools from tektronix; Oracle upgrades Java development tool; Xilinx delivers EDK for FPGA processor; Westbridge’s tool to sniff unauthorized XML; SpeechStudio’s telephony development tools. Computer, 35(12):118–119, December 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2002/12/rz118.
htm;   http://csdl.computer.
org/dl/mags/co/2002/12/rz118.
pdf.

Anonymous:2002:PXO

[Ano02c] Anonymous. Products: XMLoriented FrameMaker from Adobe Systems; Breeze Factor’s XML data-binding tool for Java; Microsoft’s speechenabling tools for Web applications; Agere Systems’s wireless LAN management tool; Macromedia debuts ColdFusion MX; OpenOffice.org unveils open source productivity suite; Integrated Measurement Systems’ validation tester. Computer, 35(6):82– 83, June 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2002/06/r6082.
htm;   http://csdl.computer.
org/dl/mags/co/2002/06/r6082.
pdf.

Anonymous:2003:PBS

[Ano03] Anonymous. Products: BEA Systems releases preview of Java API for XML; Panda appliance defends network perimeters; Parasoft enhances Java test and analysis tool. Computer, 36(12):108, December 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2003/12/rz108.
pdf.

Apple:2000:QWH

[App00] Apple Computer, Inc. Quick-Time for the WEB: a handson guide for Webmasters, site designers, and HTML authors. QuickTime developer series. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2000. ISBN 0-12-471255-X, 0-12-471256-8 (CDROM). xxvi + 720 + 15 pp. LCCN TK5105.888 .Q5 2000.

ArciniegasA:2002:CX

[Arc02] Fabio Arciniegas A. C++ XML. New Riders Publishing, Carmel, IN, USA, 2002. ISBN 0-7357-1052-X. xvi + 316 pp. LCCN QA76.76.H94 A73 2002.

OReilly:2002:XCB

[Ass02] O’Reilly and Inc. Associates.

XML CD Bookshelf. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00335-8. 600 (est.) pp. LCCN ???? US$119.95.

URL http://www.oreilly.
com/catalog/xmlcdbksf.

Abiteboul:2006:RQX

[ASV06] Serge Abiteboul, Luc Segoufin, and Victor Vianu. Representing and querying XML with incomplete information. ACM Transactions on Database Systems, 31(1):208–254, March 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Asakura:2002:XK

[ASY02] Hiroshi Asakura, Takeyuki Shimura, and Yohei Yamamoto. XML kyokasho. Sofuto Risachi Senta, Tokyo, 2002. ISBN 4-88373-167-7. 457 pp. LCCN ????

Auth:2002:SAX

[AvM02] Gunnar Auth and Eitel von Maur. A software architecture for XML-based metadata interchange in data warehouse systems. Lecture Notes in Computer Science, 2490:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2490/24900001.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2490/24900001.pdf.

Altamimi:2008:MQL

[AY08] Moody Ebrahem Altamimi and Abdou Youssef. A math query language with an expanded set of wildcards. Mathematics in Computer Science, 2 (2):305–331, December 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=1661-
8270&volume=2&issue=2&spage=
305.

Ayesh:2000:EDH

[Aye00]  Aladdin Ayesh.  Essential

Dynamic HTML Fast: developing an interactive Web site. Springer-Verlag Inc., New York, NY, USA, 2000. ISBN 1-85233-626-9. ix + 153 pp. LCCN QA76.76.H94 A99 2000. US$28.00.

Amer-Yahia:2003:PEA

[AYFSX03a] Sihem Amer-Yahia, Mary Fernández, Divesh Srivastava, and Yu Xu. PIX: exact and approximate phrase matching in XML. In ACM [ACM03a], page 664. ISBN ???? LCCN ????

Amer-Yahia:2003:PMX

[AYFSX03b] Sihem Amer-Yahia, Mary F.

Fernandez, Divesh Srivastava, and Yu Xu. Phrase matching in XML. In Freytag et al. [FLA^{+}03], pages 177– 188. ISBN 0-12-722442-4. URL

http://www.vldb.org/dblp/
db/indices/a-tree/a/Amer=
Yahia:Sihem.html.

Bernstein:2002:VPT

[B^{+}02] Philip A. Bernstein et al., editors. VLDP 2002: proceedings of the Twenty-Eighth International Conference on Very Large Data Bases, Hong Kong SAR, China, 20–23 August 2002. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2002. ISBN 1-55860-869-9. LCCN ????

Barker:2000:BBA

[Bar00] John S. Barker. Building the Bell Atlantic Tariff Management System. Dr. Dobbs Journal, 25(5):90, 92–95, 97, May 2000. CODEN DDJOEB. ISSN 1044-789X.

Baran:2001:NVCb

[Bar01] Nicholas Baran. News and views: College Board to add Java test; astronomy project beats out algorithm for science prize; consortium to release XML business specification; open source developer agreement proposed; consortium hopes to eliminate computer failures; new linear algorithm for sequence analysis; distributing computing à la SETI catches on. Dr. Dobbs Journal, 26(3):18, March 2001. CODEN DDJOEB. ISSN 1044-789X. URL

                                        http://publicscience.net/
                                        ;   http://www.ddj.com/.

Bavestrelli:2000:BHD

[Bav00] Giovanni Bavestrelli. Building HTML documents with C++. C/C++ Users Journal, 18(4): 67–??, April 2000. CODEN CCUJEX. ISSN 1075-2838.

Berstel:2002:FPX

[BB02] Jean Berstel and Luc Boasson. Formal properties of XML grammars and languages. Acta Informatica, 38(9):649– 671, August 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Baartse:2000:PAX

[BBB^{+}00] Mark Baartse, Richard Blair, Luca Bolognese, Dinar Dalvi, Steven Hahn, Corey Haines, Alex Homer, Bill Kropog, Brian Loesgen, Stephen Mohr, John Slater, Kevin Williams, and Mario Zucca. Professional ASP XML. Wrox Press, Chicago, IL, USA, 2000. ISBN 1-86100-402-8. xix + 883 pp. LCCN TK5105.8885.A26 P789 2000. US$49.99.

Berger:2003:XVP

[BBSW03] Sacha Berger, François Bry, Sebastian Schaffert, and Christoph Wieser. Xcerpt and visXcerpt: From pattern-based to visual querying of XML and semistructured data. In Freytag et al. [FLA^{+}03], pages 1053–1056. ISBN 0-12-722442-4. URL http://www.

vldb.org/dblp/db/indices/
a-tree/b/Berger:Sacha.html.

Blanc:2002:CSR

[BCD^{+}02] Mathieu Blanc, Fabien Costantini, Sébastien Dubois, Manuel Forget, Olivier Francillon, and Christian Toinard. Collaborative simulation by reuse of COTS simulators with a reflexive XML middleware1. Lecture Notes in Computer Science, 2426:142–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2426/24260142.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2426/24260142.pdf.

Bertino:2001:SXD

[BCF01] Elisa Bertino, Silvana Castano, and Elena Ferrari. Securing XML documents: the author-X project demonstration. In Sellis and Mehrotra [SM01], page 605. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Benzaken:2003:CXC

[BCF03] Véronique Benzaken, Giuseppe Castagna, and Alain Frisch.

CDuce: an XML-centric general-purpose language. ACM SIGPLAN Notices, 38(9):51– 63, September 2003. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bressan:2005:DXT

[BCH^{+}05] Stéphane Bressan, Stefano Ceri, Ela Hunt, Zachary G. Ives, Zohra Bellahsène, Michael Rys, and Rainer Unland, editors. Database and XML technologies: Third International XML Database Symposium, XSym 2005, Trondheim, Norway, August 28– 29, 2005: proceedings, volume 3671 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28583-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3.I5978 2005. URL

http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3671.

Beyer:2006:DGH

[BCH^{+}06] K. Beyer, R. Cochrane, M. Hvizdos, V. Josifovski, J. Kleewein, G. Lapis, G. Lohman, R. Lyle, M. Nicola, F. Özcan, H. Pirahesh, N. Seemann, A. Singh, T. Truong, R. C. Van der Linden, B. Vickery, C. Zhang, and G. Zhang. DB2 goes hybrid: Integrating native XML and XQuery with relational data and SQL. IBM Systems Journal, 45(2):271–298, ????

2006. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/452/beyer.html.

Braga:2002:MAR

[BCKL02] Daniele Braga, Alessandro Campi, Mika Klemettinen, and PierLuca Lanzi. Mining association rules from XML data. Lecture Notes in Computer Science, 2454:21– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2454/24540021.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2454/24540021.
pdf.

Bonifati:2001:ARX

[BCP01] Angela Bonifati, Stefano Ceri, and Stefano Paraboschi. Active rules for XML: A new paradigm for E-services. VLDB Journal: Very Large Data Bases, 10(1): 39–47, August 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL http://link.springer.
de/link/service/journals/
00778/bibs/1010001/10100039.
htm;   http://link.springer.
de/link/service/journals/
00778/papers/1010001/10100039.
pdf.

Bonifati:2002:PRS

[BCP02] Angela Bonifati, Stefano Ceri, and Stefano Paraboschi. Pushing reactive services to XML repositories using active rules. Computer Networks (Amsterdam, Netherlands: 1999), 39 (5):645–660, August 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.com/gej-ng/
                                        10/15/22/96/52/38/abstract.
                                        html.

Britt:2000:PVB

[BD00] James Britt and Teun Duynstee. Professional Visual Basic 6 XML. Wrox Press, Chicago, IL, USA, 2000. ISBN 1-86100-332-3. xi + 725 pp. LCCN QA76.73.B3 B7425 2000.

Buneman:2002:KX

[BDF^{+}02a] Peter Buneman, Susan Davidson, Wenfei Fan, Carmem Hara, and Wang-Chiew Tan. Keys for XML. Computer Networks (Amsterdam, Netherlands: 1999), 39(5):473–487, August 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/15/22/96/52/28/
                                        abstract.html.

Buneman:2002:RAK

[BDF^{+}02b] Peter Buneman, Susan Davidson, Wenfei Fan, Carmem Hara, and Wang-Chiew Tan. Reasoning about keys for XML. Lecture Notes in Computer Science, 2397:133–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2397/23970133.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2397/23970133.pdf.

Bohannon:2003:RRL

[BDG^{+}03] Philip Bohannon, Xin (Luna)

Dong, Sumit Ganguly, Henry F. Korth, Chengkai Li, P. P. S. Narayan, and Pradeep Shenoy. ROLEX: relational on-line exchange with XML. In ACM [ACM03a], page 673. ISBN ???? LCCN ????

Beasley:2002:VAD

[Bea02] Rick Beasley. Voice application development with VoiceXML. SAMS Publishing, Indianapolis, IN, USA, 2002. ISBN 0-672-32138-6. ix + 382 pp. LCCN QA76.76.H94 B45 2002.

Bean:2003:XDA

[Bea03] James Bean. XML for Data Architects: Designing for Reuse and Integration. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 1-55860-907-5. 270 (est.) pp. LCCN ???? US$39.95.

Behrens:2000:XMM

[Beh00]  Ralf Behrens.  Ein XML-

basiertes Modell zur Modellierung und Strukturierung von Informationen im Kontext eines Medienarchivs für die Lehre. (German) [An XML based model for the modelling and structuring of information in the context of a medium file for the teaching]. Thesis (doctoral), Medizinische Universität zu Lübeck, Lübeck, Germany, 2000. 182 pp.

Balmin:2006:CBO

[BEH^{+}06] A. Balmin, T. Eliaz, J. Hornibrook, L. Lim, G. M. Lohman, D. Simmen, M. Wang, and C. Zhang. Cost-based optimization in DB2 XML. IBM Systems Journal, 45(2):299– 319, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/452/balmin.html.

Bertino:2002:SSD

[BF02] Elisa Bertino and Elena Ferrari. Secure and selective dissemination of XML documents. ACM Transactions on Information and System Security, 5 (3):290–331, August 2002. CO-DEN ATISBQ. ISSN 1094-9224 (print), 1557-7406 (electronic).

Bohannon:2002:LCR

[BFH^{+}02] Philip Bohannon, Juliana Freire, Jayant R. Haritsa, Maya Ramanath, Prasan Roy, and Jérôme Siméon. LegoDB: Customizing relational storage for XML documents. In Bernstein et al. [B^{+}02], pages 1091–1094. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.
org/conf/2002/S33P13.pdf.

Brown:2002:MMW

[BFRW02] Allen Brown, Matthew Fuchs, Jonathan Robie, and Philip Wadler. MSL: a model for W3C XML Schema. Computer Networks (Amsterdam, Netherlands: 1999), 39(5):507– 521, August 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/
gej-ng/10/15/22/96/52/30/
abstract.html.

Bell:2002:FWA

[BFS^{+}02] Eric Bell, Hao Howard Feng, Edward L. W. Soong, David Zhang, and Shijia Sam Zhu. Fundamentals of Web Applications Using .NET and XML. The integrated .NET series from Object Innovations and Prentice Hall PTR. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-041790-4. xxvi + 564 pp. LCCN TK5105.888 .F86 2002. US$44.99. URL http:

//www.phptr.com/ptrbooks/
ptr_0130417904.html.

Bhatti:2005:XGX

[BGBJ05] Rafae Bhatti, Arif Ghafoor, Elisa Bertino, and James B. D. Joshi. X-GTRBAC: an XML-based policy specification framework and architecture for enterprise-wide access control. ACM Transactions on Information and System Security, 8 (2):187–227, May 2005. CO-DEN ATISBQ. ISSN 1094-9224 (print), 1557-7406 (electronic).

Buneman:2003:PQC

[BGK03] Peter Buneman, Martin Grohe, and Christoph Koch. Path queries on compressed XML. In Freytag et al. [FLA^{+}03], pages 141–152. ISBN 0-12-722442-4.

                                        URL http://www.vldb.org/
                                        dblp/db/indices/a-tree/b/
                                        Buneman:Peter.html.

Bertino:2002:ESD

[BGMT02] Elisa Bertino, Giovanna Guerrini, Marco Mesiti, and Luigi Tosetto. Evolving a set of DTDs according to a dynamic set of XML documents. Lecture Notes in Computer Science, 2490:45– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/
                                        service/series/0558/bibs/
                                        2490/24900045.htm; http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2490/24900045.pdf.

Boumphrey:2000:BX

[BGR^{+}00] Frank Boumphrey, Cassandra Greer, Dave Raggett, Jenny Raggett, Sebastian Schnitzenbaumer, and Ted Wugofski. Beginning XHTML. Wrox Press, Chicago, IL, USA, 2000. ISBN 1-86100-343-9. xxi + 733 pp. LCCN QA76.76.H94 B438 2000. US$39.99.

Balmin:2003:SKP

[BHK^{+}03] Andrey Balmin, Vagelis Hristidis, Nick Koudas, Yannis Papakonstantinou, Divesh Srivastava, and Tianqiu Wang. A system for keyword proximity search on XML databases. In Freytag et al. [FLA^{+}03], pages 1069– 1072. ISBN 0-12-722442-4.

URL http://www.vldb.org/
dblp/db/indices/a-tree/b/
Balmin:Andrey.html.

Boer:2002:PDL

[BHW^{+}02] Alexander Boer, Rinke Hoekstra, Radboud Winkels, Tom van Engers, and Frederik Willaert. Proposal for a Dutch legal XML Standard. Lecture Notes in Computer Science, 2456:142–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2456/24560142.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2456/24560142.pdf.

Binstock:2003:XSC

[Bin03] Cliff Binstock, editor. The XML schema complete reference. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-672-32374-5. xxxvii + 965 pp. LCCN QA76.76.H94 X4549 2003.

Birbeck:2001:PX

[Bir01] Mark Birbeck. Professional XML. Programmer to programmer. Wrox Press, Chicago, IL, USA, second edition, 2001. ISBN 1-86100-505-9 (paperback). xiii + 1269 pp. LCCN QA76.76.H94.P76 2001.

Bruno:2002:HTJ

[BKS02] Nicolas Bruno, Nick Koudas, and Divesh Srivastava. Holistic twig joins: optimal XML pattern matching. In Franklin et al. [FMA02], pages 310–321. ISBN ???? LCCN ???? ACM order number 475020.

Bordawekar:2005:APX

[BL05] Rajesh R. Bordawekar and Christian A. Lang. Analytical processing of XML documents: opportunities and challenges. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(2):27–32, June 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Broeglin:2002:DMS

[BLS02] Dominique Broeglin, Stéphane Lavirotte, and Peter Sander. Displaying mathematics on the Semantic Web: MathML content to SVG. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Barbosa:2006:DGS

[BM06] Denilson Barbosa and Alberto O. Mendelzon. Declarative generation of synthetic XML data. Software—Practice and Experience, 36(10): 1051–1079, August 2006. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bernardin:2002:MM

[BMH02] Laurent Bernardin, James Mc-Carron, and Douglas Harder. MathML in Maple. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Barbosa:2002:TTB

[BMKL02] Denilson Barbosa, Alberto Mendelzon, John Keenleyside, and Kelly Lyons. Toxgene: a template-based data generator for XML. In Franklin et al. [FMA02], page 616. ISBN ???? LCCN ???? ACM order number 475020.

Braband:2001:SVD

[BMS01] Claus Braband, Anders Møller, and Michael Schwartzbach. Static validation of dynamically generated HTML. In ACM [ACM01], pages 38–45. ISBN 1-58113-413-4. LCCN QA76.758 .A345 2001. URL

http://www.acm.org/pubs/
citations/proceedings/soft/
379605/p38-braband/. Sup-

plement to ACM SIGPLAN Notices.

Bordoni:2002:CCC

[Bor02] Luciana Bordoni. COVAX:

A contemporary culture virtual archive in XML. Lecture Notes in Computer Science, 2458:661–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.de/link/service/
series/0558/bibs/2458/24580661.
htm; http://link.springer.

                                        de/link/service/series/0558/
                                        papers/2458/24580661.pdf.

Brockmeier:2001:DXP

[BP01] Joe “Zonker” Brockmeier and Kara J. Pritchard. DocBook XML Publishing. Prima Publishing, Roseville, CA, USA, 2001. ISBN 0-7615-3331-1. 375 pp. LCCN ???? US$39.99. Includes CD-ROM.

Balmin:2005:SQX

[BP05] Andrey Balmin and Yannis Papakonstantinou. Storing and querying XML data using denormalized relational databases. VLDB Journal: Very Large Data Bases, 14(1): 30–49, March 2005. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Balmin:2004:IVX

[BPV04] Andrey Balmin, Yannis Papakonstantinou, and Victor Vianu. Incremental validation of XML documents. ACM Transactions on Database Systems, 29(4):710–751, December 2004. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Bailey:2002:AOE

[BPW02] James Bailey, Alexandra Poulovassilis, and Peter T. Wood. Analysis and optimisation of event-condition-action rules on XML. Computer Networks (Amsterdam, Netherlands: 1999), 39(3):239–259, June 21, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/
gej-ng/10/15/22/96/50/29/
abstract.html;   http://
www.elsevier.com/gej-ng/
10/15/22/96/50/29/article.
pdf.

Bradley:2000:XC

[Bra00] Neil Bradley. The XML companion. Addison-Wesley, Reading, MA, USA, second edition, 2000. ISBN 0-201-67486-6. x + 566 pp. LCCN QA76.76.H94 B77 2000.

Brownell:2002:S

[Bro02]  David Brownell.  SAX2.

O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00237-8. ix + 228 pp. LCCN QA76.76.H94 B795 2002 Stacks. US$29.95. URL http:/

/safari.oreilly.com/0596002378;
http://www.oreilly.com/
catalog/sax2.

Brookes:2003:XDB

[Bro03] Allen Brookes. XML data binding. Dr. Dobbs Journal, 28(3): 26, 28, 30, 32, 35–36, March 2003. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/documents/s=7826/
ddj0303c/.

Bry:2002:TDQ

[BS02] François Bry and Sebastian Schaffert. Towards a declarative query and transformation language for XML and semistructured data: Simulation unification. Lecture Notes in Computer Science, 2401:255–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.de/link/service/
series/0558/bibs/2401/24010255.
htm; http://link.springer.
de/link/service/series/0558/
papers/2401/24010255.pdf.

Box:2000:EXB

[BSL00] Don Box, Aaron Skonnard, and John Lam. Essential XML: Beyond Markup. Addison-Wes-ley, Reading, MA, USA, 2000. ISBN 0-201-70914-7. xviii + 368 pp. LCCN QA76.76.H94 B69 2000. US$34.95.

Burke:2001:JX

[Bur01] Eric M. Burke. Java and XSLT. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00143-

6. xvi + 510 pp. LCCN QA76.73.J38 B87 2001.

Burch:2002:FCS

[Bur02a] Glenda Burch. First commentary on “XML and the new design regime”. ACM Journal of Computer Documentation, 26 (2):43–44, May 2002. CODEN AJCDBH. ISSN 1527-6805.

Burke:2002:PLF

[Bur02b] Sean M. Burke. Perl and LWP:

fetching Web pages, parsing HTML, writing spiders and more. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00178-9. xv + 242 pp. LCCN QA76.73.P22 B87 2002 Stacks.

Burns:2002:HG

[Bur02c] Joe Burns. HTML goodies.

Que Corporation, Indianapolis, IN, USA, second edition, 2002. ISBN 0-7897-2611-4. xxix + 586 pp. LCCN QA76.76.H94 B87 2002 Computer.

Cagle:2000:XDH

[Cag00]  Kurt Cagle.  XML Devel-

oper’s Handbook. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2000. ISBN 0-7821-2704-5. xix + 619 pp. LCCN QA76.76.H94 C34 2000. US$49.99.

Callihan:2000:LHW

[Cal00] Steven E. Callihan. Learn HTML in a weekend. In a weekend. Prima Publishing, Roseville, CA, USA, third edition, 2000. ISBN 0-7615-2694-3. xxviii + 418 pp. LCCN QA76.76.H94 C35 2000.

Carey:2000:NPC

[Car00] Patrick Carey. New perspectives on creating Web pages with HTML: introductory. New perspectives series. Course Technology, Cambridge, MA, USA, second edition, 2000. ISBN 0-619-01967-0 (paperback). various pp. LCCN QA76.76.H94 C37 2000.

Carlisle:2002:MWU

[Car02a] David Carlisle. MathML on the Web: Using XSLT to enable cross-platform support for XHTML and MathML in current browsers. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Carminati:2002:SPX

[Car02b]  Barbara Carminati.  Se-

cure publishing of XML documents. Lecture Notes in Computer Science, 2490:587– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/
                                        service/series/0558/bibs/
                                        2490/24900587.htm; http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2490/24900587.pdf.

Castro:2000:HWW

[Cas00] Elizabeth Castro. HTML 4 for the World Wide Web. Visual quickstart guide. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, fourth edition, 2000. ISBN 0-201-35493-4. 384 pp. LCCN QA76.76.H94 C3698 2000.

Case:2006:EXS

[Cas06] P. Case. Enhancing XML search with XQuery 1.0 and XPath 2.0 full-text. IBM Systems Journal, 45(2):353– 360, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/452/case.html.

Cho:2003:LSE

[CAYLS03] SungRan Cho, Sihem Amer-Yahia, Laks V. S. Lakshmanan, and Divesh Srivastava. LockX: a system for efficiently querying secure XML. In ACM [ACM03a], page 669. ISBN ???? LCCN ????

Chalub:2004:MRS

[CB04] F. Chalub and C. Braga. A modular rewriting semantics for CML. J.UCS: Journal of Universal Computer Science, 10(7):789–??, July 28, 2004. CODEN ???? ISSN 0948-

6968. URL http://www.jucs.
org/jucs_10_7/a_modular_
rewriting_semantics.

Comai:2001:CGQ

[CDF01] Sara Comai, Ernesto Damiani, and Piero Fraternali. Computing graphical queries over XML data. ACM Transactions on Information Systems, 19(4):371– 430, October 2001. CODEN ATISET. ISSN 1046-8188.

Chen:2003:RRR

[CDHZ03] Yi Chen, Susan B. Davidson, Carmem S. Hara, and Yifeng Zheng. RRXF: Redundancy reducing XML storage in relations. In Freytag et al. [FLA^{+}03], pages 189–

200. ISBN 0-12-722442-4. URL

http://www.vldb.org/dblp/
db/indices/a-tree/c/Chen:
Yi.html.

Cerami:2002:WSE

[Cer02]  Ethan Cerami.  Web Ser-

vices Essentials: Distributed Applications with XML-RPC, SOAP, UDDI & WSDL. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00224-6. xiii + 288 pp. LCCN TK5105.888 .C47 2002. US$29.95. URL http://

safari.oreilly.com/0596002246;
 http://www.oreilly.com/
catalog/webservess.

Chan:2002:TPA

[CFF^{+}02] Chee Yong Chan, Wenfei Fan, Pascal Felber, Minos N. Garofalakis, and Rajeev Rastogi. Tree pattern aggregation for scalable XML data dissemination. In Bernstein et al. [B^{+}02], pages 826–837. ISBN 1-55860-869-9. LCCN ???? URL http:

//www.vldb.org/conf/2002/
S23P02.pdf.

Chan:2002:EFX

[CFGR02] C.-Y. Chan, P. Felber, M. Garofalakis, and R. Rastogi. Efficient filtering of XML documents with XPath expressions. VLDB Journal: Very Large Data Bases, 11(4):354– 379, December 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL http://link.springer.
de/link/service/journals/
00778/bibs/2011004/20110354.
htm;   http://link.springer.
de/link/service/journals/
00778/papers/2011004/20110354.
pdf.

Colazzo:2004:TPC

[CGMS04] Dario Colazzo, Giorgio Ghelli, Paolo Manghi, and Carlo Sartiani. Types for path correctness of XML queries. ACM SIGPLAN Notices, 39(9):126– 137, September 2004. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cagle:2000:BX

[CH00] Kurt Cagle and David Hunter.

Beginning XML. Wrox Press, Chicago, IL, USA, 2000. ISBN 1-86100-341-2. xix + 823 pp. LCCN QA76.76.H94 B44 2000.

Chang:2006:SXD

[CH06] Tao-Ku Chang and Gwan-Hwan Hwang. To secure XML documents with the extension function of XSLT. Software— Practice and Experience, 36(5):

539–555, April 25, 2006. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chang:2000:OXH

[Cha00]  Ben Chang.  Oracle XML

handbook. Osborne/McGraw-Hill, Berkeley, CA, USA, 2000. ISBN 0-07-212489-X. xxii + 357 pp. LCCN QA76.76.H94 O73 2000.

Chamberlin:2002:XXQ

[Cha02a]  D. Chamberlin.  XQuery:

An XML query language. IBM Systems Journal, 41(4): 597–??, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/414/chamberlin.pdf.

Chang:2002:XDR

[Cha02b] Jae-Woo Chang. An XML document retrieval system supporting structure- and contentbased queries. Lecture Notes in Computer Science, 2465:320– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2465/24650320.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2465/24650320.pdf.

Chamberlin:2003:XQL

[Cha03] Don Chamberlin. XQuery: a query language for XML. In ACM [ACM03a], page 682. ISBN ???? LCCN ????

Chatvichienchai:2002:TAA

[CIK02] S. Chatvichienchai, M. Iwaihara, and Y. Kambayashi. Translating access authorizations for transformed XML documents. Lecture Notes in Computer Science, 2453:290– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2453/24530290.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2453/24530290.
pdf.

Carey:2001:NPC

[CK01] Patrick Carey and Shirley E.

Kaiser. New perspectives on creating Web pages with Dynamic HTML: introductory. New perspectives series. Course Technology, Cambridge, MA, USA, 2001. ISBN 0-619-01918-2. various pp. LCCN QA76.76.H94 C373 2001.

Chung:2002:EII

[CK02] Tae-Sun Chung and Hyoung-Joo Kim. Extracting indexing information from XML DTDs. Information Processing Letters, 81(2):97–103, January 31, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.elsevier.com/
gej-ng/10/23/20/85/33/34/
abstract.html.

Chung:2003:EPX

[CKK03] Yon Dohn Chung, Jong Wook Kim, and Myoung Ho Kim. Efficient preprocessing of XML queries using structured signatures. Information Processing Letters, 87(5):257–264, September 15, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cohen:2002:LDX

[CKM02] Edith Cohen, Haim Kaplan, and Tova Milo. Labeling dynamic XML trees. In ACM [ACM02a], pages 271–281. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Chaudhuri:2003:RSX

[CKN03] Surajit Chaudhuri, Raghav Kaushik, and Jeffrey F. Naughton. On relational support for XML publishing: beyond sorting and tagging. In ACM [ACM03a], pages 611–622. ISBN ???? LCCN ????

Cohen:2002:SPQ

[CKS02a] Sara Cohen, Yaron Kanza, and Yehoshua Sagiv. Selectproject queries over XML documents. Lecture Notes in Computer Science, 2382:2–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2382/23820002.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/

0558/papers/2382/23820002.
pdf.

Cohen:2002:SFQ

[CKS02b] Sara Cohen, Yaron Kanza, and Yehoshua Sagiv. SQL4X: A flexible query language for XML and relational databases. Lecture Notes in Computer Science, 2397:263–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2397/23970263.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2397/23970263.pdf.

Chen:2004:CSI

[CL04] Cheng-Wei Chen and Jenq Kuen Lee. Case study: an infrastructure for C/ATLAS environments with object-oriented design and XML representation. The Journal of Systems and Software, 71(1–2):83–95, April 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cha:2002:IXB

[CLCC02] Si-Ho Cha, Jae-Oh Lee, Young-Keun Choi, and Kook-Hyun Cho. Implementing an XMLbased universal network management system in Java. Lecture Notes in Computer Science, 2343:652–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2343/23430652.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2343/23430652.pdf.

Cleaveland:2001:PGJ

[Cle01a] J. Craig Cleaveland. Program generators with Java and XML. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-025878-4. xxviii + 415 pp. LCCN QA76.73.J38 C54 2001. Includes CD-ROM.

Cleaveland:2001:PGX

[Cle01b] J. Craig Cleaveland. Program generators with XML and Java. The Charles F. Goldfarb series on open information management. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-025878-4. xxviii + 415 pp. LCCN QA76.73.J38 C54 2001. Includes CD-ROM.

Chen:2002:DVX

[CLL02] Ya Bing Chen, Tok Wang Ling, and Mong Li Lee. Designing valid XML views. Lecture Notes in Computer Science, 2503:463–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2503/25030463.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2503/25030463.pdf.

Cluet:2003:WNX

[Clu03] Sophie Cluet. Who needs XML databases? In Freytag et al. [FLA^{+}03], page 1045. ISBN 0-12-722442-4.

URL http://www.vldb.org/
dblp/db/indices/a-tree/c/
Cluet:Sophie.html.

Cohen:2003:XSS

[CMKS03] Sara Cohen, Jonathan Mamou, Yaron Kanza, and Yehoshua Sagiv. XSEarch: A semantic search engine for XML. In Freytag et al. [FLA^{+}03], pages 45–56. ISBN 0-12-722442-4.

URL http://www.vldb.org/
dblp/db/indices/a-tree/c/
Cohen:Sara.html.

Chung:2002:AAP

[CMS02] Chin-Wan Chung, Jun-Ki Min, and Kyuseok Shim. APEX: an adaptive path index for XML data. In Franklin et al. [FMA02], pages 121–132. ISBN ???? LCCN ???? ACM order number 475020.

Crawford:2002:FEE

[CNB^{+}02] Diane Crawford, Srinivas Nedunuri, Adrian Bowyer, Arnd Weber, Greg A. Woods, and Mark Adler. Forum: Embrace the engineering metaphor; credit for crypto’s parallel development; enough PDF: Give me HTML. Communications of the ACM, 45(8):11–14, August 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cochran:2001:NV

[Coc01] Shannon Cochran. News and views: Curl programming contest winners announced; prior art database will protect open software; AI researcher [Kenneth Mark Colby] dies; personal router for surfing wireless services; XML schema formalized; thought-controlled computers. Dr. Dobbs Journal, 26 (8):18, August 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/.

Cox:2001:PX

[Cox01a] David Cox. Parsing XML. Dr. Dobbs Journal, 26(1):96, 98, 100, January 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2001/2001_01/minixml.
                                        txt;   http://www.ddj.com/
                                        ftp/2001/2001_01/minixml.
                                        zip.

Cox:2001:XIC

[Cox01b] David Cox. XML for instrument control and monitoring. Dr. Dobbs Journal, 26(11):83– 85, November 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2001/2001_11/ixml.txt.

Coyle:2002:XWS

[Coy02] Frank P. Coyle. XML, Web Services, and the Data Revolution. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-77641-3. xxxii + 356 pp. LCCN QA76.76.H94 C69 2002.

Ciaccia:2002:AFS

[CP02] Paolo Ciaccia and Wilma Penzo. Adding flexibility to structure similarity queries on XML data. Lecture Notes in Computer Science, 2522:124– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2522/25220124.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2522/25220124.pdf.

Crane:2005:AA

[CPJ05] Dave Crane, Eric Pascarello, and Darren James. Ajax in Action. Manning Publications, Greenwich, CT, USA, 2005. ISBN 1-932394-61-3. 680 (est.) pp. LCCN ???? US$22.50 (electronic), US$44.95 (print + electronic). URL http://www.

manning.com/books/crane.

Cromwell:2001:PBD

[Cro01] Jeff Cromwell. Programmer’s bookshelf: Days of our data: A server SOAP opera. Dr. Dobbs Journal, 26(7):138, 140, July 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/.

Chen:2002:XSC

[CRW02] Li Chen, Elke A. Rundensteiner, and Song Wang. XCache: a semantic caching system for XML queries. In Franklin et al. [FMA02], page 618. ISBN ???? LCCN ???? ACM order number 475020.

Chang:2001:OXH

[CSK01] Ben Chang, Mark Scardina, and Stefan Kiritzov. Oracle9i XML handbook. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-213495-X. xxiii + 495 pp. LCCN QA76.76 .H94 C53 2001. CD-ROM includes the official Oracle XML Developer Kit (XDK).

Chien:2002:ESM

[CTZ02] S.-Y. Chien, V. J. Tsotras, and C. Zaniolo. Efficient schemes for managing multiversion XML documents. VLDB Journal: Very Large Data Bases, 11(4):332–353, December 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL http://link.springer.
de/link/service/journals/
00778/bibs/2011004/20110332.
htm;   http://link.springer.
de/link/service/journals/
00778/papers/2011004/20110332.
pdf.

Chien:2002:ESJ

[CVZ^{+}02] Shu-Yao Chien, Zografoula Vagena, Donghui Zhang, Vassilis J. Tsotras, and Carlo Zaniolo. Efficient structural joins on indexed XML documents. In Bernstein et al. [B^{+}02], pages 263–274. ISBN 1-55860-869-

9. LCCN ???? URL http:

//www.vldb.org/conf/2002/
S08P03.pdf.

Chaudhri:2001:SOD

[CZ01] Akmal B. Chaudhri and Roberto Zicari. Succeeding with object databases: a practical look at today’s implementations with Java and XML. Wiley, New York, NY, USA, 2001. ISBN 0-471-38384-8 (cloth). xxi + 442 pp. LCCN QA76.9.D3 C3848 2001.

Diao:2003:PSP

[DAF^{+}03] Yanlei Diao, Mehmet Altinel, Michael J. Franklin, Hao Zhang, and Peter Fischer. Path sharing and predicate evaluation for high-performance XML filtering. ACM Transactions on Database Systems, 28(4): 467–516, December 2003. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Danielson:2000:FAD

[Dan00] Richard Karl Danielson. Flash animation and DHTML: the best work from the Web. Web site graphics. Rockport Publishers, Gloucester, MA, USA, 2000. ISBN 1-56496-722-0. 96 pp. LCCN TK5105.888.D359 2000. US$20.00.

Daum:2003:MBO

[Dau03]  Berthold Daum.  Modeling

Business Objects with XML Schema. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 1-55860-816-8. xxv + 535 pp. LCCN QA76.76.H94 D418 2003. US$49.95.

Davis:2001:DSE

[Dav01] Don Davis. Defective sign-and-encrypt: Can you really trust S/MIME, PCKS#7, PGP, and XML? Dr. Dobbs Journal, 26 (11):30, November 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/.

David:2003:ASH

[Dav03]  Michael M. David.  ANSI

SQL hierarchical processing can fully integrate native XML. SIGMOD, 32(1):41–46, March 2003.

Damiani:2002:FGA

[DDPS02] Ernesto Damiani, Sabrina De Capitani di Vimercati, Stefano Paraboschi, and Pierangela Samarati. A fine-grained access control system for XML documents. ACM Transactions on Information and System Security, 5(2):169–202, May 2002. CODEN ATISBQ. ISSN 1094-9224 (print), 1557-7406 (electronic).

Deitel:2001:XHP

[Dei01] Harvey M. Deitel. XML: how to program. How to program series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-028417-3. xxxviii + 934 pp. LCCN QA76.76.H94 X45 2001. Includes CD-ROM.

Dellwig:2002:H

[Del02] Ingo Dellwig. HTML 4.0. Addison-Wesley nitty gritty programming series. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-75877-6 (paperback). xiii + 258 pp. LCCN QA76.76.H94 D46 2002 Computer.

Desmarais:2000:AXL

[Des00] Norman Desmarais. The ABCs of XML: the librarian’s guide to the extensible markup language. New Technology Press, Houston, TX, 2000. ISBN 0-9675942-0-0 (paperback). x + 206 pp. LCCN QA76.76.H94 D48 2000.

Diao:2003:QPH

[DF03] Yanlei Diao and Michael J.

Franklin. Query processing for high-volume XML message brokering. In Freytag et al. [FLA^{+}03], pages 261– 272. ISBN 0-12-722442-4. URL

http://www.vldb.org/dblp/
db/indices/a-tree/d/Diao:
Yanlei.html.

Delalandre:2002:SSA

[DHA^{+}02] Mathieu Delalandre, Pierre Héroux, Sébastien Adam, Eric Trupin, and Jean-Marc Ogier. A statistical and structural approach for symbol recognition, using XML modelling. Lecture Notes in Computer Science, 2396:281–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2396/23960281.
htm;   http://link.springer-

                                        ny.com/link/service/series/
                                        0558/papers/2396/23960281.
                                        pdf.

Diehl:2001:DVW

[Die01] Stephan Diehl. Distributed virtual worlds: foundations and implementation techniques using VRML, Java, and CORBA. Springer-Verlag Inc., New York, NY, USA, 2001. ISBN 3-540-67624-4. xii + 166 pp. LCCN QA76.9.D5 D427 2001.

Dix:2001:WSS

[Dix01]  Chris Dix.  Working with

SOAP, the Simple Object Access Protocol. C/C++ Users Journal, 19(1):22–??, January 2001. CODEN CCUJEX. ISSN 1075-2838.

Demey:2002:CML

[DJM02] Jan Demey, Mustafa Jarrar, and Robert Meersman. A conceptual markup language that supports interoperability between business rule modeling systems. Lecture Notes in Computer Science, 2519:19– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/
                                        service/series/0558/bibs/
                                        2519/25190019.htm; http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2519/25190019.pdf.

Dong:2004:HXS

[DL04] Ying Dong and Mingshu Li. HyO-XTM: a set of hypergraph operations on XML Topic Map toward knowledge management. Future Generation Computer Systems, 20(1): 81–100, January 15, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Davenport:2008:FEO

[DL08] James H. Davenport and Paul Libbrecht. The freedom to extend OpenMath and its utility. Mathematics in Computer Science, 2(2):379–398, December 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://www.

springerlink.com/openurl.
asp?genre=article&issn=1661-
8270&volume=2&issue=2&spage=
379.

Dong:2003:XBS

[DLS^{+}03] Jin Song Dong, Yuan Fang Li, Jing Sun, Jun Sun, and Hai Wang. XML-based static type checking and dynamic visualization for TCOZ. Lecture Notes in Computer Science, 2495:311–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2495/24950311.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2495/24950311.pdf.

Daum:2003:SAX

[DM03] Berthold Daum and Udo Merten. System architecture with XML. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 1-55860-745-5, 3-89864-196-1. xxii + 458 pp. LCCN QA76.9.A73 D384 2003.

Dooley:2002:T

[Doo02a] Sam Dooley. techexplorer. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Dooley:2002:BMC

[Doo02b] Samuel S. Dooley. Bringing MathML content and presentation markup to the Web with the IBM MathML Expression Editor. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Dournaee:2002:XS

[Dou02] Blake Dournaee. XML security. RSA Press, Berkeley, CA, USA, 2002. ISBN 0-07-219399-9. xiii + 379 pp. LCCN QA76.76.H94 D68 2002.

Daconta:2000:XDJ

[DS00] Michael C. Daconta and Al Saganich. XML development with Java 2. SAMS Publishing, Indianapolis, IN, USA, 2000. ISBN 0-672-31653-6. ix + 429 pp. LCCN QA76.76.H94 D32 2000.

Deutsch:2003:MSP

[DT03] Alin Deutsch and Val Tannen. MARS: A system for publishing XML from mixed and redundant storage. In Freytag et al. [FLA^{+}03], pages 201–212. ISBN 0-12-722442-4.

URL http://www.vldb.org/
dblp/db/indices/a-tree/d/
Deutsch:Alin.html.

Deutsch:2005:XQC

[DT05] Alin Deutsch and Val Tannen. XML queries and constraints, containment and reformulation. Theoretical Computer Science, 336(1):57–87, May 25, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deMeo:2004:XGS

[dTU04] P. de Meo, G. Terracina, and D. Ursino. X-Global: a system for the “almost automatic” and semantic integration of XML sources at various flexibility levels. J.UCS: Journal of Universal Computer Science, 10(9):1065–??, September 28, 2004. CODEN ???? ISSN 0948-

6968. URL http://www.jucs.
org/jucs_10_9/x_global_a_
system.

Dwelly:2000:JXL

[Dwe00a] Andrew Dwelly. Java, XML, and literate programming. Dr. Dobbs Journal, 25(2):62, 64–66, 68, February 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_02/litjava.
txt;   http://www.ddj.com/
ftp/2000/2000_02/litjava.
zip.

Dwelly:2000:XRP

[Dwe00b] Andrew Dwelly. XML, reflective pattern matching, and Java. Dr. Dobbs Journal, 25 (6):46, 49–52, 54, June 2000. CODEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_06/
marius05.zip;   http://www.
ddj.com/ftp/2000/2000_06/
xmljava.txt.

Eckstein:2001:XPR

[EC01] Robert Eckstein and Michel Casabianca. XML pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2001. ISBN 0-596-00133-9. vi + 96 pp. LCCN QA76.76.H94 E25 2001. Covers XSTL.

Eddy:2000:XPEa

[ED00] Sandra E. Eddy and B. K. De-Long. XML In Plain English. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, second edition, 2000. ISBN 0-7645-4744-5. xx + 860 pp. LCCN QA76.76.H94 E33 2001. US$19.99.

Eddy:2001:XPE

[ED01] Sandra E. Eddy and B. K. Delong. XML in plain English. Professional mindware. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, second edition, 2001. ISBN 0-7645-4744-5 (paperback). xx + 860 pp. LCCN QA76.76.H94 E33 2001.

Eddy:2000:XPEb

[Edd00] Sandra E. Eddy. XHTML in plain English. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 2000. ISBN 0-7645-4743-7. xxiii + 830 pp. LCCN QA76.76.H94

E35 2000. URL http://www.
hungryminds.com/mandtbooks/
extras/xhtmlipe/index.html.

Edgar:2001:VHU

[Edg01] Bob Edgar. The VoiceXML Handbook: Understanding and Building the Phone-Enabled Web. CMP Books, 6600 Silacci Way, Gilroy, CA 95020, USA, 2001. ISBN 1-57820-084-9. 481 pp. LCCN QA76.76.H94 E38 2001. US$39.95.

Eberhart:2002:JTU

[EF02] Andreas Eberhart and Stefan Fischer. Java tools: using XML, EJB, CORBA, Servlets and SOAP. Wiley, New York, NY, USA, 2002. ISBN 0-471-48666-3. 493 pp. LCCN QA76.73.J38 E24

2002. URL http://www.i-u.
de/schools/eberhart/book/.

Ehlen:2000:HC

[Ehl00] Anamary Ehlen. HTML complete. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, second edition, 2000. ISBN 0-7821-2801-7. xxvi + 1010 pp. LCCN QA76.76.H94 H763 2000.

Eisenberg:2002:SE

[Eis02] J. David Eisenberg. SVG Essentials. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00223-8. xvii + 335 pp. LCCN T385 .E47 2002. US$34.95. URL http://

                                        safari.oreilly.com/0596002238;
                                         http://www.oreilly.com/
                                        catalog/svgess.

Eisenberg:2002:SXM

[EM02] Andrew Eisenberg and Jim Melton. SQL/XML is making good progress. SIGMOD, 31 (2):101–108, June 2002.

Ennser:2000:IXD

[EMS00] Luis Ennser, Christophe Delporte Mikio, and Oba K. M. Sunil. Integrating XML with DB2 XML Extender and DB2 Text Extender. IBM Redbooks, 11400 Burnet Road, Austin, TX 78758-3493, USA, 2000. ISBN 0-7384-2143-X. 340 pp. LCCN ???? URL

                                        http://www.redbooks.ibm.
                                        com/pubs/pdfs/redbooks/sg246130.
                                        pdf;   http://www.redbooks.
                                        ibm.com/redbooks/SG246130.

html. IBM order number SG24-

6130-00.

Endres:2000:GCM

[End00]  Stephan Endres.  Genetic

Coding mittels XML in der medizinischen Dokumentation.

(German) [Genetic Coding by means of XML in medical documentation]. Thesis (doctoral), Ludwig-Maximilians-Universitat zu München, München, Germany, 2000. 94 pp.

Eggert:2005:PEN

[EP05] P. R. Eggert and D. S. Parker.

Perturbing and evaluating numerical programs without recompilation — the wonglediff way. Software—Practice and Experience, 35(4):313–322, April 10, 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Erickson:2001:EWA

[Eri01] Jonathan Erickson. Editorial:

R_{XML} for what ails healthcare. Dr. Dobbs Journal, 26 (11):8, November 2001. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/.

Erickson:2003:DDE

[Eri03] Jon Erickson. 2003 Dr. Dobb’s Excellence in Programming Award. Dr. Dobbs Journal, 28(4):16–17, April 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7827/
ddj0304b/.

Evjen:2007:PX

[EST^{+}07] Bill Evjen, Kent Sharkey, Thiru Thangarathinam, Michael Kay, Alessandro Vernet, and Sam Ferguson, editors. Professional XML. Wrox professional guides. Wrox Press, Chicago, IL, USA, 2007. ISBN 0-471-77777-3 (paperback). xxix + 856 pp. LCCN QA76.76.H94 P7638

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0713/2007006214-d.html;
 http://www.loc.gov/catdir/
enhancements/fy0741/2007006214-
b.html;   http://www.loc.
gov/catdir/toc/ecip0710/
2007006214.html.

Embley:2002:AEO

[ETL02] David W. Embley, Cui Tao, and Stephen W. Liddle. Automatically extracting ontologically specified data from HTML tables of unknown structure. Lecture Notes in Computer Science, 2503:322– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2503/25030322.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2503/25030322.pdf.

Elmasri:2002:CMC

[EWH^{+}02] Ramez Elmasri, Yu-Chi Wu, Babak Hojabri, Charley Li, and Jack Fu. Conceptual modeling for customized XML schemas. Lecture Notes in Computer Science, 2503:429– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/

2503/25030429.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2503/25030429.pdf.

Finkelstein:2000:BCP

[FA00] Clive Finkelstein and Peter Aiken. Building corporate portals using XML. McGraw-Hill, New York, NY, USA, 2000. ISBN 0-07-913705-9. xxvii + 529 pp. LCCN QA76.76.H94F56 1999.

Falco:2000:JBX

[Fal00a] Joe Falco. Java-based XML utility for the NIST machine tool data repository. ????, November 2000. 13 pp. Shipping list number 2001-0146-M.

Falco:2000:JXU

[Fal00b] Joe Falco. Java-based XML utility for the NIST machine tool data repository. Gaithersburg, MD, USA, November 2000. 13 pp. Shipping list no.: 2001-0146-M.

Freire:2004:MXD

[FB04] Juliana Freire and Michael Benedikt. Managing XML data: An abridged overview. Computing in Science and Engineering, 6(4):12–19, July/ August 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

URL http://csdl.computer.
org/dl/mags/cs/2004/04/c4012.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/04/c4012.
pdf.

Feng:2002:SNB

[FCD02] Ling Feng, Elizabeth Chang, and Tharam Dillon. A semantic network-based design methodology for XML documents. ACM Transactions on Information Systems, 20 (4):390–421, 2002. CODEN ATISET. ISSN 1046-8188.

Feng:2003:STO

[FCD03] L. Feng, E. Chang, and T. S. Dillon. Schemata transformation of object-oriented conceptual models to XML. International Journal of Computer Systems Science and Engineering, 18(1):??, January 2003. CODEN CSSEEI. ISSN 0267-6192.

Feng:2004:STO

[FCD04] L. Feng, E. Chang, and T. S. Dillon. Schemata transformation of object-oriented conceptual models to XML. International Journal of Computer Systems Science and Engineering, 19(1):??, January 2004. CODEN CSSEEI. ISSN 0267-6192.

Feinberg:2005:NXD

[Fei05] George Feinberg. Native XML database storage and retrieval. Linux Journal, 2005(137):??, September 2005. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Fiebig:2002:ANX

[FHK^{+}02] T. Fiebig, S. Helmer, C.-C. Kanne, G. Moerkotte, J. Neumann, R. Schiele, and T. Westmann. Anatomy of a native XML base management system. VLDB Journal: Very Large Data Bases, 11(4):292– 314, December 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL http://link.springer.
de/link/service/journals/
00778/bibs/2011004/20110292.
htm;   http://link.springer.
de/link/service/journals/
00778/papers/2011004/20110292.
pdf.

Freire:2002:SMX

[FHR^{+}02] Juliana Freire, Jayant R. Haritsa, Maya Ramanath, Prasan Roy, and Jérôme Siméon. StatiX: making XML count. In Franklin et al. [FMA02], pages 181–191. ISBN ???? LCCN ???? ACM order number 475020.

Fiebig:2000:UCA

[Fie00]  Annegret Fiebig.  Urkun-

dentext: computergestutzte Auswertung deutschsprachiger Urkunden der Kuenringer auf Basis der eXtensible Markup Language (XML). (German) [Document text: Computeraided analysis of Germanlanguage documents of the Kuenringer on the basis of eXtensible Markup Language (XML)]. Number 333 in Schriften zur sudwestdeutschen Landeskunde. DRW-Verlag, Leinfelden, Germany, 2000. ISBN 3-87181-433-4. ix + 246 pp. LCCN QA76.9.T48 F53 2000. Originally presented as the author’s thesis (doctoral)–Freie Universität, Berlin, Wintersemester, 1996 and 1997, with title: Urkundentext als Textsystem: computergesturte Auswertung mittelhochdeutscher Urkunden als Beitrag zur volkssprachlichen Schriftlichkeit.

Fitzgerald:2001:XE

[Fit01] Michael Fitzgerald. XSL essentials. Wiley, New York, NY, USA, 2001. ISBN 0-471-41620-7. 312 pp. LCCN QA76.76.H94 F473 2001.

Fitzgerald:2003:LX

[Fit03] Michael Fitzgerald. Learning XSLT. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00327-7. xiii + 352 pp. LCCN QA76.73.X58 F58 2003. US$34.95, CDN$54.95.

                                        URL http://press.oreilly.
                                        com/pub/pr/1126.

Feng:2004:PEX

[FJ04] L. Feng and W. Jonker. Preparations for encrypted XML metadata querying. International Journal of Computer Systems Science and Engineering, 19(3):??, May 2004. CO-DEN CSSEEI. ISSN 0267-6192.

Fan:2002:UCM

[FKS02a] Wenfei Fan, Gabriel M. Kuper, and Jérôme Siméon. A unified constraint model for XML. Computer Networks (Amsterdam, Netherlands: 1999), 39(5):489–505, August 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/
gej-ng/10/15/22/96/52/29/
abstract.html.

Fernandez:2002:SFP

[FKS^{+}02b] Mary Fernández, Yana Kadiyska, Dan Suciu, Atsuyuki Morishima, and Wang-Chiew Tan. SilkRoute: A framework for publishing relational data in XML. ACM Transactions on Database Systems, 27(4):438– 493, December 2002. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Funderburk:2002:XBR

[FKS^{+}02c] J. E. Funderburk, G. Kiernan, J. Shanmugasundaram, E. Shekita, and C. Wei. XTA-BLES: Bridging relational technology and XML. IBM Systems Journal, 41(4):616–??, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/414/funderburk.pdf.

Funderburk:2003:TNX

[FKS^{+}03] J. E. Funderburk, G. Kiernan, J. Shanmugasundaram, E. Shekita, and C. Wei. Technical note—XTABLES: Bridging relational technology and XML. IBM Systems Journal, 42(3):538–541, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/423/funderburk.pdf.

Fan:2002:XIC

[FL02] Wenfei Fan and Leonid Libkin. On XML integrity constraints in the presence of DTDs. Journal of the ACM, 49(3):368–406, May 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Freytag:2003:VPI

[FLA^{+}03] Johann Christoph Freytag, Peter C. Lockemann, Serge Abiteboul, Michael J. Carey, Patricia G. Selinger, and Andreas Heuer, editors. VLDB 2003: Proceedings of 29th International Conference on Very Large Data Bases, September 9–12, 2003, Berlin, Germany. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 0-12-722442-4.

Fuhr:2005:AXI

[FLMS05] Norbert Fuhr, Mounia Lalmas, Saadia Malik, and Zoltán Szlávik, editors. Advances in XML information retrieval: third international workshop of the initiative for the evaluation of XML retrieval, INEX 2004, Dagstuhl Castle, Germany, December 6–8, 2004. Revised selected papers, volume 3493 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-26166-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.
springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=3493;   http:
//www.springerlink.com/openurl.
asp?genre=volume&id=doi:
10.1007/b136250.

Floyd:2000:BWS

[Flo00] Michael Floyd. Building Web sites with XML. The Charles F. Goldfarb series on open information management. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-086601-6. xxx + 418 pp. LCCN QA76.76.H94 F58 2000.

Franklin:2002:PAS

[FMA02] Michael Franklin, Bongki Moon, and Anastassia Ailamaki, editors. Proceedings of the ACM SIGMOD International Conference on Management of Data, June 3– 6, 2002, Madison, WI, USA. ACM Press, New York, NY, USA, 2002. ISBN ???? LCCN ???? ACM order number 475020.

Ferreras:2002:HVF

[FMP02] César González Ferreras, David Escudero Mancebo, and Valentín Cardeñoso Payo. From HTML to VoiceXML: A first approach. Lecture Notes in Computer Science, 2448:441– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2448/24480441.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2448/24480441.
                                        pdf.

Fialli:2003:LXD

[FMPL^{+}03] Joe Fialli, Mike Morley, Eduardo Pelegri-Llopart, Arpad Elo, Jr., and Joseph Roy D. North. Letters: XML data binding; XML-based programming; block-structured languages; history lesson. Dr. Dobbs Journal, 28(6):10, June 2003. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/documents/s=8213/
                                        ddj0306l/.

Funderburk:2002:XPS

[FMR02] J. E. Funderburk, S. Malaika, and B. Reinwald. XML programming with SQL/XML and XQuery. IBM Systems Journal, 41(4):642–??, ???? 2002. CODEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/414/reinwald.pdf.

Fernandez:2001:EEX

[FMS01] Mary Fernandez, Atsuyuki Morishima, and Dan Suciu. Efficient evaluation of XML middle-ware queries. In Sellis and Mehrotra [SM01], pages 103–114. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Formica:2008:SXS

[For08]  Anna Formica.  Similarity

of XML-schema elements: A structural and information content approach. The Computer Journal, 51(2):240–254, March 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/2/240;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/2/240;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/2/240.

Fox:2002:XIB

[Fox02] Geoffrey Fox. XML and the importance of being an object. Computing in Science and Engineering, 4(3):96–98, May/ June 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/comp/
mags/cs/2002/03/c3096abs.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/03/c3096.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/03/c3096.
pdf.

Fraternali:2000:MDD

[FP00] Piero Fraternali and Paolo Paolini. Model-driven development of Web applications: the AutoWeb system. ACM Transactions on Information Systems, 18(4): 323–382, 2000. CODEN ATISET. ISSN 1046-8188.

                                        URL http://www.acm.org/
                                        pubs/citations/journals/
                                        tois/2000-18-4/p323-fraternali/

.

Freeby:2001:CDJ

[Fre01] James M. Freeby. A comparison of DHTML and JAVA applets. Thesis (M.S.), University of California, Santa Cruz, Santa Cruz, CA, USA, 2001.

Fung:2000:XWX

[Fun00] Khun Yee Fung. XSLT: working with XML and HTML. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-71103-

6. xx + 441 pp. LCCN QA76.76.H94 F86 2001. CD-ROM includes a case study illustrating all the major concepts and techniques in the book, all the self-contained examples in the book, and code for the case study and examples.

Gross-Amblard:2003:QPW

[GA03] David Gross-Amblard. Querypreserving watermarking of relational databases and XML documents. In ACM [ACM03b], pages 191–201. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Grose:2002:MXJ

[GDB02] Timothy J. Grose, Gary C. Doney, and Stephen A. Brodsky. Mastering XMI: Java programming with XMI, XML, and UML. Wiley, New York, NY, USA, 2002. ISBN 0-471-38429-1. xli + 434 pp. LCCN QA76.73.J38 G76 2002 Bar. Includes CD-ROM.

Georgescu:2002:CGT

[Geo02] Cristian Georgescu. Code generation templates using XML and XSL. C/C++ Users Journal, 20(1):??, January 2002. CODEN CCUJEX. ISSN 1075-2838.

Garelli:2002:DMM

[GF02] Francesco Garelli and Carlo Ferrari. A dynamic model for mapping XML elements in a object-oriented fashion. Lecture Notes in Computer Science, 2519:1255–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2519/25191255.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2519/25191255.pdf.

Green:2004:PXS

[GGM^{+}04] Todd J. Green, Ashish Gupta, Gerome Miklau, Makoto Onizuka, and Dan Suciu. Processing XML streams with deterministic automata and stream indexes. ACM Transactions on Database Systems, 29(4):752– 788, December 2004. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Gilorien:2000:DJ

[Gil00]  Gilorien.  DHTML and

JavaScript. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-086334-3. xx + 1142 pp. LCCN QA76.76.H94 G55 2000.

Guha:2002:AXJ

[GJK^{+}02] Sudipto Guha, H. V. Jagadish, Nick Koudas, Divesh Srivastava, and Ting Yu. Approximate XML joins. In Franklin et al. [FMA02], pages 287–298. ISBN ???? LCCN ???? ACM order number 475020.

Guha:2006:IXD

[GJK^{+}06] Sudipto Guha, H. V. Jagadish, Nick Koudas, Divesh Srivastava, and Ting Yu. Integrating XML data sources using approximate joins. ACM Transactions on Database Systems, 31 (1):161–207, March 2006. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Garofalakis:2003:CXD

[GK03] Minos Garofalakis and Amit Kumar. Correlating XML data streams using tree-edit distance embeddings. In ACM [ACM03b], pages 143–154. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Garofalakis:2005:XSP

[GK05] Minos Garofalakis and Amit Kumar. XML stream processing using tree-edit distance embeddings. ACM Transactions on Database Systems, 30(1): 279–332, March 2005. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Gottlob:2005:CXQ

[GKPS05] Georg Gottlob, Christoph Koch, Reinhard Pichler, and Luc Segoufin. The complexity of XPath query evaluation and XML typing. Journal of the ACM, 52(2):284–335, March 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gabeler-Lee:2003:LMM

[GLFO^{+}03] Matthew Gabeler-Lee, Daniel Fremberg, J. G. Owen, Gregory A. Janie, and Mike Moretti. Letters: Mithra; more business process outsourcing; SquareList; XML data binding. Dr. Dobbs Journal, 28(7):10, July 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/.

Goncalves:2002:XLS

[GLS^{+}02] Marcos André Gonçalves, Ming Luo, Rao Shen, Mir Farooq Ali, and Edward A. Fox. An XML log standard and tool for digital library logging analysis. Lecture Notes in Computer Science, 2458:129– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/

2458/24580129.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2458/24580129.pdf.

Gkoutos:2001:CMX

[GMRRW01] Georgios V. Gkoutos, Peter Murray-Rust, Henry S. Rzepa, and Michael Wright. Chemical markup, XML and the World-Wide Web. Part III: Towards a signed semantic Chemical Web of Trust. Journal of Chemical Information and Computer Sciences, 41(5):1124–1130, ???? 2001. CODEN JCISD8. ISSN 0095-2338. URL http:

//pubs.acs.org/subscribe/
journals/jcisd8/suppinfo/
ci000406v/exe/xsignarticle.
html;   http://pubs.acs.
org/subscribe/journals/jcisd8/
suppinfo/ci000406v/exe/xsignarticle.
pdf;   http://pubs.acs.org/
subscribe/journals/jcisd8/
suppinfo/ci000406v/exe/xsignarticle.
xml;   http://pubs.acs.org/
subscribe/journals/jcisd8/
suppinfo/ci000406v/exe/xsignarticle_
xslt.xml.

Garruzzo:2002:XCX

[GMRU02] Salvatore Garruzzo, Stefano Modafferi, Domenico Rosaci, and Domenico Ursino. Xcompass: An XML agent for supporting user navigation on the Web. Lecture Notes in Computer Science, 2522:197– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/

2522/25220197.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2522/25220197.pdf.

Goldman:2000:LDM

[GMW00] Roy Goldman, Jason McHugh, and Jennifer Widom. Lore: A database management system for XML. Dr. Dobbs Journal, 25(4):76, 78–80, April 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2000/2000_04/
lore.txt.

Goldberg:2009:XVQ

[Gol09] Kevin Howard Goldberg. XML:

Visual QuickStart Guide. Visual quickstart guide. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, second edition, 2009. ISBN 0-321-55967-3 (paperback). xviii + 269 pp. LCCN QA76.76.H94 G64 2009.

Goodman:2002:DHD

[Goo02] Danny Goodman. Dynamic HTML: The Definitive Reference: A Comprehensive Resource for HTML, CSS, DOM & JavaScript. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00316-1. 1418 (est.) pp. LCCN ???? US$59.95. URL http:/

/www.oreilly.com/catalog/
dhtmlref2.

Goschl:2003:JXB

[Gös03] Siegfried Göschl. Jelly: An XML-based scripting language. Dr. Dobbs Journal, 28(7): 66–67, July 2003. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2003/2003_07/
                                        jelly.zip.

Goldfarb:2000:XH

[GP00] Charles F. Goldfarb and Paul Prescod. The XML Handbook. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 2000. ISBN 0-13-014714-1. lvi + 1013 pp. LCCN QA76.76.H94 G65 2000. US$44.99.

Goldfarb:2001:CFG

[GP01a] Charles F. Goldfarb and Paul Prescod. Charles F. Goldfarb’s XML handbook. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, fourth edition, 2001. ISBN 0-13-065198-2. lvii + 1147 pp. LCCN QA76.76.H94 G65 2001. Includes CD-ROM.

Goldfarb:2001:XH

[GP01b] Charles F. Goldfarb and Paul Prescod. The XML handbook. The definitive XML series from Charles F. Goldfarb. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, third edition, 2001. ISBN 0-13-055068-X (paperback/CD-ROM). lxiv + 988 pp. LCCN QA76.76.H94 G65 2001.

Goldfarb:2002:CFG

[GP02] Charles F. Goldfarb and Paul Prescod. Charles F. Goldfarb’s XML handbook. Charles F. Goldfarb definitive XML series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, fourth edition, 2002. ISBN 0-13-065198-2. lvii + 1147 pp. LCCN QA76.76.H94 G65 2002.

Goldfarb:2004:CFG

[GP04] Charles F. Goldfarb and Paul Prescod. Charles F. Goldfarb’s XML handbook. Charles F. Goldfarb definitive XML series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, fifth edition, 2004. ISBN 0-13-049765-7. lv + 1222 pp. LCCN ????

Gardner:2002:XXG

[GR02] John Robert Gardner and Zarella L. Rendon. XSLT and XPath: a guide to XML transformations. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-040446-2. xxii + 558 pp. LCCN QA76.76.H94 R46 2002.

Graham:2000:XLD

[Gra00a] Ian S. Graham. XHTML 1.0 Language and Design Sourcebook: the Next Generation of HTML. Wiley, New York, NY, USA, 2000. ISBN 0-471-37485-7 (paperback). xxv + 692 pp. LCCN QA76.76.H94 G7334

2000. URL http://www.loc.
gov/catdir/enhancements/
fy0706/99089119-b.html;

                                        http://www.loc.gov/catdir/
                                        toc/onix04/99089119.html.

Graham:2000:XWD

[Gra00b] Ian S. Graham. XHTML 1.0 Web development sourcebook: building better sites and applications. Wiley, New York, NY, USA, 2000. ISBN 0-471-37486-5 (paper/Website). xxii + 471 pp. LCCN QA76.76.H94 G73345 2000.

Griffith:2002:JXJ

[Gri02] Arthur Griffith. Java, XML, and JAXP. Wiley, New York, NY, USA, 2002. ISBN 0-471-20907-4 (PAPER/WEBSITE). xii + 222 pp. LCCN QA76.73.J38 G77 2002 Bar.

Gao:2003:TSE

[GS03] Dengfeng Gao and Richard T. Snodgrass. Temporal slicing in the evaluation of XML queries. In Freytag et al. [FLA^{+}03], pages 632– 643. ISBN 0-12-722442-4.

                                        URL http://www.vldb.org/
                                        dblp/db/indices/a-tree/g/
                                        Gao:Dengfeng.html.

Guo:2003:XRK

[GSBS03] Lin Guo, Feng Shao, Chavdar Botev, and Jayavel Shanmugasundaram. XRANK: ranked keyword search over XML documents. In ACM [ACM03a], pages 16–27. ISBN ???? LCCN ????

Gupta:2003:VSP

[GSH03] Ashish Kumar Gupta, Dan Suciu, and Alon Y. Halevy.

The view selection problem for XML content based routing. In ACM [ACM03b], pages 68–77. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Gottleber:2000:MEH

[GT00] Timothy T. Gottleber and Timothy N. Trainor. More excellent HTML with an introduction to JavaScript. Irwin/ McGraw Hill, Boston, MA, USA, 2000. ISBN 0-07-233745-1, 0-07-233839-3. xxiii + 550 pp. LCCN QA76.76.H94G69 2000.

Gunton:2001:SSD

[Gun01a] Neil Gunton. SOAP: Simplifying distributed development: Putting the Simple Object Access Protocol to work. Dr. Dobbs Journal, 26(9):89, 91– 92, 95, September 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_09/
soap.txt.

Gunton:2001:WSS

[Gun01b] Neil Gunton. Web site searching and indexing in Perl. Dr. Dobbs Journal, 26(4):106, 108– 109, 111–112, April 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_04/
mysearch.txt;   http://www.
ddj.com/ftp/2001/2001_04/
mysearch.zip.

Guruge:2002:CPE

[Gur02] Anura Gurugé. Corporate Portals Empowered with XML and Web Services. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2002. ISBN 1-55558-280-X. xx + 313 pp. LCCN TK5105.875.I6 G87 2003. US$39.99.

Gutierrez:2004:IJU

[Gut04] Jaime Gutierrez, editor. ISAAC 2004: July 4–7, 2004, University of Cantabria, Santander, Spain: proceedings of the 2004 International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY, USA, 2004. ISBN 1-58113-827-X. LCCN ????

Gohel:2001:LGK

[GWT^{+}01] Himanshu Gohel, Thomas Weidenfeller, Andrew D. Todd, Ze’ev Atlas, and Dave Paris. Letters: GNOME, KDE clarification; parsing XML; CPRM frenzy; deadlock detecting; Perl versus the world. Dr. Dobbs Journal, 26(5):10, 12, May 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/.

Hunter:2007:BX

[H^{+}07] David Hunter et al., editors. Beginning XML. Wrox beginning guides. Wrox Press, Chicago, IL, USA, fourth edition, 2007. ISBN 0-470-11487-8 (paperback). xxxvi + 1039 pp. LCCN QA76.76.H94 B439

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0827/2007006580-b.html;
http://www.loc.gov/catdir/
enhancements/fy0827/2007006580-
d.html;   http://www.loc.
gov/catdir/toc/ecip0711/
2007006580.html.

Habel:2002:RCL

[Hab02] Geneviève Habel. Rainbow:

Complete learning environment and learning management system. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Harold:2001:XB

[Har01] Elliotte Rusty Harold. XML bible. Hungry Minds, 909 Third Avenue, New York, NY 10022, USA, second edition, 2001. ISBN 0-7645-4760-7. xxxiii + 1206 pp. LCCN QA76.76.H94 H34 2001. US$49.99. URL

http://www.cafeconleche.
org/books/bible2/.

Harris:2002:IMM

[Har02] Jason Harris. Implementing MathML in Mathematica. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Harold:2003:PXJ

[Har03] Elliotte Rusty Harold. Processing XML with Java: a guide to SAX, DOM, JDOM, JAXP, and TrAX. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-201-77186-1. xxxvii + 1071 pp. LCCN QA76.76.H94 H337 2003 Bar.

Haynes:2002:SCS

[Hay02] Mir Haynes. Second commentary on “XML and the new design regime”. ACM Journal of Computer Documentation, 26 (2):45–47, May 2002. CODEN AJCDBH. ISSN 1527-6805. See [Wil02b].

Halverson:2003:MMX

[HBG^{+}03] Alan Halverson, Josef Burger, Leonidas Galanis, Ameet Kini, Rajasekar Krishnamurthy, Ajith Nagaraja Rao, Feng Tian, Stratis Viglas, Yuan Wang, Jeffrey F. Naughton, and David J. De-Witt. Mixed mode XML query processing. In Freytag et al. [FLA^{+}03], pages 225–236. ISBN 0-12-722442-4.

URL http://www.vldb.org/
dblp/db/indices/a-tree/h/
Halverson:Alan.html.

Heires:2003:LEV

[HBH^{+}03] James T. Heires, Barry Boehm, Li Guo Huang, Paul C. Jorgensen, Laurie Williams, Alistair Cockburn, Rita Creel, Rob Schaaf, Steven Hong, William A. Arbaugh, and Mikhail S. Brikman. Letters: Earned-value management terminology; putting agile methods in perspective; standards clarification; balancing feedback and foresight; defending firewalls; analyzing XML development trends. Computer, 36(8):5–7, August 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2003/08/r8005.
htm;   http://www.computer.
org/computer/co2003/r8005.
pdf.

Hinkelman:2006:EPU

[HBZ06] S. Hinkelman, D. Buddenbaum, and L.-J. Zhang. Emerging patterns in the use of XML for information modeling in vertical industries. IBM Systems Journal, 45(2):373–378, ???? 2006. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/452/hinkelman.html.

Hunter:2000:BX

[HC^{+}00] David Hunter, Kurt Cagle, et al. Beginning XML. Wrox Press, Chicago, IL, USA, 2000. ISBN 1-86100-341-2. xix + 823 pp. LCCN QA76.76.H94 B44 2000. US$39.99.

Huang:2002:AMA

[HCC^{+}02] Yueh-Min Huang, Chin-Chiang Chang, Pin-Ying Chiu, Shin-Chang Huang, and William Cheng-Chung Chu. Applying multimedia authoring tool and XML techniques to standardized knowledge management for Web-based learning. Lecture Notes in Computer Science, 2436:386–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2436/24360386.

                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2436/24360386.
                                        pdf.

Heijl:2001:DXS

[Hei01]  Danny Heijl.  The Delphi

XML SAX2 component and MSXML 3.0. Dr. Dobbs Journal, 26(9):42, 46, 48, 50, 52, 54, September 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2001/2001_09/xmlsax2.
                                        txt;   http://www.ddj.com/
                                        ftp/2001/2001_09/xmlsax2.

zip. See correction [TEM^{+}01].

Heines:2003:EXS

[Hei03] J. M. Heines. Enabling XML storage from Java applets in a GUI programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):88– 93, 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Herzberg:2002:SX

[Her02] Amir Herzberg. Securing XML. Dr. Dobbs Journal, 27(3):56, 59–62, March 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2002/2002_03/
                                        secxml.txt.

Hosoya:2005:PPX

[HFC05] Haruo Hosoya, Alain Frisch, and Giuseppe Castagna. Parametric polymorphism for XML. ACM SIGPLAN Notices, 40(1):

50–62, January 2005. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hancock:2001:MXL

[HG01] B. Hancock and M. J. Giarlo.

Moving to XML: Latin texts XML conversion project at the Center for Electronic Texts in the Humanities. Library Hi Tech, 19(3):257–264, September 19, 2001. CODEN LI-HTD2. ISSN 0737-8831.

Hatano:2002:IRS

[HKYU02a] K. Hatano, H. Kinutani, M. Yoshikawa, and S. Uemura. Information retrieval system for XML documents. Lecture Notes in Computer Science, 2453:758–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-
ny.com/link/service/series/
0558/bibs/2453/24530758.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2453/24530758.
pdf.

Hatano:2002:EPX

[HKYU02b] Kenji Hatano, Hiroko Kinutani, Masatoshi Yoshikawa, and Shunsuke Uemura. Extraction of partial XML documents using IR-based structure and contents analysis. Lecture Notes in Computer Science, 2465:334– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/
                                        service/series/0558/bibs/
                                        2465/24650334.htm; http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2465/24650334.pdf.

Harding:2003:XRX

[HLM03] Philip J. Harding, Quanzhong Li, and Bongki Moon. XISS/R: XML indexing and storage system using RDBMS. In Freytag et al. [FLA^{+}03], pages 1073– 1076. ISBN 0-12-722442-4.

                                        URL http://www.vldb.org/
                                        dblp/db/indices/a-tree/h/
                                        Harding:Philip_J=.html.

Harold:2001:XND

[HM01] Elliotte Rusty Harold and W. Scott Means. XML in a nutshell: a desktop quick reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00058-8. xvi + 480 pp. LCCN QA76.76.H94 H36 2001.

Harold:2002:XN

[HM02] Elliotte Rusty Harold and W. Scott Means. XML in a nutshell. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00292-0. xvii + 613 pp. LCCN QA76.76.H94 H36 2002 Stacks.

Harold:2004:XND

[HM04] Elliotte Rusty Harold and W. Scott Means. XML in a nutshell: [a desktop quick reference]. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2004. ISBN 0-596-00764-7. xix + 689 pp. LCCN QA76.76.H94; Internet; QA76.76.H94 H36 2004eb; QA76.76.H94 H36 2004.

Holzner:2000:HBB

[Hol00a] Steven Holzner. HTML black book. Coriolis Group Books, Scottsdale, AZ, USA, 2000. ISBN 1-57610-617-9. xxxvi + 1195 pp. LCCN QA76.76.H94 H63 2000.

Holzschlag:2000:UH

[Hol00b] Molly E. Holzschlag. Using HTML 4. Que Corporation, Indianapolis, IN, USA, 6th, special ed edition, 2000. ISBN 0-7897-2267-4. various pp. LCCN QA76.76.H94 B76 2000.

Holzner:2001:IX

[Hol01a] Steven Holzner. Inside XML.

Inside series. New Riders Publishing, Carmel, IN, USA, 2001. ISBN 0-7357-1020-1. xxii + 1102 pp. LCCN QA76.76.H94 H64 2001.

Holzschlag:2001:SEU

[Hol01b] Molly E. Holzschlag. Special edition using XHTML. Que Corporation, Indianapolis, IN, USA, 2001. ISBN 0-7897-2431-6. xiv + 958 pp. LCCN QA76.76.H94 H656 2001.

Holzner:2002:IX

[Hol02]  Steven Holzner.  Inside

XSLT. New Riders Publishing, Carmel, IN, USA, 2002. ISBN 0-7357-1136-4 (paperback). xix + 616 pp. LCCN QA76.73.X58 H65 2002.

Hoque:2000:XRP

[Hoq00] Reaz Hoque. XML for Real Programmers. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2000. ISBN 0-12-355592-2. xv + 462 pp. LCCN QA76.76.H94 H66 2000. US$44.95.

Houlding:2001:VVD

[Hou01] David Houlding. VoiceXML and the voice-driven Internet. Dr. Dobbs Journal, 26(4):88, 90–92, 94–95, April 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_04/
                                        voicexml.zip;   http://www.
                                        ddj.com/ftp/2001/2001_04/
                                        voicexmltxt.zip.

Hosoya:2001:REP

[HP01] Haruo Hosoya and Benjamin Pierce. Regular expression pattern matching for XML. ACM SIGPLAN Notices, 36 (3):67–80, March 2001. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL

http://www.acm.org/pubs/
articles/proceedings/plan/
360204/p67-hosoya/p67-hosoya.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
plan/360204/p67-hosoya/.

Hosoya:2003:XST

[HP03] Haruo Hosoya and Benjamin C.

Pierce. XDuce: A statically typed XML processing language. ACM Transactions on Internet Technology (TOIT), 3 (2):117–148, May 2003. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Hendricks:2002:CMC

[HRB^{+}02] Brent Hendricks, Ross Reedstrom, Richard Baraniuk, Don Johnson, Bill Wilson, and Geneva Henry. Connexions: MathML and collaborative curriculum development in engineering. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Hastings:2005:XDS

[HRL^{+}05] Shannon Hastings, Matheus Ribeiro, Stephen Langella, Scott Oster, Umit Catalyurek, Tony Pan, Kun Huang, Renato Ferreira, Joel Saltz, and Tahsin Kurc. XML database support for distributed execution of data-intensive scientific workflows. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(3):50–55, September 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Huerter:2002:CFT

[HRW02] Sandy Huerter, Igor Rodionov, and Stephen Watt. Contentfaithful transformations for MathML. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Hjelm:2002:XUG

[HS02a] Johan Hjelm and Peter Stark. XSLT: the ultimate guide to transforming Web data. Professional developer’s guide series. Wiley, New York, NY, USA, 2002. ISBN 0-471-40603-1 (paperback). viii + 311 pp. LCCN QA76.73.X58 H54 2002.

Huang:2002:DXB

[HS02b] Shi-Ming Huang and Chun-Hao Su. The development of an XML-based data warehouse system. Lecture Notes in Computer Science, 2412:206– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2412/24120206.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2412/24120206.
                                        pdf.

Hongwei:2002:CPM

[HSJJ02a] Sun Hongwei, Zhang Shusheng, Zhou Jingtao, and Wang Jing.

Constraints-preserving mapping algorithm from XMLschema to relational schema. Lecture Notes in Computer Science, 2480:193–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer-ny.com/link/service/
series/0558/bibs/2480/24800193.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2480/24800193.
pdf.

Hongwei:2002:MXS

[HSJJ02b] Sun Hongwei, Zhang Shusheng, Zhou Jingtao, and Wang Jing. Mapping XML-schema to relational schema. Lecture Notes in Computer Science, 2510:322– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2510/25100322.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2510/25100322.pdf.

Huddleston:2008:HXC

[Hud08a]  Rob Huddleston.  HTML,

XHTML, and CSS: your visual blueprint for designing effective Web pages. Visual read less, learn more. Wiley, New York, NY, USA, 2008. ISBN 0-470-27436-0 (paperback). xiii + 369 pp. LCCN TK5105.888 .H828

2008. URL http://www.loc.
gov/catdir/enhancements/
fy0828/2008926447-d.html.

Huddleston:2008:XYV

[Hud08b] Rob Huddleston. XML: your visual blueprint for building expert Web sites with XML, CSS, XHTML, and XSLT. Visual read less, learn more. Wiley, New York, NY, USA, 2008. ISBN 0-471-93383-X (paperback). xii + 292 pp. LCCN QA76.76.H94 H933 2008.

Hunt:2002:WUA

[Hun02] Andrew Hunt. What’s the use? Analysis of real-world MathML and Web browser usage. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Hosoya:2000:RET

[HVP00] Haruo Hosoya, Jérôme Vouillon, and Benjamin C. Pierce. Regular expression types for XML. ACM SIGPLAN Notices, 35(9):11–22, September 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/proceedings/fp/
                                        351240/p11-hosoya/p11-hosoya.
                                        pdf;   http://www.acm.org/
                                        pubs/citations/proceedings/
                                        fp/351240/p11-hosoya/.

Huang:2004:STI

[HYC04] Bo Huang, Shanzhen Yi, and Weng Tat Chan. Spatiotemporal information integration in XML. Future Generation Computer Systems, 20(7): 1157–1170, October 1, 2004.

CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ives:2002:XQE

[IHW02] Zachary G. Ives, A. Y. Halevy, and D. S. Weld. An XML query engine for network-bound data. VLDB Journal: Very Large Data Bases, 11(4):380–402, December 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL http://link.springer.
de/link/service/journals/
00778/bibs/2011004/20110380.
htm;   http://link.springer.
de/link/service/journals/
00778/papers/2011004/20110380.
pdf.

ISO:2000:IIIb

[Int00] International Organization for Standardization. ISO/IEC 13673:2000: Information technology — Document processing and related communication — Conformance testing for Standard Generalized Markup Language (SGML) systems. International Organization for Standardization, Geneva, Switzerland, 2000. ISBN ???? 33 pp. LCCN ???? CHF

110. URL http://www.iso.
ch/cate/d22553.html.

Jackson:2003:EFO

[Jac03] Quinn Tyler Jackson. Efficient formalism-only parsing of XML/HTML using the §- calculus. ACM SIGPLAN Notices, 38(2):29–35, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jaksic:2004:MBS

[Jak04] Mirjana Jakšić. Mapping of bibliographical standards into XML. Software—Practice and Experience, 34(11):1051–1064, September 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jagadish:2002:TNX

[JAKC^{+}02] H. V. Jagadish, S. Al-Khalifa, A. Chapman, L. V. S. Lakshmanan, A. Nierman, S. Paparizos, J. M. Patel, D. Srivastava, N. Wiwatwattana, Y. Wu, and C. Yu. TIM-BER: A native XML database. VLDB Journal: Very Large Data Bases, 11(4):274–291, December 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00778/bibs/2011004/20110274.
                                        htm;   http://link.springer.
                                        de/link/service/journals/
                                        00778/papers/2011004/20110274.
                                        pdf.

Janulaitis:2001:CHU

[Jan01] Michael Janulaitis. Creating HTML user interfaces for server programs. Dr. Dobbs Journal, 26(6):101–102, 104, 106, 108–109, June 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_06/

htmlui.txt;   http://www.
ddj.com/ftp/2001/2001_06/
htmlui.zip.

Jones:2002:PX

[JD02] Christopher A. Jones and Fred L. Drake. Python and XML. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00128-2. xvi + 360 pp. LCCN QA76.73.P98 J66 2002.

Josifovski:2005:QXS

[JFB05] Vanja Josifovski, Marcus Fontoura, and Attila Barta. Querying XML streams. VLDB Journal: Very Large Data Bases, 14 (2):197–210, April 2005. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Jamsa:2002:HWD

[JKA02] Kris A. Jamsa, Konrad King, and Andy Anderson. HTML and Web design: tips and techniques. Osborne/McGraw-Hill, Berkeley, CA, USA, 2002. ISBN 0-07-219394-8 (paperback). xx + 665 pp. LCCN QA76.76.H94 J355 2002.

Jin:2004:CSD

[JLP04] Lee Hwa Jin, Ho-Joon Lee, and Jong C. Park. Constructing SSML documents with automatically generated intonation information in a combinatory categorial grammar framework. International Journal of Computer Processing of Oriental Languages (IJCPOL), 17(4): 223–??, December 2004. CO-DEN ???? ISSN 0219-4279.

Jagadish:2002:TTA

[JLST02] H. V. Jagadish, Laks V. S. Lakshmanan, Divesh Srivastava, and Keith Thompson. TAX: A tree algebra for XML. Lecture Notes in Computer Science, 2397:149–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2397/23970149.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2397/23970149.pdf.

Johnston:2002:CME

[Joh02] Clifford Johnston. Collecting mathematical expressions with Web forms: Converting plain text to MathML. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Jung:2002:EIX

[JOKA02] Jong-Seok Jung, Dong-Ik Oh, Yong-Hae Kong, and Jong-Keun Ahn. Extracting information from XML documents by reverse generating a DTD. Lecture Notes in Computer Science, 2510:314– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/

service/series/0558/bibs/
2510/25100314.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2510/25100314.pdf.

Jones:2003:ELX

[Jon03a] M. Tim Jones. An embeddable lightweight XML-RPC server. Dr. Dobbs Journal, 28(6):60, 62–64, 66–67, June 2003. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/documents/s=8213/
ddj0306g/.

Jonker:2003:XSD

[Jon03b] W. Jonker. XML and secure data management in an ambient world. International Journal of Computer Systems Science and Engineering, 18(5): ??, September 2003. CODEN CSSEEI. ISSN 0267-6192.

Jeong:2004:JBS

[JSSM04] J. Jeong, D. Shin, D. Shin, and K. Moon. Java-based single sign-on library supporting SAML (Security Assertion Markup Language) for distributed Web services. Lecture Notes in Computer Science, 3007:891–894, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jiang:2003:HTJ

[JWLY03] Haifeng Jiang, Wei Wang, Hongjun Lu, and Jeffrey Xu Yu. Holistic twig joins on indexed XML documents. In Freytag et al. [FLA^{+}03], pages 273–284. ISBN 0-12-722442-4.

                                        URL http://www.vldb.org/
                                        dblp/db/indices/a-tree/j/
                                        Jiang:Haifeng.html.

Kay:2000:X

[Kay00] Michael Kay. XSLT. Wrox Press, Chicago, IL, USA, 2000. ISBN 1-86100-312-9. 600 (est.) pp. LCCN ???? US$34.99.

Kay:2001:XPR

[Kay01] Michael Kay. XSLT programmer’s reference. Programmer to programmer. Wrox Press, Chicago, IL, USA, second edition, 2001. ISBN 1-86100-506-

7. xxxiii + 939 pp. LCCN

QA76.73.X58 K38 2001.

Kay:2008:XXP

[Kay08] Michael Kay. XSLT 2.0 and XPath 2.0: programmer’s reference. Wrox programmer to programmer. Wiley, New York, NY, USA, fourth edition, 2008. ISBN 0-470-19274-7. xlix + 1316 pp. LCCN QA76.73.X58.

                                        URL http://www.loc.gov/
                                        catdir/enhancements/fy0814/
                                        2007049362-d.html;   http:
                                        //www.loc.gov/catdir/enhancements/
                                        fy0814/2007049362-t.html.

Kudrass:2002:MXD

[KC02] Thomas Kudrass and Matthias Conrad. Management of XML documents in object-relational databases. Lecture Notes in Computer Science, 2490:210– ??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2490/24900210.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2490/24900210.pdf.

Kim:2001:TSJ

[Kim01] Eugene Eric Kim. A triumph of simplicity: James Clark on markup languages and XML. Dr. Dobbs Journal, 26(7):56, 58–60, July 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/.

Kitiyakara:2002:ATH

[Kit02]  Narti Kitiyakara.  Accep-

tance testing HTML. Lecture Notes in Computer Science, 2418:112–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.de/link/service/
series/0558/bibs/2418/24180112.
htm; http://link.springer.
de/link/service/series/0558/
papers/2418/24180112.pdf.

Kim:2002:LSE

[KiY02] Sungrim Kim and Yong ik Yoon.

A levelized schema extraction for XML document using user-defined graphs. Lecture Notes in Computer Science, 2510:434–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/

                                        service/series/0558/bibs/
                                        2510/25100434.htm; http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2510/25100434.pdf.

Kerer:2002:XPG

[KKK02] Clemens Kerer, Engin Kirda, and Christopher Krügel. XGuide — A practical guide to XML-based Web engineering. Lecture Notes in Computer Science, 2376:104–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/
                                        service/series/0558/bibs/
                                        2376/23760104.htm; http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2376/23760104.pdf.

Kempa:2002:XBA

[KL02] Martin Kempa and Volker Linnemann. XML-based applications using XML schema. Lecture Notes in Computer Science, 2490:67–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/
                                        service/series/0558/bibs/
                                        2490/24900067.htm; http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2490/24900067.pdf.

Klein:2003:BEH

[Kle03] J. M. E. Klein. Building Enhanced HTML Help with DHTML & CSS [Book Review].

IEEE Transactions on Professional Communication, 46(1): 68–69, March 2003. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kim:2002:IPV

[KLL02] Sang-Kyun Kim, Myungcheol Lee, and Kyu-Chul Lee. Immediate and partial validation mechanism for the conflict resolution of update operations in XML databases. Lecture Notes in Computer Science, 2419:387– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2419/24190387.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2419/24190387.
pdf.

Kanne:2006:ISC

[KM06] C. C. Kanne and G. Moerkotte. The importance of sibling clustering for efficient bulkload of XML document trees. IBM Systems Journal, 45(2): 321–334, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/452/kanne.html.

Knuckles:2001:IIP

[Knu01] Craig D. Knuckles. Introduction to interactive programming on the Internet: using HTML and JavaScript. Wiley, New York, NY, USA, 2001. ISBN 0-471-38366-X (paperback). xviii + 423 pp. LCCN QA76.625 .K58 2001.

Koch:2003:EPE

[Koc03]  Christoph Koch.  Efficient

processing of expressive nodeselecting queries on XML data in secondary storage: A tree automata-based approach. In Freytag et al. [FLA^{+}03], pages 249–260. ISBN 0-12-722442-4.

                                        URL http://www.vldb.org/
                                        dblp/db/indices/a-tree/k/
                                        Koch:Christoph.html.

Kim:2002:SIH

[KP02a] Hyon Hee Kim and Seung Soo Park. Semantic integration of heterogeneous XML data sources. Lecture Notes in Computer Science, 2425:95– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2425/24250095.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2425/24250095.
                                        pdf.

Kim:2002:FMR

[KP02b] Jaehoon Kim and Seog Park. Flexible modification of relational schema by X2RMap in storing XML into relations. Lecture Notes in Computer Science, 2510:330–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2510/25100330.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2510/25100330.pdf.

Koloniari:2004:FXB

[KP04] Georgia Koloniari and Evaggelia Pitoura. Filters for XMLbased service discovery in pervasive computing. The Computer Journal, 47(4):461– 474, July 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_04/470461.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_04/pdf/
470461.pdf.

Koloniari:2005:PPM

[KP05] Georgia Koloniari and Evaggelia Pitoura. Peer-to-peer management of XML data: issues and research challenges. SIG-MOD Record (ACM Special Interest Group on Management of Data), 34(2):6–17, June 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kratky:2002:GFE

[KPSS02] Michal Krátký, Jaroslav Pokorný, Tomáš Skopal, and Václav Snášel. The geometric framework for exact and similarity querying XML data.

Lecture Notes in Computer Science, 2510:35–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2510/25100035.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2510/25100035.pdf.

Kroeker:2000:PCL

[Kro00a]  Kirk L. Kroeker.  Prod-

ucts: Controlling the look and feel of Windows; sophisticated project estimation for $99; Web applications for the Internet-challenged; networkbased digital asset manager; Linux middleware framework; XML-based information portal; Java development for small computing devices; build installations for Windows 2000. Computer, 33(3):105– 107, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.
org/co/books/co2000/pdf/
r3105.pdf.

Kroeker:2000:PDD

[Kro00b] Kirk L. Kroeker. Products: Digital dashboard concept now affordable; defect-tracking tool adds custom workflow; active application source code browser for MFC; customizable set of dials and gauges for Motif; fist IDE for PlayStation2; new XML editor hits channels; the disk optimization controversy. Computer, 33(2):120– 122, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.
org/co/books/co2000/pdf/
r2120.pdf.

Klettke:2002:MXD

[KSH02] Meike Klettke, Lars Schneider, and Andreas Heuer. Metrics for XML document collections. Lecture Notes in Computer Science, 2490:15– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2490/24900015.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2490/24900015.pdf.

Kim:2002:DMA

[KSK^{+}02] Sung Wan Kim, Pan Seop Shin, Youn Hee Kim, Jaeho Lee, and Hae Chull Lim. A data model and algebra for document-centric XML document. Lecture Notes in Computer Science, 2344:714– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2344/23440714.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2344/23440714.pdf.

Kohlhase:2002:ACM

[KSSS02] Michael Kohlhase, Matthew Szudzik, Dana Scott, and Klaus Sutner. Acquisition of content: MathML in an academic setting. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Kunkle:2002:WBI

[Kun02] Wanda Kunkle. A Web-based integral evaluator: A demonstration of the successful integration of WebEQ, Maple, and Java. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Kuzniarek:2002:PSD

[Kuz02] Andre Kuzniarek. Publicon 1.0: Structured documentauthoring system demo. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Kotok:2002:ENG

[KW02] Alan Kotok and David R. R. Webber. ebXML: the new global standard for doing business over the Internet. New Riders Publishing, Carmel, IN, USA, 2002. ISBN 0-7357-1117-8. xxxix + 339 pp. LCCN QA76.76.H94 K686 2002.

Kha:2002:ARP

[KYU02a] D. D. Kha, M. Yoshikawa, and S. Uemura. Application of rUID in processing XML queries on structure and keyword. Lecture Notes in Computer Science, 2453:279– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2453/24530279.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2453/24530279.
pdf.

Kha:2002:SNS

[KYU02b] Dao Dinh Kha, Masatoshi Yoshikawa, and Shunsuke Uemura. A structural numbering scheme for XML data. Lecture Notes in Computer Science, 2490:91–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2490/24900091.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2490/24900091.pdf.

Ladd:2001:PEU

[Lad01] Eric Ladd. Platinum edition using XHTML, XML and Java 2. Que Corporation, Indianapolis, IN, USA, platinum edition, 2001. ISBN 0-7897-2473-1. xxiv + 1413 pp. LCCN QA76.76.H94 P585 2001.

Ludascher:2002:TLT

[LAG02] Bertram Ludäscher, Ilkay Altintas, and Amarnath Gupta. Time to leave the trees: From syntactic to conceptual querying of XML. Lecture Notes in Computer Science, 2490:148– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/
                                        service/series/0558/bibs/
                                        2490/24900148.htm; http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2490/24900148.pdf.

Larson:2003:VID

[Lar03] James A. Larson. VoiceXML: introduction to developing speech applications. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2003. ISBN 0-13-009262-2. xii + 324 pp. LCCN QA76.76.H94 L35 2003.

Latteier:2002:ZPT

[Lat02] Amos Latteier. Zope page templates. Dr. Dobbs Journal, 27 (2):67, 69–70, 73–75, February 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/2002/2002_02/
                                        zope.zip.

Lawrence:2002:NXD

[Law02] Ramon Lawrence. Naming in XML documents. Lecture Notes in Computer Science, 2519:1287–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/
                                        service/series/0558/bibs/
                                        2519/25191287.htm; http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2519/25191287.pdf.

Lawson:2004:VSW

[Law04]  K. Lawson.  Visualizing

the Semantic Web: XMLbased Internet and Information Visualization [book review]. IEEE Distributed Systems Online, 5(2):6/1–6/3, February 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/28454/01270786.pdf?
isnumber=28454&prod=JNL&
arnumber=1270786&arSt=+6%
2F1&ared=+6%2F3&arAuthor=
+Lawson%2C+K.;   http://
ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=28454&
arnumber=1270786&count=6&
index=4.

Lee:2003:AXT

[LB03] B.-S. Lee and B. R. Bryant.

Applying XML technology for implementation of natural language specifications. International Journal of Computer Systems Science and Engineering, 18(5):??, September 2003. CODEN CSSEEI. ISSN 0267-6192.

Li:2003:CXT

[LBN03] Chengkai Li, Philip Bohannon, and P. P. S. Narayan. Composing XSL transformations with XML publishing views. In ACM [ACM03a], pages 515– 526. ISBN ???? LCCN ????

Lee:2000:TXD

[LC00] Dongwon Lee and Wesley W.

Chu. Transforming XML document type definition to relational schema while preserving semantic constraints. Los Angeles, CA, 2000. 23 pp.

Lu:2004:XMS

[LC04] Eric Jui-Lin Lu and Rai-Fu Chen. An XML multisignature scheme. Applied Mathematics and Computation, 149(1): 1–14, February 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lee:2002:DRE

[LCC^{+}02] Kyong-Ho Lee, Yoon-Chul Choy, Sung-Bae Cho, Xiao Tang, and Victor McCrary. Document reverse engineering: From paper to XML. Lecture Notes in Computer Science, 2423:503–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-
ny.com/link/service/series/
0558/bibs/2423/24230503.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2423/24230503.
pdf.

Lemay:2001:STY

[LCT01] Laura Lemay, Rafe Colburn, and Denise Tyler. Sams teach yourself Web publishing with HTML and XHTML in 21 days. SAMS Publishing, Indianapolis, IN, USA, third edition, 2001. ISBN 0-672-32077-0. 834 pp. LCCN QA76.76.H94 L453 2001.

Li:2004:WAS

[LCZ04] H. Li, J. Cai, and R. Zhu. Wireless applications service platform based on XML and Java. Journal- Zhejiang University of Technology, 32(2):182–185, 2004. CODEN ???? ISSN 1006-4303.

Le:2000:DAF

[Le00]  Chuong Cam Le.  A dy-

namic approach to formulation of customized HTMLbased XMAS template queries. Thesis (M.S.), University of California, San Diego, San Diego, CA, USA, 2000.

Lear:2000:NBT

[Lea00] Anne C. Lear. News briefs:

Transmeta unveils Crusoe mobile processor; new Web protocol combines HTML and XML; as UML evolves, its applications expand. Computer, 33(3): 24–26, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.
org/co/books/co2000/pdf/
r3024.pdf.

Lehnert:2002:WWG

[Leh02] Wendy G. Lehnert. The Web wizard’s guide to HTML. Addison Wesley’s Web wizard series. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-74172-5. x + 189 pp. LCCN QA76.76.H94 L43 2002.

Lenz:2005:XPR

[Len05] Evan Lenz. XSLT 1.0 pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10008-6 (paperback). 104 (est.) pp. LCCN ????

Lewis:2002:EXR

[Lew02] Bryn Lewis. Extraction of XML from relational databases. Lecture Notes in Computer Science, 2490:228–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2490/24900228.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2490/24900228.pdf.

Larson:2001:XDM

[LFG^{+}01] Per-Åke Larson, Dana Florescu, Goetz Graefe, Guido Moerkotte, Hamid Pirahesh, and Harald Schöning. XML data management (panel session): go native or spruce up relational systems? In Sellis and Mehrotra [SM01], page 620. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Lloyd:2008:UHR

[LHO08] Ian Lloyd, Lachlan Hunt, and Tommy Olsson. The ultimate HTML reference. Sitepoint, Collingwood, Victoria, Australia, 2008. ISBN 0-9802858-8-7. x + 537 pp. LCCN QA76.76.H94 L56 2008.

Lin:2003:DEA

[Lin03] W. Lin. Development of electronic acquisition model for project scheduling (e-AMPS) using Java-XML. NIST Special Publication SP, 989:41–46, 2003. CODEN ???? ISSN 1048-776X.

Liu:2000:QPX

[Liu00] Xiaoyu Liu. Querying and presenting XML data for MIX system. Vita thesis (ph.d.), Computer Science Department, University of California, San Diego, San Diego, CA, USA, 2000.

Liu:2002:CSH

[Liu02]  M. Liu.  Capturing se-

mantics in HTML documents. Lecture Notes in Computer Science, 2453:103– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2453/24530103.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2453/24530103.
pdf.

Lux:2002:XMI

[LKB^{+}02] Mathias Lux, Werner Klieber, Jutta Becker, Klaus Tochtermann, Harald Mayer, Helmut Neuschmied, and Werner Haas. XML and MPEG-7 for interactive annotation and retrieval using semantic metadata. J.UCS: Journal of Universal Computer Science, 8 (10):965–??, October 28, 2002. CODEN ???? ISSN 0948-6968.

URL http://www.jucs.org/
jucs_8_10/xml_and_mpeg_7.

Lee:2002:IRX

[LL02a] Sung Wan Kim Jaeho Lee and Hae Chull Lim. Indexing and retrieval of XML-encoded structured documents in dynamic environment. Lecture Notes in Computer Science, 2480:141–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-
ny.com/link/service/series/
0558/bibs/2480/24800141.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2480/24800141.
pdf.

Liu:2002:DWE

[LL02b] Jixue Liu and Chengfei Liu. A declarative way of extracting XML data in XSL. Lecture Notes in Computer Science, 2435:374–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-
ny.com/link/service/series/

0558/bibs/2435/24350374.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2435/24350374.
pdf.

Lujan-Mora:2002:WOA

[LMMT02] Sergio Luján-Mora, Enrique Medina, and Juan Trujillo. A Web-oriented approach to manage multidimensional models through XML schemas and XSLT. Lecture Notes in Computer Science, 2490:29– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2490/24900029.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2490/24900029.pdf.

Ludascher:2002:TBX

[LMP02] Bertram Ludäscher, Pratik Mukhopadhyay, and Yannis Papakonstantinou. A transducer-based XML query processor. In Bernstein et al. [B^{+}02], pages 227–238. ISBN 1-55860-869-9. LCCN ???? URL

http://www.vldb.org/conf/
2002/S07P03.pdf.

Lopez:2002:XBD

[LMY02] Javier López, Antonio Maña, and Mariemma I. Yagüe. XMLbased distributed access control system. Lecture Notes in Computer Science, 2455:203– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2455/24550203.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2455/24550203.
                                        pdf.

Lamiroy:2002:SXU

[LN02] Bart Lamiroy and Laurent Najman. Scan-to-XML: Using software component algebra for intelligent document generation. Lecture Notes in Computer Science, 2390:211– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/
                                        service/series/0558/bibs/
                                        2390/23900211.htm; http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2390/23900211.pdf.

Leobacher:2002:MAI

[LP02] G. Leobacher and F. Pillichshammer. A method for approximate inversion of the hyperbolic CDF. Computing, 69(4):291–303, December 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lau:2002:DXW

[LR02] C. Lau and A. Ryman. Developing XML Web services with WebSphere Studio Application Developer. IBM Systems Journal, 41(2):178–??, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/412/lau.html.

Lee:2003:ERM

[LS03] Jae Kyu Lee and Mye M. Sohn.

The eXtensible Rule Markup Language. Communications of the ACM, 46(5):59–64, May 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lemberg:2001:CPC

[LSS01] Ingrid Lemberg, Bernhard Schroder, and Angelika Storrer. Chancen und Perspektiven computergestutzter Lexikographie: Hypertext, Internet und SGML/XML für die Produktion und Publikation digitaler Worterbucher. (German) [Opportunities and perspectives of computer-aided lexicography: Hypertext, Internet and SGML/XML for the production and publication of digital dictionaries], volume 107 of Lexicographica. Series maior ISSN: 0175-9264; 107. Niemeyer, Tübingen, Germany, 2001. ISBN 3-484-39107-3 (paperback). ISSN 0175-9264. vi + 270 pp. LCCN P327.5.T43 C43 2001; P327.5.T43 C436 2001. In German; abstracts in English and French.

Lucas:2000:VWB

[Luc00] Bruce Lucas. VoiceXML for Web-based distributed conversational applications. Communications of the ACM, 43(9):

53–57, September 2000. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-9/p53-lucas/.

Lutz:2002:BXC

[Lut02] Michael J. Lutz. Bookshelf: The XML culture clash [System Architecture with XML]; broadband’s core technologies [Integrated Broadband Networks: TCP/IP, ATM, SDH/SONET, and WDM/ Optics]; efficient IT management [Building Operational Excellence]; making CORBA secure [Developing Secure Distributed Systems with CORBA]. Computer, 35(10): 72, October 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/co/2002/10/rx072.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/co/2002/10/rx072.
                                        pdf.

Lu:2004:RXE

[LW04] Eric Jui-Lin Lu and Chang-Chuan Wu. A ReScUE XML/ EDI model. Software—Practice and Experience, 34(3):315– 338, March 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lim:2002:XLS

[LWP^{+}02] Lipyeow Lim, Min Wang, Sriram Padmanabhan, Jeffrey Scott Vitter, and Ronald Parr. XPathLearner: An online self-tuning Markov histogram for XML path selectivity estimation. In Bernstein et al. [B^{+}02], pages 442–453. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.
org/conf/2002/S13P01.pdf.

Lv:2002:PED

[LWY^{+}02] Jianhua Lv, Guoren Wang, Jeffrey X. Yu, Ge Yu, Hongjun Lu, and Bing Sun. Performance evaluation of a DOMbased XML database: Storage, indexing and query optimization. Lecture Notes in Computer Science, 2419:13– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2419/24190013.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2419/24190013.
pdf.

Liu:2005:VKC

[LYT^{+}05] Yunfeng Liu, Dongqing Yang, Shiwei Tang, Tengjiao Wang, and Jun Gao. Validating key constraints over XML document using XPath and structure checking. Future Generation Computer Systems, 21 (4):583–595, April 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Liu:2003:RDE

[LZZ03] H. Liu, X. Zhou, and B. Zhong. The realization of data exchange between different system platforms using Java and XML under the environment of Lotus Domino. Journal — Sichuan Normal University Natural Science Edition, 26(3): 327–330, 2003. CODEN ???? ISSN 1001-8395.

Milo:2003:EIX

[MAA^{+}03] Tova Milo, Serge Abiteboul, Bernd Amann, Omar Benjelloun, and Fred Dang Ngoc. Exchanging intensional XML data. In ACM [ACM03a], pages 289–300. ISBN ???? LCCN ????

Milo:2005:EIX

[MAA^{+}05] Tova Milo, Serge Abiteboul, Bernd Amann, Omar Benjelloun, and Fred Dang Ngoc. Exchanging intensional XML data. ACM Transactions on Database Systems, 30(1):1– 40, March 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Min:2003:EES

[MAC03] Jun-Ki Min, Jae-Yong Ahn, and Chin-Wan Chung. Efficient extraction of schemas for XML documents. Information Processing Letters, 85(1):7–12, January 16, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mamlin:2001:OSX

[Mam01] B. Mamlin. An open-source XML-based Java implementation of the medical gopher order and note writing tool. Journal of Biomedical Informatics, 35 (SUPP):826–??, 2001. CODEN JBIOBL. ISSN 1532-0464.

Mangano:2002:XC

[Man02a] Sal Mangano. XSLT cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00372-2. xiv + 654 pp. LCCN QA76.73.X58 M36 2003.

Mansfield:2002:EIM

[Man02b] Paul Mansfield. Essential ingredients for mathematical content deployment on the Web: The MapleNet experience. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Marchal:2000:XE

[Mar00] Benoît Marchal. XML by example. Que Corporation, Indianapolis, IN, USA, 2000. ISBN 0-7897-2242-9. xv + 505 pp. LCCN QA76.76.H94 M2545 2000. US$24.99.

Martinsson:2001:SXW

[Mar01] Tobias Martinsson. Scripting XML and WMI for Microsoft SQL Server 2000: professional developer’s guide. Professional developer’s guide series. Wiley, New York, NY, USA, 2001. ISBN 0-471-39951-5 (paperback). xviii + 413 pp. LCCN QA76.76.H94 M279 2001.

Markus:2004:LXF

[Mar04] Arjen Markus. Have a look at XML files. ACM Fortran Forum, 23(3):2–10, December 2004. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Markus:2005:SLX

[Mar05] Arjen Markus. Have a second look at XML files. ACM Fortran Forum, 24(1):2–5, April 2005. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Mastidoro:2002:IDL

[Mas02] Nicola Mastidoro. The IntraText digital library: XMLdriven online library based on high accessibility, lexical hypertextualization and scholarly accuracy in philological/textual notations. Lecture Notes in Computer Science, 2458:659– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2458/24580659.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2458/24580659.pdf.

May:2002:RBQ

[May02]  Wolfgang May.  A rule-

based querying and updating language for XML. Lecture Notes in Computer Science, 2397:165–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.de/link/service/
series/0558/bibs/2397/23970165.
htm; http://link.springer.
de/link/service/series/0558/
papers/2397/23970165.pdf.

Murthy:2003:XSO

[MB03] Ravi Murthy and Sandeepan Banerjee. XML schemas in Oracle XML DB. In Freytag et al. [FLA^{+}03], pages 1009– 1018. ISBN 0-12-722442-4.

URL http://www.vldb.org/
dblp/db/indices/a-tree/m/
Murthy:Ravi.html.

Melton:2006:QXX

[MB06] Jim Melton and Stephen Buxton. Querying XML: XQuery, XPath, and SQL/XML in context. The Morgan Kaufmann series in data management systems. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2006. ISBN 1-55860-711-0 (paperback). xxvii + 815 pp. LCCN QA76.76.H94 M448

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0668/2006298085-d.html.

McFedries:2000:CIG

[McF00] Paul McFedries. The complete idiot’s guide to creating a Web page. Que Corporation, Indianapolis, IN, USA, fourth edition, 2000. ISBN 0-7897-2256-9. xxx + 414 pp. LCCN QA76.76.H94 M387 2000.

McGrath:2000:XPP

[McG00] Sean McGrath. XML processing with Python. The Charles F. Goldfarb series on open information management. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-021119-2. xxiv + 527 pp. LCCN QA76.76.H94 M3885 2000. US$44.99. URL http:

//www.phptr.com/ptrbooks/
ptr_0130211192.html.  In-

cludes CD-ROM.

McLaughlin:2000:JX

[McL00] Brett McLaughlin. Java and XML. Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 0-596-00016-2. xvi + 479 pp. LCCN QA76.73.J38 M39 2000.

McLaughlin:2001:JX

[McL01] Brett McLaughlin. Java and XML. Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2001. ISBN 0-596-00197-5 (paperback). xv + 509 pp.

LCCN QA76.73.J38 M39 2001. Also available via the Internet.

McLaughlin:2002:JXD

[McL02] Brett McLaughlin. Java and XML data binding. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00278-5. xi + 200 pp. LCCN QA76.73.J38 M39 2002 Stacks.

McLaughlin:2007:JX

[McL07] Brett McLaughlin. Java and XML. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2007. ISBN 0-596-10149-X (paperback). 465 pp. LCCN QA76.73.J38 M39

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0715/2007273941-d.html;
 http://www.loc.gov/catdir/
toc/fy0710/2007273941.html.

Martinez:2003:XMS

[MdlFD03] M. M. Martínez, P. de la Fuente, and J.-C. Derniame. XML as a means to support information extraction from legal documents. International Journal of Computer Systems Science and Engineering, 18(5): ??, September 2003. CODEN CSSEEI. ISSN 0267-6192.

Melton:2001:SMA

[ME01] Jim Melton and Andrew Eisenberg. SQL multimedia and application packages (SQL/MM). SIGMOD, 30(4):97–102, December 2001.

Malerba:2002:MHP

[MEC02] Donato Malerba, Floriana Esposito, and Michelangelo Ceci. Mining HTML pages to support document sharing in a cooperative system. Lecture Notes in Computer Science, 2490:420–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/
                                        service/series/0558/bibs/
                                        2490/24900420.htm; http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2490/24900420.pdf.

Mercer:2001:XBG

[Mer01] Dave Mercer. XML: a beginner’s guide. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-212740-6. xxii + 400 pp. LCCN QA76.76.H94 M46 2001.

Mercer:2002:SOH

[Mer02] Dave Mercer. Schaum’s outline of HTML. Schaum’s outline series. McGraw-Hill, New York, NY, USA, 2002. ISBN 0-07-137365-9 (paperback). viii + 232 pp. LCCN QA76.76.H94 M45 2002.

Meyer:2000:CSS

[Mey00] Eric A. Meyer. Cascading Style Sheets: the definitive guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-622-6. xiv + 453 pp. LCCN QA76.76.H94 M47 2000.

Meyer:2001:CSS

[Mey01a] Eric A. Meyer. Cascading Style Sheets 2.0: programmer’s reference. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-213178-0. xvi + 334 pp. LCCN QA76.76.H94 M47 2001.

Meyer:2001:CPR

[Mey01b] Eric A. Meyer. CSS pocket reference. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00120-7. 91 pp. LCCN QA76.76.H94 M472 2001.

Miner:2002:OMD

[Min02a] Robert Miner. An object model for dynamic math. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Miner:2002:TWA

[Min02b] Robert Miner. Two ways to author for MathPlayer with WebEQ. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Musciano:2000:HXD

[MK00] Chuck Musciano and Bill Kennedy. HTML and XHTML, the definitive guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2000. ISBN 0-596-00026-X. xviii + 660 pp. LCCN QA76.76.H94 M875 2000.

Musciano:2002:HXD

[MK02] Chuck Musciano and Bill Kennedy. HTML & XHTML: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fifth edition, 2002. ISBN 0-596-00382-X. xxi + 645 pp. LCCN QA76.76.H94 M875 2002 Stacks. US$39.95. URL

http://safari.oreilly.com/
059600382X;   http://www.
oreilly.com/catalog/html5.

Martin:2001:LXC

[MKR^{+}01] Larry Martin, Eugene Kim, Toby Reyelts, Al Stevens, Rob Chamberlin, Michael Brinkley, Michael Wojcik, and Jeff Duntemann. Letters: XML and the 21st Century; passing the C++ test; Ruby, Ruby; PHP and hosted applications; shared libraries aren’t all bad. Dr. Dobbs Journal, 26(3):10, 12, March 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/.

Morgenthal:2001:EAI

[MlF01] J. P. (Jeffrey P.) Morgenthal and Bill la Forge. Enterprise Application Integration with XML and Java. The definitive XML series from Charles F. Goldfarb. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-085135-3 (paperback/CD-ROM). xix + 504 pp. LCCN QA76.76.H94 M6 2001. US$44.99. URL http:/

/www.phptr.com/ptrbooks/
ptr_0130851353.html.  In-

cludes CD-ROM.

Meng:2003:OSB

[MLLA03] Xiaofeng Meng, Daofeng Luo, Mong-Li Lee, and Jing An. OrientStore: A schema based native XML storage system. In Freytag et al. [FLA^{+}03], pages 1057–1060. ISBN 0-12-722442-4. URL

http://www.vldb.org/dblp/
db/indices/a-tree/m/Meng:
Xiaofeng.html.

Medina:2002:HCM

[MLMT02] Enrique Medina, Sergio Luján-Mora, and Juan Trujillo. Handling conceptual multidimensional models using XML through DTDs. Lecture Notes in Computer Science, 2405:66– ??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2405/24050066.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2405/24050066.
                                        pdf.

Melton:2001:SME

[MMJ^{+}01] Jim Melton, Jan-Eike Michels, Vanja Josifovski, Krishna Kulkarni, Peter Schwarz, and Kathy Zeidenstein. SQL and management of external data. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(1):70–77, March 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Melton:2002:SMS

[MMJ^{+}02] Jim Melton, Jan Eike Michels, Vanja Josifovski, Krishna Kulkarni, and Peter Schwarz. SQL/MED: a status report. SIGMOD, 31(3):81–89, September 2002.

Martens:2005:CTT

[MN05] Wim Martens and Frank Neven. On the complexity of typechecking top-down XML transformations. Theoretical Computer Science, 336(1):153– 180, May 25, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Moreno:2000:HDT

[Mor00] Carlos Moreno. HTML document templates for CGI applications. C/C++ Users Journal, 18(9):10–??, September 2000. CODEN CCUJEX. ISSN 1075-2838.

Min:2003:XQC

[MPC03] Jun-Ki Min, Myung-Jae Park, and Chin-Wan Chung. XPRESS: a queriable compression for XML data. In ACM [ACM03a], pages 122–133. ISBN ???? LCCN ????

Murray-Rust:2001:CMX

[MRR01] Peter Murray-Rust and Henry S.

Rzepa. Chemical markup, XML and the World-Wide Web. 2. Information objects and the CMLDOM. Journal of Chemical Information and Computer Sciences, 41(??): 1113–1123, ???? 2001. CODEN JCISD8. ISSN 0095-2338.

Murray-Rust:2003:CMX

[MRR03] Peter Murray-Rust and Henry S.

Rzepa. Chemical markup, XML and the Worldwide Web. Part 4. CML schema. Journal of Chemical Information and Computer Sciences, 43(3):757– 772, ???? 2003.

Murray-Rust:2004:CMX

[MRRWW04] Peter Murray-Rust, Henry S.

Rzepa, Mark J. Williamson, and Egon L. Willighagen. Chemical markup, XML, and the World Wide Web. 5. Applications of chemical metadata in RSS aggregators. Journal of Chemical Information and Computer Sciences, 44(??):

462–469, ???? 2004. CODEN JCISD8. ISSN 0095-2338.

Marian:2003:PXD

[MS03] Amélie Marian and Jérôme Siméon. Projecting XML documents. In Freytag et al. [FLA^{+}03], pages 213– 224. ISBN 0-12-722442-4.

                                        URL http://www.vldb.org/
                                        dblp/db/indices/a-tree/m/
                                        Marian:Am=eacute=lie.html.

Muench:2000:BOX

[Mue00] Steve Muench. Building Oracle XML applications. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-691-9. xvi + 792 pp. LCCN QA76.76.H94 M75 2000.

Muffke:2001:CPE

[Müf01] Friedger Müffke. The Curl programming environment: Making the Internet interactive. Dr. Dobbs Journal, 26(9):66, 68, 70–71, September 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_09/
                                        curllist.zip.

Munroe:2000:BVI

[Mun00] Kevin D. Munroe. BBQ: a visual interface and API for integrated browsing and querying of XML. Thesis (m.s.), Computer Science Department, University of California, San Diego, San Diego, CA, USA, 2000.

Murphy:2000:TBX

[Mur00] A. J. Murphy. Three books on XML [book reviews]. IEEE Transactions on Professional Communication, 43(1):108– 109, March 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Musayev:2001:SSA

[Mus01] Eldar A. Musayev. SAX2:

The Simple API for XML. Dr. Dobbs Journal, 26(2):130, 132–133, February 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_02/
sax2.txt;   http://www.ddj.
com/ftp/2001/2001_02/sax2.
zip.

Myllymaki:2002:EWD

[Myl02] Jussi Myllymaki. Effective Web data extraction with standard XML technologies. Computer Networks (Amsterdam, Netherlands: 1999), 39(5):635– 644, August 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/
gej-ng/10/15/22/96/52/37/
abstract.html.

Naccarato:2002:XQX

[Nac02] Giuseppe Naccarato. XSLT querying and XML documents. Dr. Dobbs Journal, 27(12):24, 26, 28, December 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/documents/s=7718/
                                        ddj0212c/.

Nguyen:2001:MXD

[NACP01] Benjamin Nguyen, Serge Abiteboul, Grégory Cobena, and Mihaí Preda. Monitoring XML data on the Web. In Sellis and Mehrotra [SM01], pages 437– 448. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Naedele:2003:SSX

[Nae03]  Martin Naedele.  Security:

Standards for XML and Web services security. Computer, 36(4):96–98, April 2003. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/co/
                                        2003/04/r4096.htm;   http:
                                        //csdl.computer.org/dl/mags/
                                        co/2003/04/r4096.pdf.

Naylor:2002:IBO

[Nay02a] Bill Naylor. Interactions between OpenMath and MathML under the Maple environment. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Naylor:2002:MBP

[Nay02b] Bill Naylor. Mappings between presentation markup and semantic markup for variablesize objects. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Neven:2002:ALX

[Nev02] Frank Neven. Automata, logic, and XML. Lecture Notes in Computer Science, 2471:2– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2471/24710002.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2471/24710002.
pdf.

Ng:2002:MCI

[Ng02] Wilfred Ng. Maintaining consistency of integrated XML trees. Lecture Notes in Computer Science, 2419:145–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2419/24190145.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2419/24190145.
pdf.

Niederst:2000:HPR

[Nie00]  Jennifer Niederst.  HTML

pocket reference. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-579-3. iii + 92 pp. LCCN QA76.76 .N54 2000.

Niederst:2001:LWD

[Nie01] Jennifer Niederst. Learning Web design: a beginner’s guide to HTML, graphics, and beyond. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00036-

7. xii + 388 pp. LCCN TK5105.888.N545 2001.

Niederst:2002:HPR

[Nie02a]  Jennifer Niederst.  HTML

Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00296-

3. v + 95 pp. LCCN QA76.76.H94 N53 2002 Stacks. US$12.95. URL http://

                                        safari.oreilly.com/0596002963;
                                         http://www.oreilly.com/
                                        catalog/htmlpr2.

Niederst:2002:WDC

[Nie02b] Jennifer Niederst. The Web design CD bookshelf: six bestselling books on CD-ROM, 2002. ISBN 0-596-00271-8.

Nierman:2002:PPD

[NJ02] Andrew Nierman and H. V. Jagadish. ProTDB: Probabilistic data in XML. In Bernstein et al. [B^{+}02], pages 646–657.

ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.
org/conf/2002/S18P03.pdf.

Nambiar:2002:EXD

[NLB^{+}02] Ullas Nambiar, Zo’e Lacroix, Stéphane Bressan, Mong Li Lee, and Ying Guang Li. Efficient XML data management: An analysis. Lecture Notes in Computer Science, 2455:87– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2455/24550087.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2455/24550087.
pdf.

Norvaag:2002:ATQ

[Nør02] Kjetil Nørvåg. Algorithms for temporal query operators in XML databases. Lecture Notes in Computer Science, 2490:169–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2490/24900169.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2490/24900169.pdf.

Nachouki:2002:EIA

[NQ02] Gilles Nachouki and Mohamed Quafafou. Extracting, interconnecting, and accessing heterogeneous data sources:

An XML query based approach. Lecture Notes in Computer Science, 2510:442– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2510/25100442.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2510/25100442.pdf.

Naciri:2002:FMP

[NR02] Hanane Naciri and Laurence Rideau. Formal mathematical proof explanations in natural language using MathML: An application to proofs in Arabic. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Navarro:2000:MX

[NWB00] Ann Navarro, Chuck White, and Linda Burman. Mastering XML. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2000. ISBN 0-7821-2266-3 (paperback). xxxviii + 882 pp. LCCN QA76.76.H94 N3853

2000. URL http://corpitk.
earthweb.com/reference/0782122663.
html.

Neumann:2002:PBD

[NZ02] Gustaf Neumann and Uwe Zdun. Pattern-based design and implementation of an XML and RDF parser and interpreter: A case study. Lecture Notes in Computer Science, 2374:392–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2374/23740392.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2374/23740392.
pdf.

Obasanjo:2003:XRD

[Oba03] Dare Obasanjo. XML and relational databases. Dr. Dobbs Journal, 28(5):54, 56–58, May 2003. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/documents/s=7927/
ddj0305e/.

Oh:2002:ITM

[OJCH02] Yoon-Jung Oh, Hong-Taek Ju, Mi-Jung Choi, and James Won-Ki Hong. Interaction translation methods for XML/SNMP gateway. Lecture Notes in Computer Science, 2506:54– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2506/25060054.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2506/25060054.pdf.

Perez:2002:XXB

[PAB02] J. M. Pérez, M. J. Aramburu, and R. Berlanga. XRL: A XML-based query language for advanced services in digital libraries. Lecture Notes in Computer Science, 2453:300– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2453/24530300.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2453/24530300.
pdf.

Padovani:2002:SAR

[Pad02]  Luca Padovani.  A stand-

alone rendering engine for MathML. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Paparizos:2003:TNS

[PAKC^{+}03] Stelios Paparizos, Shurug Al-Khalifa, Adriane Chapman, H. V. Jagadish, Laks V. S. Lakshmanan, Andrew Nierman, Jignesh M. Patel, Divesh Srivastava, Nuwee Wiwatwattana, Yuqing Wu, and Cong Yu. TIMBER: a native system for querying XML. In ACM [ACM03a], page 672. ISBN ???? LCCN ????

Paparizos:2002:GX

[PAKJ^{+}02] Stelios Paparizos, Shurug Al-Khalifa, H. V. Jagadish, Laks Lakshmanan, Andrew Nierman, Divesh Srivastava, and Yuqing Wu. Grouping in XML. Lecture Notes in Computer Science, 2490:128–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/

service/series/0558/bibs/
2490/24900128.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2490/24900128.pdf.

Pankowski:2002:XSX

[Pan02] Tadeusz Pankowski. XML-SQL: An XML query language based on SQL and path tables. Lecture Notes in Computer Science, 2490:184– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2490/24900184.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2490/24900184.pdf.

Pawson:2002:XFM

[Paw02] Dave Pawson. XSL-FO: Making XML Look Good in Print. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00355-

2. xv + 263 pp. LCCN QA76.76.H94 P38 2002 Stacks. US$34.95. URL http://

safari.oreilly.com/0596003552;
http://www.oreilly.com/
catalog/xslfo.

Payne:2002:TCS

[Pay02] John Payne. Third commentary on “XML and the new design regime”. ACM Journal of Computer Documentation, 26 (2):48–49, May 2002. CODEN AJCDBH. ISSN 1527-6805. See [Wil02b].

Park:2002:XQP

[PCK02] Sangwon Park, Yoonra Choi, and Hyoung-Joo Kim. XML query processing using signature and DTD. Lecture Notes in Computer Science, 2455:162– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2455/24550162.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2455/24550162.
pdf.

Phillips:2002:TLM

[PD02] Ivor Phillips and Stan Devitt. TEX and LaTEX in a MathML context. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Paepen:2002:UXR

[PE02] Bert Paepen and Jan Engelen. Using XML as a reading enabler for visually impaired persons. Lecture Notes in Computer Science, 2398:382– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2398/23980382.
htm;   http://link.springer-
ny.com/link/service/series/

0558/papers/2398/23980382.
pdf.

Polyzotis:2002:SSG

[PG02a] Neoklis Polyzotis and Minos Garofalakis. Statistical synopses for graph-structured XML databases. In Franklin et al. [FMA02], pages 358–369. ISBN ???? LCCN ???? ACM order number 475020.

Polyzotis:2002:SVS

[PG02b] Neoklis Polyzotis and Minos N.

Garofalakis. Structure and value synopses for XML data graphs. In Bernstein et al. [B^{+}02], pages 466–477. ISBN 1-55860-869-9. LCCN ???? URL

http://www.vldb.org/conf/
2002/S13P03.pdf.

Phillips:2001:VVW

[Phi01] Lee Anne Phillips. VoiceXML and the voice/Web environment. Dr. Dobbs Journal, 26 (10):91–92, 94, 96, October 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2001/2001_10/
voicexml.zip;   http://www.
ddj.com/ftp/2001/2001_10/
voicexml_txt.zip.

Passi:2002:MXS

[PLM^{+}02] Kalpdrum Passi, Louise Lane, Sanjay Madria, Bipin C. Sakamuri, Mukesh Mohania, and Sourav Bhowmick. A model for XML schema integration. Lecture Notes in Computer Science, 2455:193– ??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2455/24550193.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2455/24550193.
pdf.

Park:2002:SFI

[PMC02] Chang-Won Park, Jun-Ki Min, and Chin-Wan Chung. Structural function inlining technique for structurally recursive XML queries. In Bernstein et al. [B^{+}02], pages 83– 94. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.
org/conf/2002/S04P01.pdf.

Perkins:2006:GEP

[PMK^{+}06] E. Perkins, M. Matsa, M. G. Kostoulas, A. Heifets, and N. Mendelsohn. Generation of efficient parsers through direct compilation of XML Schema grammars. IBM Systems Journal, 45(2):225–244, ???? 2006. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/452/perkins.html.

Porter:2003:MDX

[Por03] Brett Porter. Managing devices with XML-RPC. Dr. Dobbs Journal, 28(4):66, 68–70, April 2003. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/documents/s=7827/
ddj0304h/.

Powell:2001:HCR

[Pow01] Thomas A. Powell. HTML: the complete reference. Osborne/McGraw-Hill, Berkeley, CA, USA, third edition, 2001. ISBN 0-07-212951-4. xxiv + 1208 pp. LCCN QA76.76.H94 P684 2001.

Petropoulos:2002:BXQ

[PPV02] Michalis Petropoulos, Yannis Papakonstantinou, and Vasilis Vassalos. Building XML query forms and reports with XQForms. Computer Networks (Amsterdam, Netherlands: 1999), 39(5):541–558, August 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/
gej-ng/10/15/22/96/52/32/
abstract.html.

Pedersen:2002:CME

[PRP02] Dennis Pedersen, Karsten Riis, and Torben Bach Pedersen. Cost modeling and estimation for OLAP-XML federations. Lecture Notes in Computer Science, 2454:245– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2454/24540245.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2454/24540245.
pdf.

Psaila:2002:EEI

[Psa02a] Giuseppe Psaila. ERX: An experience in integrating entityrelationship models, relational databases, and XML technologies. Lecture Notes in Computer Science, 2490:242– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2490/24900242.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2490/24900242.pdf.

Psaila:2002:EQQ

[Psa02b] Giuseppe Psaila. ERX-QL: Querying an entity-relationship DB to obtain XML documents. Lecture Notes in Computer Science, 2397:281– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2397/23970281.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2397/23970281.pdf.

Polak:2002:MPX

[PSK02] S. Polak, R. Slota, and J. Kitowski. Message passing in XML-based language for creating multimedia presentations (research note). Lecture Notes in Computer Science, 2400:826–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-
ny.com/link/service/series/
0558/bibs/2400/24000826.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2400/24000826.
pdf.

Powell:2001:HPR

[PW01] Thomas A. Powell and Dan Whitworth. HTML: programmer’s reference. Osborne/McGraw-Hill, Berkeley, CA, USA, second edition, 2001. ISBN 0-07-213232-9. xi + 436 pp. LCCN QA76.76.H94 P495 2001.

Peng:2002:XBD

[PWK02] Xu Peng, Yang Wenjun, and Wang Kehong. XML-based data rendering engine for Web and mobile application scenarios. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15 (4):433–??, 2002. CODEN ???? ISSN 0219-4279.

Qu:2002:TOS

[QN02] Changtao Qu and Wolfgang Nejdl. Towards open standards: The evolution of an XML/JSP/WebDAV based collaborative courseware generating system. Lecture Notes in Computer Science, 2436:167– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2436/24360167.
htm;   http://link.springer-

                                        ny.com/link/service/series/
                                        0558/papers/2436/24360167.
                                        pdf.

Quin:2000:OSX

[Qui00] Liam Quin. Open source XML database toolkit: resources and techniques for improved development. Wiley, New York, NY, USA, 2000. ISBN 0-471-37522-5 (paperback). xiv + 434 pp. LCCN QA76.76.H94 Q56 2000.

Quint:2002:MLA

[QV02] Vincent Quint and Irène Vatton. MathML in E-learning with Amaya. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Rahman:2001:AVW

[Rah01] Sharif Md Abidur Rahman. Application of VRML in Web design: a 3D Webspace for Hotel Sonargaon. Project (M. Arch.), UCLA School of Architecture, Los Angeles, CA, USA, 2001.

Ramanan:2003:CIX

[Ram03] Prakash Ramanan. Covering indexes for XML queries: Bisimulation — simulation = negation. In Freytag et al. [FLA^{+}03], pages 165– 176. ISBN 0-12-722442-4.

                                        URL http://www.vldb.org/
                                        dblp/db/indices/a-tree/r/
                                        Ramanan:Prakash.html.

Ray:2001:LXC

[Ray01] Erik T. Ray. Learning XML: Creating Self-Describing Data. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00046-4. xii + 354 pp. LCCN QA76.76.H94 R3945 2001. US$34.95.

Roth:2006:XMT

[RHC^{+}06] M. Roth, M. A. Hernandez, P. Coulthard, L. Yan, L. Popa, H. C.-T. Ho, and C. C. Salter. XML mapping technology: Making connections in an XML-centric world. IBM Systems Journal, 45(2): 389–409, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/452/roth.html.

Ray:2002:PX

[RM02] Erik T. Ray and Jason McIntosh. Perl and XML. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00205-X. xii + 202 pp. LCCN QA76.73.P22 R39 2002 Stacks.

Rao:2006:SXD

[RM06] Praveen Rao and Bongki Moon.

Sequencing XML data and query twigs for fast pattern matching. ACM Transactions on Database Systems, 31(1): 299–345, March 2006. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Rockwell:2001:XXJ

[Roc01] Westy Rockwell. XML, XSLT, Java and JSP: a case study in developing a Web application. New Riders Publishing, Carmel, IN, USA, 2001. ISBN 0-7357-1089-9. xviii + 746 pp. LCCN QA76.76.H94 R62 2000.

Rollins:2000:AXA

[Rol00] Sami Rollins. Audio XmL: aural interaction with XML documents. Thesis (m.s.), University of California, Santa Barbara, Santa Barbara, CA, USA, 2000.

Runapongsa:2002:SQX

[RP02] Kanda Runapongsa and Jignesh M. Patel. Storing and querying XML data in object-relational DBMSs. Lecture Notes in Computer Science, 2490:266–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/
                                        service/series/0558/bibs/
                                        2490/24900266.htm; http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2490/24900266.pdf.

Ray:2000:HDQ

[RR00] Deborah S. Ray and Eric J. Ray. HTML 4 for dummies quick reference. IDG Books, San Mateo, CA, USA, second edition, 2000. ISBN 0-7645-0721-4. xii + 224 pp. LCCN QA76.76.H94 R38 2000.

Ramalho:2003:XXL

[RRB03] F. Ramalho, J. Robin, and R. Barros. XOCL — an XML language for specifying logical constraints in object oriented models. J.UCS: Journal of Universal Computer Science, 9(8):956–??, August 28, 2003. CODEN ???? ISSN 0948-

6968. URL http://www.jucs.
org/jucs_9_8/xocl_an_xml_
language.

Rollins:2000:ACA

[RS00] Sami Rollins and Neel Sundaresan. AVoN calling: AXL for voice-enabled Web navigation. Computer Networks (Amsterdam, Netherlands: 1999), 33 (1–6):533–551, June 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/15/22/48/25/
62/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/15/22/48/25/62/article.
pdf;   http://www9.org/w9cdrom/
55/55.html.

Raynal:2002:CHB

[RS02] Mathieu Raynal and Mathieu Serrurier. CYNTHIA: An HTML browser for visually handicapped people. Lecture Notes in Computer Science, 2398:353–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-

                                        ny.com/link/service/series/
                                        0558/bibs/2398/23980353.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2398/23980353.
                                        pdf.

Rodionov:2002:TTT

[RW02a] Igor Rodionov and Stephen Watt. Tool for translating TEX/LaTEX to MathML. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Rolsky:2002:EPH

[RW02b] Dave Rolsky and Ken Williams. Embedding Perl in HTML with Mason. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00225-4. xviii + 297 pp. LCCN QA76.73.P22 R65 2002. US$34.95. URL http:/

                                        /www.oreilly.com/catalog/
                                        perlhtmlmason.

Synodinos:2003:LHE

[SA03] D. G. Synodinos and P. Avgeriou. Leveraging hypermedia engineering for multi-channel access with the use of XML. International Journal of Computer Systems Science and Engineering, 18(5):??, September 2003. CODEN CSSEEI. ISSN 0267-6192.

Sahuguet:2001:KMT

[Sah01] Arnaud Sahuguet. Kweelt: more than just “yet another framework to query XML!”. In Sellis and Mehrotra [SM01], page 602. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Sandhu:2003:MH

[San03] Pavi Sandhu. The MathML handbook. Charles River Media, Hingham, MA, USA, 2003. ISBN 1-58450-249-5. xviii + 518 pp. LCCN QA41.6 .S36 2002. US$41.95, CAN$64.95.

Savourel:2001:XIL

[Sav01] Yves Savourel. XML Internationalization and Localization. SAMS Publishing, Indianapolis, IN, USA, 2001. ISBN 0-672-32096-7. xi + 519 pp. LCCN QA76.76.H94 S252 2001.

Spencer:2000:PWV

[SB00] Stephen N. Spencer and Don Brutzman, editors. Proceedings Web3D — VRML 2000: Fifth Symposium on the Virtual Reality Modeling Language, Monterey, Calif., February 21–24, 2000. ACM Press, New York, NY, USA, 2000. ISBN 1-58113-211-5. URL

http://www.acm.org/pubs/
contents/proceedings/graph/
330160.

Schrefl:2002:AXS

[SB02] Michael Schrefl and Martin Bernauer. Active XML schemas. Lecture Notes in Computer Science, 2465:363– ??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2465/24650363.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2465/24650363.pdf.

Sundaram:2002:XBP

[SC02] Babu Sundaram and Barbara M. Chapman. XMLbased policy engine framework for usage policy management in Grids. Lecture Notes in Computer Science, 2536:194– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2536/25360194.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2536/25360194.pdf.

Snoeren:2001:MBC

[SCG01] Alex C. Snoeren, Kenneth Conley, and David K. Gifford. Mesh-based content routing using XML. Operating Systems Review, 35(5):160–173, December 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schmitt:2002:ASM

[Sch02] Hermann Schmitt. An application server for Mathematica. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

StLaurent:2000:XMT

[SD00] Simon St. Laurent and B. K.

Delong. XHTML: Moving Toward XML. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 2000. ISBN 0-7645-4709-7. xx + 435 pp. LCCN QA76.76.H94 S725

2000. URL http://164.109.
153.102/product.asp?isbn=
0764547097.

Sunderland:2004:FXB

[SDC04] R. M. Sunderland, R. I.

Damper, and R. M. Crowder. Flexible XML-based configuration of physical simulations. Software—Practice and Experience, 34(12):1149–1155, October 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Searls:2000:NBE

[Sea00] Doc Searls. The next bang:

The explosive combination of embedded Linux, XML and instant messaging. Linux Journal, 77:??, September 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.
com/lj-issues/issue77/4195.
html.

Seely:2002:SCP

[See02]  Scott Seely.  SOAP: cross

platform Web service development using XML. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-090763-4. xiv + 391 pp. LCCN QA76.73.S62 S44 2002. Includes CD-ROM.

Segoufin:2003:TQX

[Seg03] Luc Segoufin. Typing and querying XML documents: some complexity bounds. In ACM [ACM03b], pages 167– 178. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Selinger:2002:IIX

[Sel02] Patricia G. Selinger. Information integration and XML in IBM’s DB2. In Bernstein et al. [B^{+}02], pages 906–907. ISBN 1-55860-869-9. LCCN

                                        ???? URL http://www.vldb.
                                        org/conf/2002/S27P01.pdf.

Semeczko:2002:BRS

[Sem02] George Semeczko. Book review: Securing XML: A review of XML Security by Blake Dournaee. IEEE Distributed Systems Online, 3(10), 2002. ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0210/d/bks_c.htm.

StLaurent:2005:XPR

[SFE05] Simon St. Laurent, Michael Fitzgerald, and Robert Eckstein. XML pocket reference. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2005. ISBN 0-596-10050-7 (paperback). iv + 171 pp. LCCN QA76.76.H94 E25 2005.

Shen:2002:JBD

[SG02a] J. Shen and G. Gu. Javabased design and implementation of the XML parser. Mini-Micro Systems, 23(12): 1449–1452, 2002. CODEN XWJXEH. ISSN 1000-1220.

Skonnard:2002:EXQ

[SG02b] Aaron Skonnard and Martin Gudgin. Essential XML quick reference: a programmer’s reference to XML, XPath, XSLT, XML Schema, SOAP, and more. DevelopMentor series. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-74095-

8. xxv + 402 pp. LCCN QA76.76.H94 S59 2002.

Shegalov:2001:XEW

[SGW01] German Shegalov, Michael Gillmann, and Gerhard Weikum. XML-enabled workflow management for e-services across heterogeneous platforms. VLDB Journal: Very Large Data Bases, 10(1):91–103, August 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//link.springer.de/link/
service/journals/00778/bibs/
1010001/10100091.htm;   http:
//link.springer.de/link/
service/journals/00778/papers/
1010001/10100091.pdf.

Stephens:2002:VBN

[SH02] Rod Stephens and Brian Hochgurtel. Visual Basic .NET and XML: harness the power of XML in VB.NET applications. Wiley, New York, NY, USA, 2002. ISBN 0-471-12060-X. xxiv + 503 pp. LCCN QA76.73.B3 S8342 2002.

Sidje:2002:MAO

[Sid02]  Roger B. Sidje.  MathML

amidst open Web standards: Mozilla’s building blocks for today and tomorrow. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Simon:2001:X

[Sim01a] Solomon H. Simon. XML. Emerging business technology series. McGraw-Hill, New York, NY, USA, 2001. ISBN 0-07-137188-5. xxviii + 259 pp. LCCN HD30.37 .S53 2001.

Simpson:2001:JX

[Sim01b]  John E. Simpson.  Just

XML. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 2001. ISBN 0-13-018554-X. xv + 417 pp. LCCN QA76.76.H94 S57 2001.

Simpson:2002:JX

[Sim02a] John E. Simpson. Just XSL. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-060311-2 (paperback). xix + 595 pp. LCCN QA76.76.H94 S573 2002.

Simpson:2002:XXL

[Sim02b] John E. Simpson. XPath and XPointer: Locating Content in XML Documents. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00291-2. x + 196 pp. LCCN QA76.76.H94 S578 2002 Stacks. US$24.95. URL http:/

/safari.oreilly.com/0596002912;
http://www.oreilly.com/
catalog/xpathpointer.

Sintes:2000:XSC

[Sin00] Tony Sintes. XML and software configuration. Dr. Dobbs Journal, 25(7):56, 58–62, July 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2000/2000_07/
xmlcfg.txt;   http://www.
ddj.com/ftp/2000/2000_07/
xmlcfg.zip.

StLaurent:2001:PWS

[SJD01] Simon St. Laurent, Joe Johnston, and Edd Dumbill. Programming Web services with XML-RPC. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00119-3. 213 pp. LCCN QA76.76.H94 S7 2001.

Sharma:2002:VST

[SK02] Chetan Sharma and Jeff Kunins. VoiceXML: strategies and techniques for effective voice application development with VoiceXML 2.0. Professional developer’s guide series. Wiley, New York, NY, USA, 2002. ISBN 0-471-41893-5 (paperback). xxv + 469 pp. LCCN QA76.76.H94 S259 2002 Bar. Includes CD-ROM.

Sellis:2001:PAS

[SM01] Timos Sellis and Sharad Mehrotra, editors. Proceedings of the 2001 ACM SIG-MOD International Conference on Management of Data 2001, Santa Barbara, California, United States, May 21–24, 2001. ACM Press, New York, NY, USA, 2001. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Sundaresan:2002:APM

[SM02] Neel Sundaresan and Reshad Moussa. Algorithms and programming models for efficient representation of XML for Internet applications. Computer Networks (Amsterdam, Netherlands: 1999), 39(5):681–697, August 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/
gej-ng/10/15/22/96/52/40/
abstract.html.

Sonneck:2002:MUW

[SMM02] Georg Sonneck, Renate Motschnig, and Thomas Mueck. Mapping UML Web navigation stereotypes to XML data skeletons. Lecture Notes in Computer Science, 2455:294–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2455/24550294.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2455/24550294.
pdf.

Schengili-Roberts:2000:CC

[SR00] Keith Schengili-Roberts. Core CSS. Prentice Hall PTR core series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-083456-4 (paperback). xxxvii + 635 + 4 pp. LCCN TK5105.888. S32 2000.

Shabo:2006:RIX

[SRCV06] A. Shabo, S. Rabinovici-Cohen, and P. Vortman. Revolutionary impact of XML on biomedical information interoperability. IBM Systems Journal, 45 (2):361–372, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/452/shabo.html.

Scribner:2002:ASI

[SS02] Kenn Scribner and Mark C.

Stiver. Applied SOAP: implementing .NET XML Web services. SAMS Publishing, Indianapolis, IN, USA, 2002. ISBN 0-672-32111-4. ix + 415 pp. LCCN QA76.76.H94 S2575 2002.

Shanmugasundaram:2001:EPR

[SSB^{+}01] Jayavel Shanmugasundaram, Eugene Shekita, Rimon Barr, Michael Carey, Bruce Lindsay, Hamid Pirahesh, and Berthold Reinwald. Efficiently publishing relational data as XML documents. VLDB Journal: Very Large Data Bases, 10(2–3):133– 154, September 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL http://link.springer.
de/link/service/journals/
00778/bibs/1010002/10100133.
htm;   http://link.springer.
de/link/service/journals/
00778/papers/1010002/10100133.
pdf.

Samwel:2000:LDS

[SSC^{+}00] Bart Samwel, Jiri Soukup, Glenn Crist, Evan Easton, Ron Ruble, David A. Rogers, Al Stevens, Bruce MacDonald, and Scott Venckus. Letters: Data structures as objects; real (Netscape) time; riding the XML bandwagon; porting to CE; nothing new about Open Source; Y2K worries?; version control. Dr. Dobbs Journal, 25 (2):12, 14, February 2000. CO-DEN DDJOEB. ISSN 1044-789X.

StLaurent:2000:XES

[St.00] Simon St. Laurent. XML elements of style. McGraw-Hill, New York, NY, USA, 2000. ISBN 0-07-212220-X. xiii + 298 pp. LCCN QA76.76.H94 S724

Standefer:2000:EXC

[Sta00] Robert Standefer. Enterprise XML Clearly Explained. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2000. ISBN 0-12-663355-X. xi + 321 pp. LCCN QA76.76.H94 S82 200. US$44.95.

Standefer:2001:EXC

[Sta01] Robert Standefer. Enterprise XML clearly explained. Academic Press, New York, USA, 2001. ISBN 0-12-663355-X (paperback). xi + 321 pp. LCCN QA76.76.H94 S82 2001.

Stayton:2003:DXC

[Sta03] Bob Stayton. DocBook XSL:

The Complete Guide. Sagehill Enterprises, PO Box 2911, Santa Cruz CA 95063-2911, USA, 2003. ISBN 0-9741521-1-0. xvi + 394 pp. LCCN ????

URL http://www.sagehill.
net/docbookxsl/index.html.

Stayton:2006:DXC

[Sta06] Bob Stayton. DocBook XSL:

The Complete Guide. Sagehill Enterprises, PO Box 2911, Santa Cruz CA 95063-2911, USA, 2006. ISBN 0-9741521-2-9. 560 (est.) pp. LCCN ???? US$49.95.

Stayton:2007:DXC

[Sta07] Bob Stayton. DocBook XSL:

the Complete Guide. Sagehill Enterprises, PO Box 2911, Santa Cruz CA 95063-2911, USA, fourth edition, 2007.

ISBN 0-9741521-3-7. 588 (est.) pp. LCCN ????

Snell:2002:PWS

[STK02] James Snell, Doug Tidwell, and Pavel Kulchenko. Programming Web Services with SOAP. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00095-2. xiii + 244 pp. LCCN QA76.625 .S64 2002 Bar. US$34.95. URL http://

safari.oreilly.com/0596000952;
 http://www.oreilly.com/
catalog/progwebsoap.

Strobel:2002:XSR

[Str02] Michael Ströbel. An XML schema representation for the communication design of electronic negotiations. Computer Networks (Amsterdam, Netherlands: 1999), 39(5):661–680, August 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/
gej-ng/10/15/22/96/52/39/
abstract.html.

Sturm:2000:DXS

[Stu00] Jake Sturm. Developing XML solutions. Microsoft Press, Bellevue, WA, USA, 2000. ISBN 0-7356-0796-6. xiii + 415 pp. LCCN QA76.76.H94 S748

Suciu:2002:STQ

[Suc02]  Dan Suciu.  From search-

ing text to querying XML streams. Lecture Notes in Computer Science, 2476:11– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2476/24760011.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2476/24760011.pdf.

Segoufin:2002:VSX

[SV02] Luc Segoufin and Victor Vianu.

Validating streaming XML documents. In ACM [ACM02a], pages 53–64. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Sipani:2004:DHP

[SVMAM04] Sudhanshu Sipani, Kunal Verma, John A. Miller, and Boanerges Aleman-Meza. Designing a high-performance database engine for the ‘Db4XML’ native XML database system. The Journal of Systems and Software, 69(1–2):87– 104, January 1, 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sauers:2001:XE

[SW01] Michael Sauers and R. Allen Wyke. XHTML essentials. Wiley, New York, NY, USA, 2001. ISBN 0-471-41764-5 (paperback). xvi + 383 pp. LCCN QA76.76.H94 S24 2001.

Simeon:2002:EXP

[SW02a] Jérôme Siméon and Philip Wadler. The essence of XML (preliminary version). Lecture Notes in Computer Science, 2441:21–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/
                                        service/series/0558/bibs/
                                        2441/24410021.htm; http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2441/24410021.pdf.

Smirnova:2002:MTC

[SW02b] Elena Smirnova and Stephen Watt. MathML to TEX conversion: Conserving high-level semantics in translation. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Simeon:2003:EX

[SW03] Jérôme Siméon and Philip Wadler. The essence of XML. ACM SIGPLAN Notices, 38 (1):1–13, January 2003. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schmidt:2002:XBX

[SWK^{+}02] Albrecht Schmidt, Florian Waas, Martin L. Kersten, Michael J. Carey, Ioana Manolescu, and Ralph Busse. XMark: A benchmark for XML data management. In Bernstein et al. [B^{+}02], pages 974–985. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.
org/conf/2002/S30P01.pdf.

Seng:2004:TMG

[SY04] Jia-Kang Seng and Jing Yu.

Toward a more generic design of XML benchmark workload. International Journal of Computer Systems Science and Engineering, 19(5):??, September 2004. CODEN CSSEEI. ISSN 0267-6192.

Syed:2002:PUN

[Sye02] Safdar Syed. The powerful use of new database technologies in indexing mathematics and displaying dynamic math on the Web. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Toshniwal:2004:TRM

[TA04] Rishi Toshniwal and Dharma P.

Agrawal. Tracing the roots of markup languages. Communications of the ACM, 47(5): 95–98, May 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tamura:2000:DWP

[Tam00] Randall A. Tamura. Domino 5 Web Programming With XML, Java and JavaScript. Que Corporation, Indianapolis, IN, USA, 2000. ISBN 0-7897-2275-5. xvi + 912 pp. LCCN QA76.625 .T36 2000. US$50.

Tannenbaum:2002:MSU

[Tan02] Adrienne Tannenbaum. Metadata solutions: using metamodels, repositories, XML, and enterprise portals to generate information on demand. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-71976-2. xxv + 490 pp. LCCN T58.5 .T37 2002.

Turner:2000:HJP

[TB00] Shane E. Turner and Karl Barksdale. HTML and JavaScript: programming concepts. International Thomson Publishing, 5101 Madison Road, Cincinnati, OH 45227, USA; Bonn, Germany, 2000. ISBN 0-538-68822-X. xi + 154 pp. LCCN ????

Thiruvathukal:2004:XCS

[TB04] George K. Thiruvathukal and David Beazley. XML and computational science. Computing in Science and Engineering, 6(1):74–80, January/February 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/comp/
mags/cs/2004/01/c1074abs.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/01/c1074.
pdf.

Tan:2002:MID

[TD02] Peter J. Tan and David L. Dowe. MML inference of decision graphs with multiway joins. Lecture Notes in Computer Science, 2557:131– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2557/25570131.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2557/25570131.pdf.

Teague:2001:DCW

[Tea01] Jason Cranford Teague. DHTML and CSS for the World Wide Web. Visual quickstart guide. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, second edition, 2001. ISBN 0-201-73084-7. xxii + 592 pp. LCCN QA76.76.H94 T428 2001.

Todd:2001:LSS

[TEM^{+}01] Andrew W. Todd, Jonathan Erickson, Nadine McKenzie, Chris Cleeland, Richard Huang, Ragae Ghaly, and The Editors. Letters: Shared source and shared secrets; JavaScript fix; CORBA interoperability; EJB application servers update; correction [“The Delphi XML SAX2 Component and MSXML 3.0”]. Dr. Dobbs Journal, 26(10):10, 12, October 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www. ddj.com/. See [Hei01].

Thuraisingham:2002:XDS

[Thu02] Bhavani Thuraisingham. XML databases and the Semantic Web. CRC Press, 2000 N.W.

Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2002. ISBN 0-8493-1031-8. 24 + 306 pp. LCCN QA76.9.D3 T4583 2002. US$79.95.

Tidwell:2001:X

[Tid01] Doug Tidwell. XSLT. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00053-7. xvi + 460 pp. LCCN QA76.73.X58 T53 2001.

Tidwell:2007:XMX

[Tid07] Doug Tidwell. XSLT: mastering XML transformations. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2007. ISBN 0-596-52721-7 (paperback). xviii + 965 pp. LCCN ???? US$31.99.

Tatarinov:2001:UX

[TIHW01] Igor Tatarinov, Zachary G. Ives, Alon Y. Halevy, and Daniel S. Weld. Updating XML. In Sellis and Mehrotra [SM01], pages 413–424. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Tittel:2002:MX

[Tit02] Ed Tittel. Mastering XHTML.

Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2002. ISBN 0-7821-2818-1. xxviii + 1264 pp. LCCN QA76.76.H94 M2795 2002.

Thiruvathukal:2004:NXD

[TL04] George K. Thiruvathukal and Konstantin Läufer. Natural XML for data binding, processing, and persistence. Computing in Science and Engineering, 6(2):86–92, March/ April 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/comp/
mags/cs/2004/02/c2086abs.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/02/c2086.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/02/c2086.
pdf.

Tang:2005:BDM

[TMK05] Zhaohui Tang, Jamie Maclennan, and Peter Pyungchul Kim. Building data mining solutions with OLE DB for DM and XML for analysis. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(2):80–85, June 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tanoue:2002:GBS

[TMYU02] Junko Tanoue, Noboru Matoba, Masatoshi Yoshikawa, and Shunsuke Uemura. GeneAround: A browsing system for gene annotation using XML technologies. Lecture Notes in Computer Science, 2419:236– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2419/24190236.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2419/24190236.
pdf.

Travis:2000:XDN

[Tra00] Brian E. Travis. XML for developers: November 6–8, 2000, Engineering 819.323. Los Angeles, CA, USA, 2000. various pp.

Turau:2002:CSW

[Tur02a] Volker Turau. A caching system for Web content generated from XML sources using XSLT. Lecture Notes in Computer Science, 2426:197– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2426/24260197.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2426/24260197.pdf.

Turner:2002:EGX

[Tur02b] Ronald C. Turner. The essential guide to XML technologies. Essential guide series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-065565-1. xviii + 385 pp. LCCN QA76.76.H94 T87 2002.

Tatarinov:2002:SQO

[TVB^{+}02] Igor Tatarinov, Stratis D. Viglas, Kevin Beyer, Jayavel Shanmugasundaram, Eugene Shekita, and Chun Zhang. Storing and querying ordered XML using a relational database system. In Franklin et al. [FMA02], pages 204–215. ISBN ???? LCCN ???? ACM order number 475020.

Theobald:2002:XSE

[TW02] Anja Theobald and Gerhard Weikum. The XXL search engine: ranked retrieval of XML data using indexes and ontologies. In Franklin et al. [FMA02], page 615. ISBN ???? LCCN ???? ACM order number 475020.

Toman:2005:RAS

[TW05] David Toman and Grant Weddell. On reasoning about structural equality in XML: a description logic approach. Theoretical Computer Science, 336 (1):181–203, May 25, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Umehara:2002:CBR

[UIN02] Masayuki Umehara, Koji Iwanuma, and Hidetomo Nabeshima. A case-based recognition of semantic structures in HTML documents. Lecture Notes in Computer Science, 2412:141– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2412/24120141.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2412/24120141.
pdf.

USOERI:2001:XLM

[Uni01] United States. Office of Educational Research and Improvement. XML: A Language to Manage the World Wide Web. Washington, DC, USA, March 3, 2001. Shipping List no 2001-0191-M.

Varela:2002:XKB

[VAS02] Leonilde Rocha Varela, Joaquim Nunes Aparício, and Sílvio Carmo Silva. An XML knowledge base system for scheduling problems. Lecture Notes in Computer Science, 2346:63– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2346/23460063.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2346/23460063.pdf.

vanderVlist:2002:XS

[vdV02] Eric van der Vlist. XML Schema. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel:

+1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00252-1. xiii + 379 pp. LCCN QA76.76.H94 V55 2002 Stacks.

Vela:2006:MDD

[VFMMP06] Belén Vela, Eduardo Fernández-Medina, Esperanza Marcos, and Mario Piattini. Model driven development of secure XML databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(3):22–27, September 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Vaughan-Nichols:2003:ITX

[VN03] Steven J. Vaughan-Nichols. Industry trends: XML raises concerns as it gains prominence. Computer, 36(5):14– 16, May 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2003/05/r5014.
htm;   http://csdl.computer.
org/dl/mags/co/2003/05/r5014.
pdf.

Vidakovic:2006:GCD

[VR06] Jovana Vidaković and Miloš Racković. Generating content and display of library catalogue cards using XML technology. Software—Practice and Experience, 36(5):513–524, April 25, 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Verstak:2003:BBS

[VRW^{+}03] Alex Verstak, Naren Ramakrishnan, Layne T. Watson, et al. BSML: A binding schema markup language for data interchange in problem solving environments. Scientific Programming, 11(3):199– 224, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wahlin:2002:XAN

[Wah02]  Dan Wahlin.  XML for

ASP.NET developers. SAMS Publishing, Indianapolis, IN, USA, 2002. ISBN 0-672-32039-8 (paperback). ix + 484 pp. LCCN QA76.76.H94 W339 2002.

Walmsley:2002:DXS

[Wal02] Priscilla Walmsley. Definitive XML schema. Charles F. Goldfarb definitive XML series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-065567-8. xxv + 528 pp. LCCN QA76.76.H94 W29 2002.

Westbomke:2002:TXB

[WD02] J. Westbomke and G. Dittrich. Towards an XML-based implementation of structured hypermedia documents. J.UCS: Journal of Universal Computer Science, 8(10):944–??, October 28, 2002. CODEN ???? ISSN 0948-6968. URL http:

//www.jucs.org/jucs_8_10/
towards_an_xlm_based.

Williamson:2002:DND

[WE02] Heather Williamson and Bruce A.

Epstein. Dreamweaver in a Nutshell: A Desktop Quick Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00239-4. xxvi + 442 pp. LCCN TK5105.8885.D74 W55 2002. US$29.95. URL http://

safari.oreilly.com/0596002394;
 http://www.oreilly.com/
catalog/dreamweavernut.

Wang:2002:DTH

[WH02] Yalin Wang and Jianying Hu.

Detecting tables in HTML documents. Lecture Notes in Computer Science, 2423:249– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2423/24230249.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2423/24230249.
pdf.

White:2001:DPP

[Whi01] Ken White. DBMS: Past, present, and future: Where databases have been, where they are going, and what it means to you. Dr. Dobbs Journal, 26(8):21–22, 24, 26, 28– 29, August 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/.

Wiggins:2000:SXI

[Wig00] Richard Wiggins. Standards: XML: An interview with Peter Flynn. Computer, 33(4):113– 114, 116, April 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.
                                        org/co/books/co2000/pdf/
                                        r4113.pdf.

Wilson:2000:PBC

[Wil00] Gregory V. Wilson. Programmer’s bookshelf: Classics old and new. Dr. Dobbs Journal, 25 (11):159–160, November 2000. CODEN DDJOEB. ISSN 1044-789X.

Williamson:2001:XCR

[Wil01a] Heather Williamson. XML: the complete reference. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-212734-1. xxvi + 965 pp. LCCN QA76.76.H94 W596 2001.

Wilson:2001:PBT

[Wil01b] Gregory V. Wilson. Programmer’s bookshelf: Time warps. Dr. Dobbs Journal, 26(2):159– 160, February 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/.

Wilson:2001:PBX

[Wil01c] Gregory V. Wilson. Programmer’s bookshelf: XML and then some. Dr. Dobbs Journal, 26(5):151–152, May 2001. CODEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/.

Wilkes:2002:RCS

[Wil02a] Gilbert Vanburen Wilkes IV.

Reply to commentaries on “XML and the new design regime”. ACM Journal of Computer Documentation, 26(2): 50–53, May 2002. CODEN AJCDBH. ISSN 1527-6805.

Wilkes:2002:XND

[Wil02b] Gilbert Vanburen Wilkes IV.

XML and the new design regime: disputes between designers, application developers, authors, and readers in changing technological conditions and perceptions of social and professional need. ACM Journal of Computer Documentation, 26(2):33–42, May 2002. CODEN AJCDBH. ISSN 1527-6805. See commentary [Bur02a, Hay02, Pay02, Wil02a].

Wilson:2003:PB

[Wil03a] Gregory V. Wilson. Programmer’s bookshelf. Dr. Dobbs Journal, 28(7):75, July 2003. CODEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/.

Wilson:2003:PBO

[Wil03b] Gregory V. Wilson. Programmer’s bookshelf: The observation of trifles. Dr. Dobbs Journal, 28(1):73, January 2003.

CODEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/documents/s=7758/
                                        ddj0301j/.

Wilson:2003:XBP

[Wil03c] Gregory V. Wilson. XMLbased programming systems. Dr. Dobbs Journal, 28(3):16, 18, 20, 22, 24, March 2003. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/documents/s=7826/
                                        ddj0303b/.

Wickham-Jones:2002:WMH

[WJ02] Tom Wickham-Jones. Web Mathematica: How to deliver computational and visualization services from a Web server. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Westermann:2003:AXD

[WK03] Utz Westermann and Wolfgang Klas. An analysis of XML database solutions for the management of MPEG-7 media descriptions. ACM Computing Surveys, 35(4):331– 373, December 2003. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Westermann:2006:PSA

[WK06] U. Westermann and W. Klas. PTDOM: a schema-aware XML database system for MPEG-7 media descriptions. Software— Practice and Experience, 36(8): 785–834, July 10, 2006. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wan:2002:ESX

[WL02] Changxuan Wan and Yunsheng Liu. Efficient supporting XML query and keyword search in relational database systems. Lecture Notes in Computer Science, 2419:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2419/24190001.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2419/24190001.
pdf.

Wen:2004:IDE

[WL04] Y. Wen and X. Liu. Intelligence data extraction based on Java XML and Web. Journal-Sichuan University Natural Science Edition, 41(2):318–321, 2004. CODEN ???? ISSN 0490-6756.

Wang:2003:VDI

[WPFY03] Haixun Wang, Sanghyun Park, Wei Fan, and Philip S. Yu. ViST: a dynamic index method for querying XML data by tree structures. In ACM [ACM03a], pages 110–121. ISBN ???? LCCN ????

Winkler:2002:SDS

[WS02] Karsten Winkler and Myra Spiliopoulou. Structuring domain-specific text archives by deriving a probabilistic XML DTD. Lecture Notes in Computer Science, 2431:461– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2431/24310461.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2431/24310461.pdf.

Wan:2002:IMH

[WW02] Bo Wan and Stephen Watt. An interactive mathematical handwriting recognizer for the Pocket PC. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Xie:2002:MXS

[XW02] Yuzhen Xie and Stephen Watt. A modular XML schema for MathML. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Xie:2002:LSB

[XWP02] Yuzhen Xie, Stephen Watt, and Luca Padovani. A Lisp subset based on MathML. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ????

Xu:2002:KDH

[XWW^{+}02] Haiyuan Xu, Quanyuan Wu, Huaimin Wang, Guogui Yang, and Yan Jia. KF-Diff+: Highly efficient change detection algorithm for XML documents. Lecture Notes in Computer Science, 2519:1273– ??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2519/25191273.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2519/25191273.pdf.

Xu:2002:XBD

[XYW02] Peng Xu, Wenjun Yang, and Kehong Wang. XMLbased data rendering engine for content management system. Lecture Notes in Computer Science, 2419:170–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2419/24190170.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2419/24190170.
pdf.

Yoshikawa:2001:XPB

[YASU01] Masatoshi Yoshikawa, Toshiyuki Amagasa, Takeyuki Shimura, and Shunsuke Uemura. XRel: a path-based approach to storage and retrieval of XML documents using relational databases. ACM Transactions on Internet Technology (TOIT), 1(1):110–141, August 2001. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Yang:2004:EME

[YF04] Michael Yang and Richard Fateman. Extracting mathematical expressions from PostScript documents. In Gutierrez [Gut04], pages 305– 311. ISBN 1-58113-827-X.

LCCN ???? URL http://www.
ocf.berkeley.edu/~mlyang/
papers/MichaelYangPsmath.
pdf.

Yeow:2002:EXP

[YHL02] Kelvin Yeow, R. Nigel Horspool, and Michael R. Levy. Expeditious XML processing. Lecture Notes in Computer Science, 2412:148–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2412/24120148.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2412/24120148.
pdf.

Yang:2002:WGU

[YKDC02] Jaeyoung Yang, Jungsun Kim, Kyoung-Goo Doh, and Joongmin Choi. Wrapper generation by using XML-based domain knowledge for intelligent information extraction. Lecture Notes in Computer Science, 2417:472–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2417/24170472.
htm;   http://link.springer-

ny.com/link/service/series/
0558/papers/2417/24170472.
pdf.

Yang:2003:EMX

[YLH03] Liang Huai Yang, Mong-Li Lee, and Wynne Hsu. Efficient mining of XML query patterns for caching. In Freytag et al. [FLA^{+}03], pages 69– 80. ISBN 0-12-722442-4. URL

http://www.vldb.org/dblp/
db/indices/a-tree/y/Yang:
Liang_Huai.html.

Yu:2005:MXD

[YLM^{+}05] Yijun Yu, Jianguo Lu, John Mylopoulos, Weiwei Sun, Jing-Hao Xue, and Erik H. D’Hollander. Making XML document markup international. Software—Practice and Experience, 35(1):1–14, January 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yu:2002:CAM

[YSLJ02] Ting Yu, Divesh Srivastava, Laks V. S. Lakshmanan, and H. V. Jagadish. Compressed accessibility map: Efficient access control for XML. In Bernstein et al. [B^{+}02], pages 478–489. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.
org/conf/2002/S14P01.pdf.

Yuan:2001:LPS

[Yua01] Michael Yuan. LaTeX2HTML:

Publish science to the Web. Linux Journal, 92:93–94, December 2001. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Yu:2003:AMN

[YWL^{+}03] J. X. Yu, G. Wang, H. Lu, G. Yu, J. Lv, and B. Sun. Automata match: a new XML query processing approach. International Journal of Computer Systems Science and Engineering, 18(5):??, September 2003. CODEN CSSEEI. ISSN 0267-6192.

Zhang:2003:RMX

[ZDW^{+}03] Xin Zhang, Katica Dimitrova, Ling Wang, Maged El Sayed, Brian Murphy, Bradford Pielech, Mukesh Mulchandani, Luping Ding, and Elke A. Rundensteiner. Rainbow: multi-XQuery optimization using materialized XML views. In ACM [ACM03a], page 671. ISBN ???? LCCN ????

Zhenhua:2003:BTS

[Zhe03] L. Zhenhua. A bus-type software model based on CORBA and XML. International Journal of Computer Systems Science and Engineering, 18(6): ??, December 2003. CODEN CSSEEI. ISSN 0267-6192.

Zhou:2003:TDC

[ZWG^{+}03] Aoying Zhou, Qing Wang, Zhimao Guo, Xueqing Gong, Shihui Zheng, Hongwei Wu, Jianchang Xiao, Kun Yue, and Wenfei Fan. TREX: DTDconforming XML to XML transformations. In ACM [ACM03a], page 670. ISBN ???? LCCN ????

Zheng:2002:SMN

[ZZY^{+}02] Shihui Zheng, Aoying Zhou, Jeffrey Xu Yu, Long Zhang, and Han Tao. Structural map: A new index for efficient XML path expression processing. Lecture Notes in Computer Science, 2419:25– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2419/24190025.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2419/24190025.
pdf.